Version Management with CVS

for cvs 1.11.23

Per Cederqvist et al Copyright \begin{matrix}c \\ ◯ \end{matrix} 1993, 1994, 1995, 1996, 1997, 1998, 1999, 2000, 2001, 2002, 2003, 2004, 2005, 2006, 2007, 2008 Free Software Foundation, Inc. Portions

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1999, 2000, 2001, 2002, 2003, 2004, 2005, 2006, 2007 Derek R. Price, Copyright \begin{matrix}c \\ ◯ \end{matrix} 2002, 2003, 2004, 2005, 2006, 2007 Ximbiot

           http://ximbiot.com,

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1992, 1993, 1999 Signum Support AB,

and Copyright \begin{matrix}c \\ ◯ \end{matrix} others.

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided also that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the Free Software Foundation.

Short Contents

1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 The Repository . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3 Starting a project with CVS . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 4 Revisions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 5 Branching and merging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 6 Recursive behavior . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 7 Adding, removing, and renaming files and directories . . . . . . . 53 8 History browsing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 9 Handling binary files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 10 Multiple developers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 11 Revision management. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 12 Keyword substitution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 13 Tracking third-party sources . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 14 How your build system interacts with CVS . . . . . . . . . . . . . . . 83 15 Special Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 A Guide to CVS commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 B Quick reference to CVS commands . . . . . . . . . . . . . . . . . . . . . 127 C Reference manual for Administrative files . . . . . . . . . . . . . . . . 139 D All environment variables which affect CVS . . . . . . . . . . . . . . 155 E Compatibility between CVS Versions . . . . . . . . . . . . . . . . . . . 159 F Troubleshooting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 G Credits. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 H Dealing with bugs in CVS or this manual . . . . . . . . . . . . . . . . 171 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 ii CVS—Concurrent Versions System v1.11.23

Table of Contents

1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 What is CVS? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 What is CVS not? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.3 A sample session . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

1.3.1 Getting the source. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.3.2 Committing your changes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.3.3 Cleaning up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.3.4 Viewing differences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

2 The Repository . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2.1 Telling CVS where your repository is . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.2 How data is stored in the repository . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

2.2.1 Where files are stored within the repository . . . . . . . . . . . . . . . . 8 2.2.2 File permissions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.2.3 File Permission issues specific to Windows . . . . . . . . . . . . . . . . 10 2.2.4 The attic. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.2.5 The CVS directory in the repository . . . . . . . . . . . . . . . . . . . . . . 11 2.2.6 CVS locks in the repository . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.2.7 How files are stored in the CVSROOT directory . . . . . . . . . . 13

2.3 How data is stored in the working directory . . . . . . . . . . . . . . . . . . . 13 2.4 The administrative files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

2.4.1 Editing administrative files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.5 Multiple repositories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.6 Creating a repository. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 2.7 Backing up a repository . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 2.8 Moving a repository. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 2.9 Remote repositories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

2.9.1 Server requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 2.9.2 Connecting with rsh or ssh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 2.9.3 Direct connection with password authentication. . . . . . . . . . . 21

2.9.3.1 Setting up the server for password authentication …. 21 2.9.3.2 Using the client with password authentication. . . . . . . . 24 2.9.3.3 Security considerations with password authentication

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 2.9.4 Direct connection with GSSAPI . . . . . . . . . . . . . . . . . . . . . . . . . . 26 2.9.5 Direct connection with Kerberos. . . . . . . . . . . . . . . . . . . . . . . . . . 26 2.9.6 Connecting with fork . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

2.10 Read-only repository access . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 2.11 Temporary directories for the server . . . . . . . . . . . . . . . . . . . . . . . . . . 28

3 Starting a project with CVS . . . . . . . . . . . . . . . . . 29

3.1 Setting up the files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 3.1.1 Creating a directory tree from a number of files . . . . . . . . . . . 29 3.1.2 Creating Files From Other Version Control Systems. . . . . . . 30 3.1.3 Creating a directory tree from scratch . . . . . . . . . . . . . . . . . . . . 30

3.2 Defining the module. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

4 Revisions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

4.1 Revision numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 4.2 Versions, revisions and releases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 4.3 Assigning revisions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 4.4 Tags–Symbolic revisions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 4.5 Specifying what to tag from the working directory . . . . . . . . . . . . . 36 4.6 Specifying what to tag by date or revision . . . . . . . . . . . . . . . . . . . . . 36 4.7 Deleting, moving, and renaming tags . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.8 Tagging and adding and removing files . . . . . . . . . . . . . . . . . . . . . . . . 38 4.9 Sticky tags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

5 Branching and merging . . . . . . . . . . . . . . . . . . . . . . . 41

5.1 What branches are good for . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 5.2 Creating a branch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 5.3 Accessing branches . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 5.4 Branches and revisions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 5.5 Magic branch numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 5.6 Merging an entire branch. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 5.7 Merging from a branch several times . . . . . . . . . . . . . . . . . . . . . . . . . . 46 5.8 Merging differences between any two revisions . . . . . . . . . . . . . . . . . 46 5.9 Merging can add or remove files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 5.10 Merging and keywords . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

6 Recursive behavior . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

7 Adding, removing, and renaming files and

directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

7.1 Adding files to a directory. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 7.2 Removing files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 7.3 Removing directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 7.4 Moving and renaming files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

7.4.1 The Normal way to Rename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 7.4.2 Moving the history file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 7.4.3 Copying the history file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

7.5 Moving and renaming directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

8 History browsing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

8.1 Log messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 8.2 The history database. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 8.3 User-defined logging. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

9 Handling binary files . . . . . . . . . . . . . . . . . . . . . . . . . . 61

9.1 The issues with binary files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 9.2 How to store binary files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

10 Multiple developers . . . . . . . . . . . . . . . . . . . . . . . . . . 63

10.1 File status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 10.2 Bringing a file up to date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 10.3 Conflicts example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 10.4 Informing others about commits. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 10.5 Several developers simultaneously attempting to run CVS …. 67 10.6 Mechanisms to track who is editing files . . . . . . . . . . . . . . . . . . . . . . 68

10.6.1 Telling CVS to watch certain files . . . . . . . . . . . . . . . . . . . . . . . 68 10.6.2 Telling CVS to notify you . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 10.6.3 How to edit a file which is being watched . . . . . . . . . . . . . . . . 70 10.6.4 Information about who is watching and editing . . . . . . . . . . 71 10.6.5 Using watches with old versions of CVS . . . . . . . . . . . . . . . . . 71

10.7 Choosing between reserved or unreserved checkouts . . . . . . . . . . . 71

11 Revision management. . . . . . . . . . . . . . . . . . . . . . . . 73

11.1 When to commit?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

12 Keyword substitution . . . . . . . . . . . . . . . . . . . . . . . . 75

12.1 Keyword List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 12.2 Using keywords . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 12.3 Avoiding substitution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 12.4 Substitution modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 12.5 Problems with the $Log$ keyword. . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

13 Tracking third-party sources . . . . . . . . . . . . . . . . 79

13.1 Importing for the first time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 13.2 Updating with the import command . . . . . . . . . . . . . . . . . . . . . . . . . 79 13.3 Reverting to the latest vendor release . . . . . . . . . . . . . . . . . . . . . . . . 80 13.4 How to handle binary files with cvs import . . . . . . . . . . . . . . . . . . . 80 13.5 How to handle keyword substitution with cvs import . . . . . . . . . 80 13.6 Multiple vendor branches . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

14 How your build system interacts with CVS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

15 Special Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 Appendix A Guide to CVS commands . . . . . . . . 87

A.1 Overall structure of CVS commands . . . . . . . . . . . . . . . . . . . . . . . . . . 87 A.2 CVS’s exit status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 A.3 Default options and the ~/.cvsrc file . . . . . . . . . . . . . . . . . . . . . . . . . . 88 A.4 Global options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 A.5 Common command options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 A.6 add—Add files and directories to the repository . . . . . . . . . . . . . . . 93

A.6.1 add options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 A.6.2 add examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

A.6.2.1 Adding a directory. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 A.6.2.2 Adding a file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 A.6.2.3 Undoing a remove command . . . . . . . . . . . . . . . . . . . . . . . . 93

A.7 admin—Administration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

A.7.1 admin options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 A.8 annotate—What revision modified each line of a file?. . . . . . . . . . 98

A.8.1 annotate options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 A.8.2 annotate example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98

A.9 checkout—Check out sources for editing. . . . . . . . . . . . . . . . . . . . . . . 99 A.9.1 checkout options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 A.9.2 checkout examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

A.10 commit—Check files into the repository. . . . . . . . . . . . . . . . . . . . . 101 A.10.1 commit options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 A.10.2 commit examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

A.10.2.1 Committing to a branch . . . . . . . . . . . . . . . . . . . . . . . . . . 102 A.10.2.2 Creating the branch after editing . . . . . . . . . . . . . . . . . 103

A.11 diff—Show differences between revisions . . . . . . . . . . . . . . . . . . . . 103

A.11.1 diff options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 A.11.1.1 Line group formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 A.11.1.2 Line formats. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110

A.11.2 diff examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 A.12 export—Export sources from CVS, similar to checkout . . . . . . 112

A.12.1 export options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 A.13 history—Show status of files and users. . . . . . . . . . . . . . . . . . . . . . 113

A.13.1 history options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 A.14 import—Import sources into CVS, using vendor branches … 115

A.14.1 import options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 A.14.2 import output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 A.14.3 import examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116

A.15 log—Print out log information for files . . . . . . . . . . . . . . . . . . . . . . 116 A.15.1 log options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 A.15.2 log examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118

A.16 rdiff—’patch’ format diffs between releases . . . . . . . . . . . . . . . . . . 118 A.16.1 rdiff options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 A.16.2 rdiff examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

A.17 release—Indicate that a Module is no longer in use . . . . . . . . . 120 A.17.1 release options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 A.17.2 release output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 A.17.3 release examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 A.18 remove—Remove files from active use. . . . . . . . . . . . . . . . . . . . . . . 121

A.18.1 remove options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 A.18.2 remove examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122

A.18.2.1 Removing a file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 A.18.2.2 Removing entire directories . . . . . . . . . . . . . . . . . . . . . . . 122

A.19 update—Bring work tree in sync with repository. . . . . . . . . . . . 122 A.19.1 update options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 A.19.2 update output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124

Appendix B Quick reference to CVS commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127

Appendix C Reference manual for

Administrative files . . . . . . . . . . . . . . . . . . . . . . . . . 139

C.1 The modules file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 C.1.1 Alias modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 C.1.2 Regular modules. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 C.1.3 Ampersand modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 C.1.4 Excluding directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 C.1.5 Module options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 C.1.6 How the modules file “program options” programs are run

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 C.2 The cvswrappers file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 C.3 The Trigger Scripts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142

C.3.1 The common syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 C.3.2 Security and the Trigger Scripts . . . . . . . . . . . . . . . . . . . . . . . . 143

C.4 The commit support files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 C.4.1 Commitinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 C.4.2 Verifying log messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 C.4.3 Editinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146

C.4.3.1 Editinfo example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 C.4.4 Loginfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148

C.4.4.1 Loginfo example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 C.4.4.2 Keeping a checked out copy . . . . . . . . . . . . . . . . . . . . . . . . 149

C.5 Rcsinfo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 C.6 Taginfo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 C.7 Ignoring files via cvsignore. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 C.8 The checkoutlist file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 C.9 The history file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 C.10 Expansions in administrative files. . . . . . . . . . . . . . . . . . . . . . . . . . . 152 C.11 The CVSROOT/config configuration file . . . . . . . . . . . . . . . . . . . . 153

Appendix D All environment variables which affect CVS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155

viii CVS—Concurrent Versions System v1.11.23

Appendix E Compatibility between CVS

Versions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 Appendix F Troubleshooting . . . . . . . . . . . . . . . . . 161

F.1 Partial list of error messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 F.2 Trouble making a connection to a CVS server . . . . . . . . . . . . . . . . 167 F.3 Other common problems. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168

Appendix G Credits . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 Appendix H Dealing with bugs in CVS or this

manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173

1 Overview

This chapter is for people who have never used cvs, and perhaps have never used version control software before.

If you are already familiar with cvs and are just trying to learn a particular feature or remember a certain command, you can probably skip everything here.

1.1 What is CVS?

cvs is a version control system. Using it, you can record the history of your source files.

For example, bugs sometimes creep in when software is modified, and you might not detect the bug until a long time after you make the modification. With cvs, you can easily retrieve old versions to see exactly which change caused the bug. This can sometimes be a big help.

You could of course save every version of every file you have ever created. This would however waste an enormous amount of disk space. cvs stores all the versions of a file in a single file in a clever way that only stores the differences between versions.

cvs also helps you if you are part of a group of people working on the same project. It is all too easy to overwrite each others’ changes unless you are extremely careful. Some editors, like gnu Emacs, try to make sure that two people never modify the same file at the same time. Unfortunately, if someone is using another editor, that safeguard will not work. cvs solves this problem by insulating the different developers from each other. Every developer works in his own directory, and cvs merges the work when each developer is done.

cvs started out as a bunch of shell scripts written by Dick Grune, posted to the newsgroup comp.sources.unix in the volume 6 release of July, 1986. While no actual code from these shell scripts is present in the current version of cvs much of the cvs conflict resolution algorithms come from them.

In April, 1989, Brian Berliner designed and coded cvs. Jeff Polk later helped Brian with the design of the cvs module and vendor branch support.

You can get cvs in a variety of ways, including free download from the Internet. For more information on downloading cvs and other cvs topics, see:

     http://cvs.nongnu.org/

There is a mailing list, known as info-cvs@nongnu.org, devoted to cvs. To subscribe or unsubscribe write to info-cvs-request@nongnu.org. If you prefer a Usenet group, there is a one-way mirror (posts to the email list are usually sent to the news group, but not vice versa) of info-cvs@nongnu.org at news:gnu.cvs.help. The right Usenet group for posts is news:comp.software.config-mgmt which is for cvs discussions (along with other configuration management systems). In the future, it might be possible to create a comp.software.config-mgmt.cvs, but probably only if there is sufficient cvs traffic on

news:comp.software.config-mgmt.

You can also subscribe to the bug-cvs@nongnu.org mailing list, described in more detail in Appendix H [BUGS], page 171. To subscribe send mail to bug-cvs-request@nongnu.org. There is a two-way Usenet mirror (posts to the Usenet group are usually sent to the email list and vice versa) of bug-cvs@nongnu.org named

news:gnu.cvs.bug.

1.2 What is CVS not?

cvs can do a lot of things for you, but it does not try to be everything for everyone.

cvs is not a build system.

Though the structure of your repository and modules file interact with your build system (e.g. ‘Makefile’s), they are essentially independent.

cvs does not dictate how you build anything. It merely stores files for retrieval in a tree structure you devise.

cvs does not dictate how to use disk space in the checked out working directories. If you write your ‘Makefile’s or scripts in every directory so they have to know the relative positions of everything else, you wind up requiring the entire repository to be checked out.

If you modularize your work, and construct a build system that will share files (via links, mounts, VPATH in ‘Makefile’s, etc.), you can arrange your disk usage however you like.

But you have to remember that any such system is a lot of work to construct and maintain. cvs does not address the issues involved.

Of course, you should place the tools created to support such a build system (scripts, ‘Makefile’s, etc.) under cvs.

Figuring out what files need to be rebuilt when something changes is, again, something to be handled outside the scope of cvs. One traditional approach is to use make for building, and use some automated tool for generating the dependencies which make uses.

See Chapter 14 [Builds], page 83, for more information on doing builds in conjunction with cvs.

cvs is not a substitute for management.

Your managers and project leaders are expected to talk to you frequently enough to make certain you are aware of schedules, merge points, branch names and release dates. If they don’t, cvs can’t help.

cvs is an instrument for making sources dance to your tune. But you are the piper and the composer. No instrument plays itself or writes its own music.

cvs is not a substitute for developer communication.

When faced with conflicts within a single file, most developers manage to resolve them without too much effort. But a more general definition of “conflict” includes problems too difficult to solve without communication between developers.

cvs cannot determine when simultaneous changes within a single file, or across a whole collection of files, will logically conflict with one another. Its concept of a conflict is purely textual, arising when two changes to the same base file are near enough to spook the merge (i.e., diff3) command.

cvs does not claim to help at all in figuring out non-textual or distributed conflicts in program logic.

For example: Say you change the arguments to function X defined in file ‘A’. At the same time, someone edits file ‘B’, adding new calls to function X using the old arguments. You are outside the realm of cvs’s competence. Acquire the habit of reading specs and talking to your peers.

cvs does not have change control Change control refers to a number of things. First of all it can mean bugtracking, that is being able to keep a database of reported bugs and the status of each one (Is it fixed? In what release? Has the bug submitter agreed that it is fixed?). For interfacing cvs to an external bug-tracking system, see the ‘rcsinfo’ and ‘verifymsg’ files (see Appendix C [Administrative files], page 139). Another aspect of change control is keeping track of the fact that changes to several files were in fact changed together as one logical change. If you check in several files in a single cvs commit operation, cvs then forgets that those files were checked in together, and the fact that they have the same log message is the only thing tying them together. Keeping a gnu style ‘ChangeLog’ can help somewhat. Another aspect of change control, in some systems, is the ability to keep track of the status of each change. Some changes have been written by a developer, others have been reviewed by a second developer, and so on. Generally, the way to do this with cvs is to generate a diff (using cvs diff or diff) and email it to someone who can then apply it using the patch utility. This is very flexible, but depends on mechanisms outside cvs to make sure nothing falls through the cracks.

cvs is not an automated testing program It should be possible to enforce mandatory use of a test suite using the commitinfo file. I haven’t heard a lot about projects trying to do that or whether there are subtle gotchas, however.

cvs does not have a built-in process model Some systems provide ways to ensure that changes or releases go through various steps, with various approvals as needed. Generally, one can accomplish this with cvs but it might be a little more work. In some cases you’ll want to

use the ‘commitinfo’, ‘loginfo’, ‘rcsinfo’, or ‘verifymsg’ files, to require

that certain steps be performed before cvs will allow a checkin. Also consider whether features such as branches and tags can be used to perform tasks such as doing work in a development tree and then merging certain changes over to a stable tree only once they have been proven.

1.3 A sample session

As a way of introducing cvs, we’ll go through a typical work-session using cvs. The first thing to understand is that cvs stores all files in a centralized repository (see Chapter 2 [Repository], page 7); this section assumes that a repository is set up.

Suppose you are working on a simple compiler. The source consists of a handful of C files and a ‘Makefile’. The compiler is called ‘tc’ (Trivial Compiler), and the repository is set up so that there is a module called ‘tc’.

1.3.1 Getting the source

The first thing you must do is to get your own working copy of the source for ‘tc’. For this, you use the checkout command:

     $ cvs checkout tc

This will create a new directory called ‘tc’ and populate it with the source files.

     $ cd tc
     $ ls
     CVS         Makefile    backend.c   driver.c    frontend.c  parser.c

The ‘CVS’ directory is used internally by cvs. Normally, you should not modify or remove any of the files in it.

You start your favorite editor, hack away at ‘backend.c’, and a couple of hours later you have added an optimization pass to the compiler. A note to rcs and sccs users: There is no need to lock the files that you want to edit. See Chapter 10 [Multiple developers], page 63, for an explanation.

1.3.2 Committing your changes

When you have checked that the compiler is still compilable you decide to make a new version of ‘backend.c’. This will store your new ‘backend.c’ in the repository and make it available to anyone else who is using that same repository.

     $ cvs commit backend.c

cvs starts an editor, to allow you to enter a log message. You type in “Added an optimization pass.”, save the temporary file, and exit the editor.

The environment variable $CVSEDITOR determines which editor is started. If $CVSEDITOR is not set, then if the environment variable $EDITOR is set, it will be used. If both $CVSEDITOR and $EDITOR are not set then there is a default which will vary with your operating system, for example vi for unix or notepad for Windows NT/95.

In addition, cvs checks the $VISUAL environment variable. Opinions vary on whether this behavior is desirable and whether future releases of cvs should check $VISUAL or ignore it. You will be OK either way if you make sure that $VISUAL is either unset or set to the same thing as $EDITOR.

When cvs starts the editor, it includes a list of files which are modified. For the cvs client, this list is based on comparing the modification time of the file against the modification time that the file had when it was last gotten or updated. Therefore, if a file’s modification time has changed but its contents have not, it will show up as modified. The simplest way to handle this is simply not to worry about it—if you proceed with the commit cvs will detect that the contents are not modified and treat it as an unmodified file. The next update will clue cvs in to the fact that the file is unmodified, and it will reset its stored timestamp so that the file will not show up in future editor sessions.

If you want to avoid starting an editor you can specify the log message on the command line using the ‘-m’ flag instead, like this:

     $ cvs commit -m "Added an optimization pass" backend.c

1.3.3 Cleaning up

Before you turn to other tasks you decide to remove your working copy of tc. One acceptable way to do that is of course

$ cd ..
$ rm -r tc

but a better way is to use the release command (see Section A.17 [release], page 120):

$ cd ..
$ cvs release -d tc
M driver.c
? tc
You have [1] altered files in this repository.
Are you sure you want to release (and delete) directory `tc': n
** `release' aborted by user choice.

The release command checks that all your modifications have been committed. If history logging is enabled it also makes a note in the history file. See Section C.9 [history file], page 152.

When you use the ‘-d’ flag with release, it also removes your working copy. In the example above, the release command wrote a couple of lines of output. ‘? tc’

means that the file ‘tc’ is unknown to cvs. That is nothing to worry about: ‘tc’ is the executable compiler, and it should not be stored in the repository. See Section C.7 [cvsignore], page 150, for information about how to make that warning go away. See Section A.17.2 [release output], page 120, for a complete explanation of all possible output from release.

‘M driver.c’ is more serious. It means that the file ‘driver.c’ has been modified since it was checked out.

The release command always finishes by telling you how many modified files you have in your working copy of the sources, and then asks you for confirmation before deleting any files or making any note in the history file.

You decide to play it safe and answer n ^{⟨}\underline{\overline{RET}}^{⟩} when release asks for confirmation.

1.3.4 Viewing differences

You do not remember modifying ‘driver.c’, so you want to see what has happened to that file.

$ cd tc
$ cvs diff driver.c

This command runs diff to compare the version of ‘driver.c’ that you checked out with your working copy. When you see the output you remember that you added a command line option that enabled the optimization pass. You check it in, and release the module.

$ cvs commit -m "Added an optimization pass" driver.c
Checking in driver.c;
/usr/local/cvsroot/tc/driver.c,v  <--  driver.c
new revision: 1.2; previous revision: 1.1
done
$ cd ..
$ cvs release -d tc

? tc
You have [0] altered files in this repository.
Are you sure you want to release (and delete) directory `tc': y

2 The Repository

The cvs repository stores a complete copy of all the files and directories which are under version control.

Normally, you never access any of the files in the repository directly. Instead, you use cvs commands to get your own copy of the files into a working directory, and then work on that copy. When you’ve finished a set of changes, you check (or commit) them back into the repository. The repository then contains the changes which you have made, as well as recording exactly what you changed, when you changed it, and other such information. Note that the repository is not a subdirectory of the working directory, or vice versa; they should be in separate locations.

cvs can access a repository by a variety of means. It might be on the local computer, or it might be on a computer across the room or across the world. To distinguish various ways to access a repository, the repository name can start with an access method. For example, the access method :local: means to access a repository directory, so the reposi-

tory :local:/usr/local/cvsroot means that the repository is in ‘/usr/local/cvsroot’

on the computer running cvs. For information on other access methods, see Section 2.9 [Remote repositories], page 19.

If the access method is omitted, then if the repository starts with ‘/’, then :local: is assumed. If it does not start with ‘/’ then either :ext: or :server: is assumed. For example, if you have a local repository in ‘/usr/local/cvsroot’, you can use /usr/local/cvsroot instead of :local:/usr/local/cvsroot. But if (under Windows NT, for example) your local repository is ‘c:\src\cvsroot’, then you must specify the access method, as in

:local:c:/src/cvsroot.

The repository is split in two parts. ‘$CVSROOT/CVSROOT’ contains administrative files for cvs. The other directories contain the actual user-defined modules.

2.1 Telling CVS where your repository is

There are several ways to tell cvs where to find the repository. You can name the repository on the command line explicitly, with the -d (for "directory") option:

     cvs -d /usr/local/cvsroot checkout yoyodyne/tc

Or you can set the $CVSROOT environment variable to an absolute path to the root of the repository, ‘/usr/local/cvsroot’ in this example. To set $CVSROOT, csh and tcsh users should have this line in their ‘.cshrc’ or ‘.tcshrc’ files:

     setenv CVSROOT /usr/local/cvsroot

sh and bash users should instead have these lines in their ‘.profile’ or ‘.bashrc’:

     CVSROOT=/usr/local/cvsroot
     export CVSROOT

A repository specified with -d will override the $CVSROOT environment variable. Once you’ve checked a working copy out from the repository, it will remember where its repository is (the information is recorded in the ‘CVS/Root’ file in the working copy).

The -d option and the ‘CVS/Root’ file both override the $CVSROOT environment variable. If -d option differs from ‘CVS/Root’, the former is used. Of course, for proper operation they should be two ways of referring to the same repository.

2.2 How data is stored in the repository

For most purposes it isn’t important how cvs stores information in the repository. In fact, the format has changed in the past, and is likely to change in the future. Since in almost all cases one accesses the repository via cvs commands, such changes need not be disruptive.

However, in some cases it may be necessary to understand how cvs stores data in the repository, for example you might need to track down cvs locks (see Section 10.5 [Concurrency], page 67) or you might need to deal with the file permissions appropriate for the repository.

2.2.1 Where files are stored within the repository

The overall structure of the repository is a directory tree corresponding to the directories in the working directory. For example, supposing the repository is in

/usr/local/cvsroot

here is a possible directory tree (showing only the directories):

/usr
|
+--local
|   |
|   +--cvsroot
|   |    |
|   |    +--CVSROOT
|      (administrative files)
|
+--gnu
|   |
|   +--diff
|   |   (source code to gnu diff)
|   |
|   +--rcs
|   |   (source code to rcs)
|   |
|   +--cvs
|       (source code to cvs)
|
+--yoyodyne
    |
    +--tc
    |    |
    |    +--man
    |    |
    |    +--testing
    |
    +--(other Yoyodyne software)

With the directories are history files for each file under version control. The name of the history file is the name of the corresponding file with ‘,v’ appended to the end. Here is what the repository for the ‘yoyodyne/tc’ directory might look like:

       $CVSROOT
         |
         +--yoyodyne
         |   |
         |   +--tc
         |   |   |
                 +--Makefile,v
                 +--backend.c,v
                 +--driver.c,v
                 +--frontend.c,v
                 +--parser.c,v
                 +--man
                 |    |
                 |    +--tc.1,v
                 |
                 +--testing
                      |
                      +--testpgm.t,v
                      +--test2.t,v

The history files contain, among other things, enough information to recreate any revision of the file, a log of all commit messages and the user-name of the person who committed the revision. The history files are known as RCS files, because the first program to store files in that format was a version control system known as rcs. For a full description of the file format, see the man page rcsfile(5), distributed with rcs, or the file ‘doc/RCSFILES’ in the cvs source distribution. This file format has become very common—many systems other than cvs or rcs can at least import history files in this format.

The rcs files used in cvs differ in a few ways from the standard format. The biggest difference is magic branches; for more information see Section 5.5 [Magic branch numbers], page 44. Also in cvs the valid tag names are a subset of what rcs accepts; for cvs’s rules see Section 4.4 [Tags], page 34.

2.2.2 File permissions

All ‘,v’ files are created read-only, and you should not change the permission of those files. The directories inside the repository should be writable by the persons that have permission to modify the files in each directory. This normally means that you must create a UNIX group (see group(5)) consisting of the persons that are to edit the files in a project, and set up the repository so that it is that group that owns the directory. (On some systems, you also need to set the set-group-ID-on-execution bit on the repository directories (see chmod(1)) so that newly-created files and directories get the group-ID of the parent directory rather than that of the current process.)

This means that you can only control access to files on a per-directory basis.

Note that users must also have write access to check out files, because cvs needs to create lock files (see Section 10.5 [Concurrency], page 67). You can use LockDir in CVS-ROOT/config to put the lock files somewhere other than in the repository if you want to allow read-only access to some directories (see Section C.11 [config], page 153).

Also note that users must have write access to the ‘CVSROOT/val-tags’ file. cvs uses it to keep track of what tags are valid tag names (it is sometimes updated when tags are used, as well as when they are created).

Each rcs file will be owned by the user who last checked it in. This has little significance; what really matters is who owns the directories.

cvs tries to set up reasonable file permissions for new directories that are added inside the tree, but you must fix the permissions manually when a new directory should have different permissions than its parent directory. If you set the CVSUMASK environment variable, that will control the file permissions which cvs uses in creating directories and/or files in the repository. CVSUMASK does not affect the file permissions in the working directory; such files have the permissions which are typical for newly created files, except that sometimes cvs creates them read-only (see the sections on watches, Section 10.6.1 [Setting a watch], page 68; -r, Section A.4 [Global options], page 88; or CVSREAD, Appendix D [Environment variables], page 155).

Note that using the client/server cvs (see Section 2.9 [Remote repositories], page 19), there is no good way to set CVSUMASK; the setting on the client machine has no effect. If you are connecting with rsh, you can set CVSUMASK in ‘.bashrc’ or ‘.cshrc’, as described in the documentation for your operating system. This behavior might change in future versions of cvs; do not rely on the setting of CVSUMASK on the client having no effect.

Using pserver, you will generally need stricter permissions on the cvsroot directory and directories above it in the tree; see Section 2.9.3.3 [Password authentication security], page 25.

Some operating systems have features which allow a particular program to run with the ability to perform operations which the caller of the program could not. For example, the set user ID (setuid) or set group ID (setgid) features of unix or the installed image feature of VMS. cvs was not written to use such features and therefore attempting to install cvs in this fashion will provide protection against only accidental lapses; anyone who is trying to circumvent the measure will be able to do so, and depending on how you have set it up may gain access to more than just cvs. You may wish to instead consider pserver. It shares some of the same attributes, in terms of possibly providing a false sense of security or opening security holes wider than the ones you are trying to fix, so read the documentation on pserver security carefully if you are considering this option (Section 2.9.3.3 [Password authentication security], page 25).

2.2.3 File Permission issues specific to Windows

Some file permission issues are specific to Windows operating systems (Windows 95, Windows NT, and presumably future operating systems in this family. Some of the following might apply to OS/2 but I’m not sure).

If you are using local cvs and the repository is on a networked file system which is served by the Samba SMB server, some people have reported problems with permissions. Enabling WRITE=YES in the samba configuration is said to fix/workaround it. Disclaimer: I haven’t investigated enough to know the implications of enabling that option, nor do I know whether there is something which cvs could be doing differently in order to avoid the problem. If you find something out, please let us know as described in Appendix H [BUGS], page 171.

2.2.4 The attic

You will notice that sometimes cvs stores an rcs file in the Attic. For example, if the cvsroot is ‘/usr/local/cvsroot’ and we are talking about the file ‘backend.c’ in the directory ‘yoyodyne/tc’, then the file normally would be in

     /usr/local/cvsroot/yoyodyne/tc/backend.c,v

but if it goes in the attic, it would be in

     /usr/local/cvsroot/yoyodyne/tc/Attic/backend.c,v

instead. It should not matter from a user point of view whether a file is in the attic; cvs keeps track of this and looks in the attic when it needs to. But in case you want to know, the rule is that the RCS file is stored in the attic if and only if the head revision on the trunk has state dead. A dead state means that file has been removed, or never added, for that revision. For example, if you add a file on a branch, it will have a trunk revision in dead state, and a branch revision in a non-dead state.

2.2.5 The CVS directory in the repository

The ‘CVS’ directory in each repository directory contains information such as file attributes (in a file called ‘CVS/fileattr’. In the future additional files may be added to this directory, so implementations should silently ignore additional files.

This behavior is implemented only by cvs 1.7 and later; for details see Section 10.6.5 [Watches Compatibility], page 71.

The format of the ‘fileattr’ file is a series of entries of the following form (where ‘{’ and ‘}’ means the text between the braces can be repeated zero or more times):

ent-type filename <tab> attrname = attrval {; attrname = attrval} <linefeed>

ent-type is ‘F’ for a file, in which case the entry specifies the attributes for that file.

ent-type is ‘D’, and filename empty, to specify default attributes to be used for newly added files.

Other ent-type are reserved for future expansion. cvs 1.9 and older will delete them any time it writes file attributes. cvs 1.10 and later will preserve them.

Note that the order of the lines is not significant; a program writing the fileattr file may rearrange them at its convenience.

There is currently no way of quoting tabs or line feeds in the filename, ‘=’ in attrname, ‘;’ in attrval, etc. Note: some implementations also don’t handle a NUL character in any of the fields, but implementations are encouraged to allow it.

By convention, attrname starting with ‘_’ is for an attribute given special meaning by cvs; other attrnames are for user-defined attributes (or will be, once implementations start supporting user-defined attributes).

Built-in attributes:

_watched Present means the file is watched and should be checked out read-only.

_watchers

Users with watches for this file. Value is watcher > type { , watcher > type } where watcher is a username, and type is zero or more of edit,unedit,commit separated by ‘+’ (that is, nothing if none; there is no "none" or "all" keyword).

_editors Users editing this file. Value is editor > val { , editor > val } where editor is a username, and val is time+hostname+pathname, where time is when the cvs edit command (or equivalent) happened, and hostname and pathname are for the working directory.

Example:

     Ffile1 _watched=;_watchers=joe>edit,mary>commit
     Ffile2 _watched=;_editors=sue>8 Jan 1975+workstn1+/home/sue/cvs
     D _watched=

means that the file ‘file1’ should be checked out read-only. Furthermore, joe is watching for edits and mary is watching for commits. The file ‘file2’ should be checked out readonly; sue started editing it on 8 Jan 1975 in the directory ‘/home/sue/cvs’ on the machine workstn1. Future files which are added should be checked out read-only. To represent this example here, we have shown a space after ‘D’, ‘Ffile1’, and ‘Ffile2’, but in fact there must be a single tab character there and no spaces.

2.2.6 CVS locks in the repository

For an introduction to cvs locks focusing on user-visible behavior, see Section 10.5 [Concurrency], page 67. The following section is aimed at people who are writing tools which want to access a cvs repository without interfering with other tools accessing the same repository. If you find yourself confused by concepts described here, like read lock, write lock, and deadlock, you might consult the literature on operating systems or databases.

Any file in the repository with a name starting with ‘#cvs.rfl.’ is a read lock. Any file in the repository with a name starting with ‘#cvs.wfl’ is a write lock. Old versions of cvs (before cvs 1.5) also created files with names starting with ‘#cvs.tfl’, but they are not discussed here. The directory ‘#cvs.lock’ serves as a master lock. That is, one must obtain this lock first before creating any of the other locks.

To obtain a read lock, first create the ‘#cvs.lock’ directory. This operation must be atomic (which should be true for creating a directory under most operating systems). If it fails because the directory already existed, wait for a while and try again. After obtaining the ‘#cvs.lock’ lock, create a file whose name is ‘#cvs.rfl.’ followed by information of your choice (for example, hostname and process identification number). Then remove the ‘#cvs.lock’ directory to release the master lock. Then proceed with reading the repository. When you are done, remove the ‘#cvs.rfl’ file to release the read lock.

To obtain a write lock, first create the ‘#cvs.lock’ directory, as with read locks. Then check that there are no files whose names start with ‘#cvs.rfl.’. If there are, remove ‘#cvs.lock’, wait for a while, and try again. If there are no readers, then create a file whose name is ‘#cvs.wfl’ followed by information of your choice (for example, hostname and process identification number). Hang on to the ‘#cvs.lock’ lock. Proceed with writing the repository. When you are done, first remove the ‘#cvs.wfl’ file and then the ‘#cvs.lock’ directory. Note that unlike the ‘#cvs.rfl’ file, the ‘#cvs.wfl’ file is just informational; it has no effect on the locking operation beyond what is provided by holding on to the ‘#cvs.lock’ lock itself.

Note that each lock (write lock or read lock) only locks a single directory in the repository, including ‘Attic’ and ‘CVS’ but not including subdirectories which represent other directories under version control. To lock an entire tree, you need to lock each directory (note that if you fail to obtain any lock you need, you must release the whole tree before waiting and trying again, to avoid deadlocks).

Note also that cvs expects write locks to control access to individual ‘foo,v’ files. rcs has a scheme where the ‘,foo,’ file serves as a lock, but cvs does not implement it and so taking out a cvs write lock is recommended. See the comments at rcs internal lockfile in the cvs source code for further discussion/rationale.

2.2.7 How files are stored in the CVSROOT directory

The ‘$CVSROOT/CVSROOT’ directory contains the various administrative files. In some ways this directory is just like any other directory in the repository; it contains rcs files whose names end in ‘,v’, and many of the cvs commands operate on it the same way. However, there are a few differences.

For each administrative file, in addition to the rcs file, there is also a checked out copy of the file. For example, there is an rcs file ‘loginfo,v’ and a file ‘loginfo’ which contains the latest revision contained in ‘loginfo,v’. When you check in an administrative file, cvs should print

     cvs commit: Rebuilding administrative file database

and update the checked out copy in ‘$CVSROOT/CVSROOT’. If it does not, there is something wrong (see Appendix H [BUGS], page 171). To add your own files to the files to be updated in this fashion, you can add them to the ‘checkoutlist’ administrative file (see Section C.8 [checkoutlist], page 151).

By default, the ‘modules’ file behaves as described above. If the modules file is very large, storing it as a flat text file may make looking up modules slow (I’m not sure whether this is as much of a concern now as when cvs first evolved this feature; I haven’t seen benchmarks). Therefore, by making appropriate edits to the cvs source code one can store the modules file in a database which implements the ndbm interface, such as Berkeley db or GDBM. If this option is in use, then the modules database will be stored in the files

‘modules.db’, ‘modules.pag’, and/or ‘modules.dir’.

For information on the meaning of the various administrative files, see Appendix C [Administrative files], page 139.

2.3 How data is stored in the working directory

While we are discussing cvs internals which may become visible from time to time, we might as well talk about what cvs puts in the ‘CVS’ directories in the working directories. As with the repository, cvs handles this information and one can usually access it via cvs commands. But in some cases it may be useful to look at it, and other programs, such as the jCVS graphical user interface or the VC package for emacs, may need to look at it. Such programs should follow the recommendations in this section if they hope to be able to work with other programs which use those files, including future versions of the programs just mentioned and the command-line cvs client.

The ‘CVS’ directory contains several files. Programs which are reading this directory should silently ignore files which are in the directory but which are not documented here, to allow for future expansion.

The files are stored according to the text file convention for the system in question. This means that working directories are not portable between systems with differing conventions for storing text files. This is intentional, on the theory that the files being managed by cvs probably will not be portable between such systems either.

‘Root’ This file contains the current cvs root, as described in Section 2.1 [Specifying a repository], page 7.

‘Repository’

This file contains the directory within the repository which the current directory corresponds with. It can be either an absolute pathname or a relative pathname; cvs has had the ability to read either format since at least version 1.3 or so. The relative pathname is relative to the root, and is the more sensible approach, but the absolute pathname is quite common and implementations should accept either. For example, after the command

     cvs -d :local:/usr/local/cvsroot checkout yoyodyne/tc

‘Root’ will contain

     :local:/usr/local/cvsroot

and ‘Repository’ will contain either

     /usr/local/cvsroot/yoyodyne/tc

or

     yoyodyne/tc

If the particular working directory does not correspond to a directory in the repository, then ‘Repository’ should contain ‘CVSROOT/Emptydir’.

‘Entries’ This file lists the files and directories in the working directory. The first character of each line indicates what sort of line it is. If the character is unrecognized, programs reading the file should silently skip that line, to allow for future expansion. If the first character is ‘/’, then the format is:

     /name/revision/timestamp[+conflict]/options/tagdate

where ‘[’ and ‘]’ are not part of the entry, but instead indicate that the ‘+’ and conflict marker are optional. name is the name of the file within the directory. revision is the revision that the file in the working derives from, or ‘0’ for an added file, or ‘-’ followed by a revision for a removed file. timestamp is the timestamp of the file at the time that cvs created it; if the timestamp differs with the actual modification time of the file it means the file has been modified. It is stored in the format used by the ISO C asctime() function (for example, ‘Sun Apr 7 01:29:26 1996’). One may write a string which is not in that format, for example, ‘Result of merge’, to indicate that the file should always be considered to be modified. This is not a special case; to see whether a file is modified a program should take the timestamp of the file and simply do a string compare with timestamp. If there was a conflict, conflict can be set to the modification time of the file after the file has been written with conflict markers (see Section 10.3 [Conflicts example], page 65). Thus if conflict is subsequently the same as the actual modification time of the file it means that the user has obviously not resolved the conflict. options contains sticky options (for example ‘-kb’ for a binary file). tagdate contains ‘T’ followed by a tag name, or ‘D’ for a date, followed by a sticky tag or date. Note that if timestamp contains a pair of timestamps separated by a space, rather than a single timestamp, you are dealing with a version of cvs earlier than cvs 1.5 (not documented here).

The timezone on the timestamp in CVS/Entries (local or universal) should be the same as the operating system stores for the timestamp of the file itself. For example, on Unix the file’s timestamp is in universal time (UT), so the timestamp in CVS/Entries should be too. On vms, the file’s timestamp is in local time, so cvs on vms should use local time. This rule is so that files do not appear to be modified merely because the timezone changed (for example, to or from summer time).

If the first character of a line in ‘Entries’ is ‘D’, then it indicates a subdirectory. ‘D’ on a line all by itself indicates that the program which wrote the ‘Entries’ file does record subdirectories (therefore, if there is such a line and no other lines beginning with ‘D’, one knows there are no subdirectories). Otherwise, the line looks like:

     D/name/filler1/filler2/filler3/filler4

where name is the name of the subdirectory, and all the filler fields should be silently ignored, for future expansion. Programs which modify Entries files should preserve these fields.

The lines in the ‘Entries’ file can be in any order.

‘Entries.Log’

This file does not record any information beyond that in ‘Entries’, but it does provide a way to update the information without having to rewrite the entire ‘Entries’ file, including the ability to preserve the information even if the program writing ‘Entries’ and ‘Entries.Log’ abruptly aborts. Programs which are reading the ‘Entries’ file should also check for ‘Entries.Log’. If the latter exists, they should read ‘Entries’ and then apply the changes mentioned in ‘Entries.Log’. After applying the changes, the recommended practice is to rewrite ‘Entries’ and then delete ‘Entries.Log’. The format of a line in ‘Entries.Log’ is a single character command followed by a space followed by a line in the format specified for a line in ‘Entries’. The single character command is ‘A’ to indicate that the entry is being added, ‘R’ to indicate that the entry is being removed, or any other character to indicate that the entire line in ‘Entries.Log’ should be silently ignored (for future expansion). If the second character of the line in ‘Entries.Log’ is not a space, then it was written by an older version of cvs (not documented here).

Programs which are writing rather than reading can safely ignore ‘Entries.Log’ if they so choose.

‘Entries.Backup’

This is a temporary file. Recommended usage is to write a new entries file to ‘Entries.Backup’, and then to rename it (atomically, where possible) to

‘Entries’.
‘Entries.Static’

The only relevant thing about this file is whether it exists or not. If it exists, then it means that only part of a directory was gotten and cvs will not create additional files in that directory. To clear it, use the update command with the ‘-d’ option, which will get the additional files and remove ‘Entries.Static’.

‘Tag’ This file contains per-directory sticky tags or dates. The first character is ‘T’ for a branch tag, ‘N’ for a non-branch tag, or ‘D’ for a date, or another character to mean the file should be silently ignored, for future expansion. This character is followed by the tag or date. Note that per-directory sticky tags or dates are used for things like applying to files which are newly added; they might not be the same as the sticky tags or dates on individual files. For general information on sticky tags and dates, see Section 4.9 [Sticky tags], page 38.

‘Notify’ This file stores notifications (for example, for edit or unedit) which have not yet been sent to the server. Its format is not yet documented here.

‘Notify.tmp’

This file is to ‘Notify’ as ‘Entries.Backup’ is to ‘Entries’. That is, to write ‘Notify’, first write the new contents to ‘Notify.tmp’ and then (atomically where possible), rename it to ‘Notify’.

‘Base’ If watches are in use, then an edit command stores the original copy of the file in the ‘Base’ directory. This allows the unedit command to operate even if it is unable to communicate with the server.

‘Baserev’ The file lists the revision for each of the files in the ‘Base’ directory. The format

is:

     Bname/rev/expansion

where expansion should be ignored, to allow for future expansion.

‘Baserev.tmp’

This file is to ‘Baserev’ as ‘Entries.Backup’ is to ‘Entries’. That is, to write ‘Baserev’, first write the new contents to ‘Baserev.tmp’ and then (atomically where possible), rename it to ‘Baserev’.

‘Template’

This file contains the template specified by the ‘rcsinfo’ file (see Section C.5 [rcsinfo], page 149). It is only used by the client; the non-client/server cvs consults ‘rcsinfo’ directly.

2.4 The administrative files

The directory ‘$CVSROOT/CVSROOT’ contains some administrative files. See Appendix C [Administrative files], page 139, for a complete description. You can use cvs without any of these files, but some commands work better when at least the ‘modules’ file is properly set up.

The most important of these files is the ‘modules’ file. It defines all modules in the repository. This is a sample ‘modules’ file.

     CVSROOT         CVSROOT
     modules         CVSROOT modules
     cvs             gnu/cvs
     rcs             gnu/rcs
     diff            gnu/diff
     tc              yoyodyne/tc

The ‘modules’ file is line oriented. In its simplest form each line contains the name of the module, whitespace, and the directory where the module resides. The directory is a path relative to $CVSROOT. The last four lines in the example above are examples of such lines.

The line that defines the module called ‘modules’ uses features that are not explained here. See Section C.1 [modules], page 139, for a full explanation of all the available features.

2.4.1 Editing administrative files

You edit the administrative files in the same way that you would edit any other module. Use ‘cvs checkout CVSROOT’ to get a working copy, edit it, and commit your changes in the normal way.

It is possible to commit an erroneous administrative file. You can often fix the error and check in a new revision, but sometimes a particularly bad error in the administrative file makes it impossible to commit new revisions. If and when this happens, you can correct the problem by temporarily copying a corrected administrative file directly into the $CVSROOT/CVSROOT repository directory, then committing the same correction via a checkout of the ‘CVSROOT’ module. It is important that the correction also be made via the checked out copy, or the next checkout and commit to the <code>CVSROOT</code> module will overwrite the correction that was copied directly into the repository, possibly breaking things in such a way as to prevent commits again.

2.5 Multiple repositories

In some situations it is a good idea to have more than one repository, for instance if you have two development groups that work on separate projects without sharing any code. All you have to do to have several repositories is to specify the appropriate repository, using the CVSROOT environment variable, the ‘-d’ option to cvs, or (once you have checked out a working directory) by simply allowing cvs to use the repository that was used to check out the working directory (see Section 2.1 [Specifying a repository], page 7).

The big advantage of having multiple repositories is that they can reside on different servers. With cvs version 1.10, a single command cannot recurse into directories from different repositories. With development versions of cvs, you can check out code from multiple servers into your working directory. cvs will recurse and handle all the details of making connections to as many server machines as necessary to perform the requested command. Here is an example of how to set up a working directory:

     cvs -d server1:/cvs co dir1
     cd dir1
     cvs -d server2:/root co sdir

     cvs update

The cvs co commands set up the working directory, and then the cvs update command will contact server2, to update the dir1/sdir subdirectory, and server1, to update everything else.

2.6 Creating a repository

This section describes how to set up a cvs repository for any sort of access method. After completing the setup described in this section, you should be able to access your cvs repository immediately via the local access method and several remote access methods. For more information on setting up remote access to the repository you create in this section, please read the section on See Section 2.9 [Remote repositories], page 19.

To set up a cvs repository, first choose the machine and disk on which you want to store the revision history of the source files. CPU and memory requirements are modest, so most machines should be adequate. For details see Section 2.9.1 [Server requirements], page 20.

To estimate disk space requirements, if you are importing RCS files from another system, the size of those files is the approximate initial size of your repository, or if you are starting without any version history, a rule of thumb is to allow for the server approximately three times the size of the code to be under cvs for the repository (you will eventually outgrow this, but not for a while). On the machines on which the developers will be working, you’ll want disk space for approximately one working directory for each developer (either the entire tree or a portion of it, depending on what each developer uses).

The repository should be accessible (directly or via a networked file system) from all machines which want to use cvs in server or local mode; the client machines need not have any access to it other than via the cvs protocol. It is not possible to use cvs to read from a repository which one only has read access to; cvs needs to be able to create lock files (see Section 10.5 [Concurrency], page 67).

To create a repository, run the cvs init command. It will set up an empty repository in the cvs root specified in the usual way (see Chapter 2 [Repository], page 7). For example,

     cvs -d /usr/local/cvsroot init

cvs init is careful to never overwrite any existing files in the repository, so no harm is done if you run cvs init on an already set-up repository.

cvs init will enable history logging; if you don’t want that, remove the history file after running cvs init. See Section C.9 [history file], page 152.

2.7 Backing up a repository

There is nothing particularly magical about the files in the repository; for the most part it is possible to back them up just like any other files. However, there are a few issues to consider.

The first is that to be paranoid, one should either not use cvs during the backup, or have the backup program lock cvs while doing the backup. To not use cvs, you might forbid logins to machines which can access the repository, turn off your cvs server, or similar mechanisms. The details would depend on your operating system and how you have cvs set up. To lock cvs, you would create ‘#cvs.rfl’ locks in each repository directory. See Section 10.5 [Concurrency], page 67, for more on cvs locks. Having said all this, if you just back up without any of these precautions, the results are unlikely to be particularly dire. Restoring from backup, the repository might be in an inconsistent state, but this would not be particularly hard to fix manually.

When you restore a repository from backup, assuming that changes in the repository were made after the time of the backup, working directories which were not affected by the failure may refer to revisions which no longer exist in the repository. Trying to run cvs in such directories will typically produce an error message. One way to get those changes back into the repository is as follows:

• Get a new working directory.

• Copy the files from the working directory from before the failure over to the new working directory (do not copy the contents of the ‘CVS’ directories, of course).

• Working in the new working directory, use commands such as cvs update and cvs diff to figure out what has changed, and then when you are ready, commit the changes into the repository.

2.8 Moving a repository

Just as backing up the files in the repository is pretty much like backing up any other files, if you need to move a repository from one place to another it is also pretty much like just moving any other collection of files.

The main thing to consider is that working directories point to the repository. The simplest way to deal with a moved repository is to just get a fresh working directory after the move. Of course, you’ll want to make sure that the old working directory had been checked in before the move, or you figured out some other way to make sure that you don’t lose any changes. If you really do want to reuse the existing working directory, it should be possible with manual surgery on the ‘CVS/Repository’ files. You can see Section 2.3 [Working directory storage], page 13, for information on the ‘CVS/Repository’ and ‘CVS/Root’ files, but unless you are sure you want to bother, it probably isn’t worth it.

2.9 Remote repositories

Your working copy of the sources can be on a different machine than the repository. Using cvs in this manner is known as client/server operation. You run cvs on a machine which can mount your working directory, known as the client, and tell it to communicate to a machine which can mount the repository, known as the server. Generally, using a remote repository is just like using a local one, except that the format of the repository name is:

     [:method:][[user][:password]@]hostname[:[port]]/path/to/repository

Specifying a password in the repository name is not recommended during checkout, since this will cause cvs to store a cleartext copy of the password in each created directory. cvs login first instead (see Section 2.9.3.2 [Password authentication client], page 24).

The details of exactly what needs to be set up depend on how you are connecting to the server.

If method is not specified, and the repository name contains ‘:’, then the default is ext or server, depending on your platform; both are described in Section 2.9.2 [Connecting via rsh], page 20.

2.9.1 Server requirements

The quick answer to what sort of machine is suitable as a server is that requirements are modest—a server with 32M of memory or even less can handle a fairly large source tree with a fair amount of activity.

The real answer, of course, is more complicated. Estimating the known areas of large memory consumption should be sufficient to estimate memory requirements. There are two such areas documented here; other memory consumption should be small by comparison (if you find that is not the case, let us know, as described in Appendix H [BUGS], page 171, so we can update this documentation).

The first area of big memory consumption is large checkouts, when using the cvs server. The server consists of two processes for each client that it is serving. Memory consumption on the child process should remain fairly small. Memory consumption on the parent process, particularly if the network connection to the client is slow, can be expected to grow to slightly more than the size of the sources in a single directory, or two megabytes, whichever is larger.

Multiplying the size of each cvs server by the number of servers which you expect to have active at one time should give an idea of memory requirements for the server. For the most part, the memory consumed by the parent process probably can be swap space rather than physical memory.

The second area of large memory consumption is diff, when checking in large files. This is required even for binary files. The rule of thumb is to allow about ten times the size of the largest file you will want to check in, although five times may be adequate. For example, if you want to check in a file which is 10 megabytes, you should have 100 megabytes of memory on the machine doing the checkin (the server machine for client/server, or the machine running cvs for non-client/server). This can be swap space rather than physical memory. Because the memory is only required briefly, there is no particular need to allow memory for more than one such checkin at a time.

Resource consumption for the client is even more modest—any machine with enough capacity to run the operating system in question should have little trouble.

For information on disk space requirements, see Section 2.6 [Creating a repository], page 18.

2.9.2 Connecting with rsh or ssh

cvs may use the ‘ssh’ protocol to perform these operations, so the remote user host needs to have a either an agent like ssh-agent to hold credentials or a ‘.shosts’ file which grants access to the local user. Note that the program that cvs uses for this purpose may be specified using the ‘--with-ssh’ flag to configure.

cvs uses the ‘rsh’ protocol to perform these operations, so the remote user host needs to have a ‘.rhosts’ file which grants access to the local user. Note that the program that cvs uses for this purpose may be specified using the ‘--with-rsh’ flag to configure.

For example, suppose you are the user ‘mozart’ on the local machine ‘toe.example.com’, and the server machine is ‘faun.example.org’. On faun, put the following line into the file ‘.rhosts’ in ‘bach’’s home directory:

     toe.example.com  mozart

Then test that ‘rsh’ is working with

     rsh -l bach faun.example.org 'echo $PATH'

To test that ‘ssh’ is working use

     ssh -l bach faun.example.org 'echo $PATH'

Next you have to make sure that rsh will be able to find the server. Make sure that the path which rsh printed in the above example includes the directory containing a program named cvs which is the server. You need to set the path in ‘.bashrc’, ‘.cshrc’, etc., not ‘.login’ or ‘.profile’. Alternately, you can set the environment variable CVS_ SERVER on the client machine to the filename of the server you want to use, for example

‘/usr/local/bin/cvs-1.6’.

There is no need to edit ‘inetd.conf’ or start a cvs server daemon.

There are three access methods that you use in CVSROOT for rsh or ssh. :server: specifies an internal rsh client, which is supported only by some cvs ports. :extssh: specifies an external ssh program. By default this is ssh (unless otherwise specified by the ‘--with-ssh’ flag to configure) but you may set the CVS_SSH environment variable to invoke another program or wrapper script. :ext: specifies an external rsh program. By default this is rsh (unless otherwise specified by the ‘--with-rsh’ flag to configure) but you may set the CVS_RSH environment variable to invoke another program which can access the remote server (for example, remsh on HP-UX 9 because rsh is something different). It must be a program which can transmit data to and from the server without modifying it; for example the Windows NT rsh is not suitable since it by default translates between CRLF and LF. The OS/2 cvs port has a hack to pass ‘-b’ to rsh to get around this, but since this could potentially cause problems for programs other than the standard rsh, it may change in the future. If you set CVS_RSH to SSH or some other rsh replacement, the instructions in the rest of this section concerning ‘.rhosts’ and so on are likely to be inapplicable; consult the documentation for your rsh replacement.

Continuing our example, supposing you want to access the module ‘foo’ in the repository

‘/usr/local/cvsroot/’, on machine ‘faun.example.org’, you are ready to go:
     cvs -d :ext:bach@faun.example.org:/usr/local/cvsroot checkout foo

(The ‘bach@’ can be omitted if the username is the same on both the local and remote hosts.)

2.9.3 Direct connection with password authentication

The cvs client can also connect to the server using a password protocol. This is particularly useful if using rsh is not feasible (for example, the server is behind a firewall), and Kerberos also is not available.

To use this method, it is necessary to make some adjustments on both the server and client sides.

2.9.3.1 Setting up the server for password authentication

First of all, you probably want to tighten the permissions on the ‘$CVSROOT’ and ‘$CVSROOT/CVSROOT’ directories. See Section 2.9.3.3 [Password authentication security], page 25, for more details.

On the server side, the file ‘/etc/inetd.conf’ needs to be edited so inetd knows to run the command cvs pserver when it receives a connection on the right port. By default, the port number is 2401; it would be different if your client were compiled with CVS_AUTH_PORT defined to something else, though. This can also be specified in the CVSROOT variable (see Section 2.9 [Remote repositories], page 19) or overridden with the CVS CLIENT PORT environment variable (see Appendix D [Environment variables], page 155).

If your inetd allows raw port numbers in ‘/etc/inetd.conf’, then the following (all on a single line in ‘inetd.conf’) should be sufficient:

     2401  stream  tcp  nowait  root  /usr/local/bin/cvs
     cvs -f --allow-root=/usr/cvsroot pserver

(You could also use the ‘-T’ option to specify a temporary directory.)

The ‘--allow-root’ option specifies the allowable cvsroot directory. Clients which attempt to use a different cvsroot directory will not be allowed to connect. If there is more than one cvsroot directory which you want to allow, repeat the option. Unfortunately, many versions of inetd have very small limits on the number of arguments and/or the total length of the command. The usual solution to this problem is to have inetd run a shell script which then invokes cvs with the necessary arguments:

     #! /bin/sh
     exec /usr/local/bin/cvs -f \
          --allow-root=/repo1 \
          --allow-root=/repo2 \
          ...
          --allow-root=/repoN \
          pserver

If your inetd wants a symbolic service name instead of a raw port number, then put

this in ‘/etc/services’:
     cvspserver      2401/tcp
and put cvspserver instead of 2401 in ‘inetd.conf’.

If your system uses xinetd instead of inetd, the procedure is slightly different. Create a file called ‘/etc/xinetd.d/cvspserver’ containing the following:

     service cvspserver
     {
        port        = 2401
        socket_type = stream
        protocol    = tcp
        wait        = no
        user        = root
        passenv     = PATH
        server      = /usr/local/bin/cvs
        server_args = -f --allow-root=/usr/cvsroot pserver
     }

(If cvspserver is defined in ‘/etc/services’, you can omit the port line.)

Once the above is taken care of, restart your inetd, or do whatever is necessary to force it to reread its initialization files.

If you are having trouble setting this up, see Section F.2 [Connection], page 167.

Because the client stores and transmits passwords in cleartext (almost—see Section 2.9.3.3 [Password authentication security], page 25, for details), a separate cvs password file is generally used, so people don’t compromise their regular passwords when they access the repository. This file is ‘$CVSROOT/CVSROOT/passwd’ (see Section 2.4 [Intro administrative files], page 16). It uses a colon-separated format, similar to ‘/etc/passwd’ on Unix systems, except that it has fewer fields: cvs username, optional password, and an optional system username for cvs to run as if authentication succeeds. Here is an example ‘passwd’ file with five entries:

     anonymous:
     bach:ULtgRLXo7NRxs
     spwang:1sOp854gDF3DY
     melissa:tGX1fS8sun6rY:pubcvs
     qproj:XR4EZcEs0szik:pubcvs

(The passwords are encrypted according to the standard Unix crypt() function, so it is possible to paste in passwords directly from regular Unix ‘/etc/passwd’ files.)

The first line in the example will grant access to any cvs client attempting to authenticate as user anonymous, no matter what password they use, including an empty password. (This is typical for sites granting anonymous read-only access; for information on how to do the "read-only" part, see Section 2.10 [Read-only access], page 27.)

The second and third lines will grant access to bach and spwang if they supply their respective plaintext passwords.

The fourth line will grant access to melissa, if she supplies the correct password, but her cvs operations will actually run on the server side under the system user pubcvs. Thus, there need not be any system user named melissa, but there must be one named pubcvs.

The fifth line shows that system user identities can be shared: any client who successfully authenticates as qproj will actually run as pubcvs, just as melissa does. That way you could create a single, shared system user for each project in your repository, and give each developer their own line in the ‘$CVSROOT/CVSROOT/passwd’ file. The cvs username on each line would be different, but the system username would be the same. The reason to have different cvs usernames is that cvs will log their actions under those names: when melissa commits a change to a project, the checkin is recorded in the project’s history under the name melissa, not pubcvs. And the reason to have them share a system username is so that you can arrange permissions in the relevant area of the repository such that only that account has write-permission there.

If the system-user field is present, all password-authenticated cvs commands run as that user; if no system user is specified, cvs simply takes the cvs username as the system username and runs commands as that user. In either case, if there is no such user on the system, then the cvs operation will fail (regardless of whether the client supplied a valid password).

The password and system-user fields can both be omitted (and if the system-user field is omitted, then also omit the colon that would have separated it from the encrypted password). For example, this would be a valid ‘$CVSROOT/CVSROOT/passwd’ file:

     anonymous::pubcvs
     fish:rKa5jzULzmhOo:kfogel
     sussman:1sOp854gDF3DY

When the password field is omitted or empty, then the client’s authentication attempt will succeed with any password, including the empty string. However, the colon after the cvs username is always necessary, even if the password is empty.

cvs can also fall back to use system authentication. When authenticating a password, the server first checks for the user in the ‘$CVSROOT/CVSROOT/passwd’ file. If it finds the user, it will use that entry for authentication as described above. But if it does not find the user, or if the cvs ‘passwd’ file does not exist, then the server can try to authenticate the username and password using the operating system’s user-lookup routines (this "fallback" behavior can be disabled by setting SystemAuth=no in the cvs ‘config’ file, see Section C.11 [config], page 153). Be aware, however, that falling back to system authentication might be a security risk: cvs operations would then be authenticated with that user’s regular login password, and the password flies across the network in plaintext. See Section 2.9.3.3 [Password authentication security], page 25 for more on this.

Right now, the only way to put a password in the cvs ‘passwd’ file is to paste it there from somewhere else. Someday, there may be a cvs passwd command.

Unlike many of the files in ‘$CVSROOT/CVSROOT’, it is normal to edit the ‘passwd’ file in-place, rather than via cvs. This is because of the possible security risks of having the ‘passwd’ file checked out to people’s working copies. If you do want to include the ‘passwd’ file in checkouts of ‘$CVSROOT/CVSROOT’, see Section C.8 [checkoutlist], page 151.

2.9.3.2 Using the client with password authentication

To run a cvs command on a remote repository via the password-authenticating server, one specifies the pserver protocol, optional username, repository host, an optional port number, and path to the repository. For example:

     cvs -d :pserver:faun.example.org:/usr/local/cvsroot checkout someproj

or

     CVSROOT=:pserver:bach@faun.example.org:2401/usr/local/cvsroot
     cvs checkout someproj

However, unless you’re connecting to a public-access repository (i.e., one where that username doesn’t require a password), you’ll need to supply a password or log in first. Logging in verifies your password with the repository and stores it in a file. It’s done with the login command, which will prompt you interactively for the password if you didn’t supply one as part of $CVSROOT:

     cvs -d :pserver:bach@faun.example.org:/usr/local/cvsroot login
     CVS password:

or

     cvs -d :pserver:bach:p4ss30rd@faun.example.org:/usr/local/cvsroot login

After you enter the password, cvs verifies it with the server. If the verification succeeds, then that combination of username, host, repository, and password is permanently recorded, so future transactions with that repository won’t require you to run cvs login. (If verification fails, cvs will exit complaining that the password was incorrect, and nothing will be recorded.)

The records are stored, by default, in the file ‘$HOME/.cvspass’. That file’s format is human-readable, and to a degree human-editable, but note that the passwords are not stored in cleartext—they are trivially encoded to protect them from "innocent" compromise (i.e., inadvertent viewing by a system administrator or other non-malicious person).

You can change the default location of this file by setting the CVS_PASSFILE environment variable. If you use this variable, make sure you set it before cvs login is run. If you were to set it after running cvs login, then later cvs commands would be unable to look up the password for transmission to the server.

Once you have logged in, all cvs commands using that remote repository and username will authenticate with the stored password. So, for example

     cvs -d :pserver:bach@faun.example.org:/usr/local/cvsroot checkout foo

should just work (unless the password changes on the server side, in which case you’ll have to re-run cvs login).

Note that if the ‘:pserver:’ were not present in the repository specification, cvs would assume it should use rsh to connect with the server instead (see Section 2.9.2 [Connecting via rsh], page 20).

Of course, once you have a working copy checked out and are running cvs commands from within it, there is no longer any need to specify the repository explicitly, because cvs can deduce the repository from the working copy’s ‘CVS’ subdirectory.

The password for a given remote repository can be removed from the CVS_PASSFILE by using the cvs logout command.

2.9.3.3 Security considerations with password authentication

The passwords are stored on the client side in a trivial encoding of the cleartext, and transmitted in the same encoding. The encoding is done only to prevent inadvertent password compromises (i.e., a system administrator accidentally looking at the file), and will not prevent even a naive attacker from gaining the password.

The separate cvs password file (see Section 2.9.3.1 [Password authentication server], page 21) allows people to use a different password for repository access than for login access. On the other hand, once a user has non-read-only access to the repository, she can execute programs on the server system through a variety of means. Thus, repository access implies fairly broad system access as well. It might be possible to modify cvs to prevent that, but no one has done so as of this writing.

Note that because the ‘$CVSROOT/CVSROOT’ directory contains ‘passwd’ and other files which are used to check security, you must control the permissions on this directory as tightly as the permissions on ‘/etc’. The same applies to the ‘$CVSROOT’ directory itself and any directory above it in the tree. Anyone who has write access to such a directory will have the ability to become any user on the system. Note that these permissions are typically tighter than you would use if you are not using pserver.

In summary, anyone who gets the password gets repository access (which may imply some measure of general system access as well). The password is available to anyone who can sniff network packets or read a protected (i.e., user read-only) file. If you want real security, get Kerberos.

2.9.4 Direct connection with GSSAPI

GSSAPI is a generic interface to network security systems such as Kerberos 5. If you have a working GSSAPI library, you can have cvs connect via a direct tcp connection, authenticating with GSSAPI.

To do this, cvs needs to be compiled with GSSAPI support; when configuring cvs it tries to detect whether GSSAPI libraries using Kerberos version 5 are present. You can also use the ‘--with-gssapi’ flag to configure.

The connection is authenticated using GSSAPI, but the message stream is not authenticated by default. You must use the -a global option to request stream authentication.

The data transmitted is not encrypted by default. Encryption support must be compiled into both the client and the server; use the ‘--enable-encrypt’ configure option to turn it on. You must then use the -x global option to request encryption.

GSSAPI connections are handled on the server side by the same server which handles the password authentication server; see Section 2.9.3.1 [Password authentication server], page 21. If you are using a GSSAPI mechanism such as Kerberos which provides for strong authentication, you will probably want to disable the ability to authenticate via cleartext passwords. To do so, create an empty ‘CVSROOT/passwd’ password file, and set SystemAuth=no in the config file (see Section C.11 [config], page 153).

The GSSAPI server uses a principal name of cvs/hostname, where hostname is the canonical name of the server host. You will have to set this up as required by your GSSAPI mechanism.

To connect using GSSAPI, use the ‘:gserver:’ method. For example,

     cvs -d :gserver:faun.example.org:/usr/local/cvsroot checkout foo

2.9.5 Direct connection with Kerberos

The easiest way to use Kerberos is to use the Kerberos rsh, as described in Section 2.9.2 [Connecting via rsh], page 20. The main disadvantage of using rsh is that all the data needs to pass through additional programs, so it may be slower. So if you have Kerberos installed you can connect via a direct tcp connection, authenticating with Kerberos.

This section concerns the Kerberos network security system, version 4. Kerberos version 5 is supported via the GSSAPI generic network security interface, as described in the previous section.

To do this, cvs needs to be compiled with Kerberos support; when configuring cvs it tries to detect whether Kerberos is present or you can use the ‘--with-krb4’ flag to configure.

The data transmitted is not encrypted by default. Encryption support must be compiled into both the client and server; use the ‘--enable-encryption’ configure option to turn it on. You must then use the -x global option to request encryption.

You need to edit ‘inetd.conf’ on the server machine to run cvs kserver. The client uses port 1999 by default; if you want to use another port specify it in the CVSROOT (see Section 2.9 [Remote repositories], page 19) or the CVS_CLIENT_PORT environment variable (see Appendix D [Environment variables], page 155) on the client.

When you want to use cvs, get a ticket in the usual way (generally kinit); it must be a ticket which allows you to log into the server machine. Then you are ready to go:

     cvs -d :kserver:faun.example.org:/usr/local/cvsroot checkout foo

Previous versions of cvs would fall back to a connection via rsh; this version will not do so.

2.9.6 Connecting with fork

This access method allows you to connect to a repository on your local disk via the remote protocol. In other words it does pretty much the same thing as :local:, but various quirks, bugs and the like are those of the remote cvs rather than the local cvs.

For day-to-day operations you might prefer either :local: or :fork:, depending on your preferences. Of course :fork: comes in particularly handy in testing or debugging cvs and the remote protocol. Specifically, we avoid all of the network-related setup/configuration, timeouts, and authentication inherent in the other remote access methods but still create a connection which uses the remote protocol.

To connect using the fork method, use ‘:fork:’ and the pathname to your local repository. For example:

     cvs -d :fork:/usr/local/cvsroot checkout foo

As with :ext:, the server is called ‘cvs’ by default, or the value of the CVS_SERVER environment variable.

2.10 Read-only repository access

It is possible to grant read-only repository access to people using the password-authenticated server (see Section 2.9.3 [Password authenticated], page 21). (The other access methods do not have explicit support for read-only users because those methods all assume login access to the repository machine anyway, and therefore the user can do whatever local file permissions allow her to do.)

A user who has read-only access can do only those cvs operations which do not modify the repository, except for certain “administrative” files (such as lock files and the history file). It may be desirable to use this feature in conjunction with user-aliasing (see Section 2.9.3.1 [Password authentication server], page 21).

Unlike with previous versions of cvs, read-only users should be able merely to read the repository, and not to execute programs on the server or otherwise gain unexpected levels of access. Or to be more accurate, the known holes have been plugged. Because this feature is new and has not received a comprehensive security audit, you should use whatever level of caution seems warranted given your attitude concerning security.

There are two ways to specify read-only access for a user: by inclusion, and by exclusion.

"Inclusion" means listing that user specifically in the ‘$CVSROOT/CVSROOT/readers’ file, which is simply a newline-separated list of users. Here is a sample ‘readers’ file:

     melissa
     splotnik
     jrandom

(Don’t forget the newline after the last user.)

"Exclusion" means explicitly listing everyone who has write access—if the file

     $CVSROOT/CVSROOT/writers

exists, then only those users listed in it have write access, and everyone else has read-only access (of course, even the read-only users still need to be listed in the cvs ‘passwd’ file). The ‘writers’ file has the same format as the ‘readers’ file.

Note: if your cvs ‘passwd’ file maps cvs users onto system users (see Section 2.9.3.1 [Password authentication server], page 21), make sure you deny or grant read-only access using the cvs usernames, not the system usernames. That is, the ‘readers’ and ‘writers’ files contain cvs usernames, which may or may not be the same as system usernames.

Here is a complete description of the server’s behavior in deciding whether to grant read-only or read-write access:

If ‘readers’ exists, and this user is listed in it, then she gets read-only access. Or if ‘writers’ exists, and this user is NOT listed in it, then she also gets read-only access (this is true even if ‘readers’ exists but she is not listed there). Otherwise, she gets full read-write access.

Of course there is a conflict if the user is listed in both files. This is resolved in the more conservative way, it being better to protect the repository too much than too little: such a user gets read-only access.

2.11 Temporary directories for the server

While running, the cvs server creates temporary directories. They are named

     cvs-servpid

where pid is the process identification number of the server. They are located in the directory specified by the ‘-T’ global option (see Section A.4 [Global options], page 88), the TMPDIR environment variable (see Appendix D [Environment variables], page 155), or, failing that, ‘/tmp’.

In most cases the server will remove the temporary directory when it is done, whether it finishes normally or abnormally. However, there are a few cases in which the server does not or cannot remove the temporary directory, for example:

• If the server aborts due to an internal server error, it may preserve the directory to aid in debugging

• If the server is killed in a way that it has no way of cleaning up (most notably, ‘kill -KILL’ on unix).

• If the system shuts down without an orderly shutdown, which tells the server to clean up.

In cases such as this, you will need to manually remove the ‘cvs-servpid’ directories. As long as there is no server running with process identification number pid, it is safe to do so.

3 Starting a project with CVS

Because renaming files and moving them between directories is somewhat inconvenient, the first thing you do when you start a new project should be to think through your file organization. It is not impossible to rename or move files, but it does increase the potential for confusion and cvs does have some quirks particularly in the area of renaming directories. See Section 7.4 [Moving files], page 56.

What to do next depends on the situation at hand.

3.1 Setting up the files

The first step is to create the files inside the repository. This can be done in a couple of different ways.

3.1.1 Creating a directory tree from a number of files

When you begin using cvs, you will probably already have several projects that can be put under cvs control. In these cases the easiest way is to use the import command. An example is probably the easiest way to explain how to use it. If the files you want to install in cvs reside in ‘wdir’, and you want them to appear in the repository as

‘$CVSROOT/yoyodyne/rdir’, you can do this:
     $ cd wdir
     $ cvs import -m "Imported sources" yoyodyne/rdir yoyo start

Unless you supply a log message with the ‘-m’ flag, cvs starts an editor and prompts for a message. The string ‘yoyo’ is a vendor tag, and ‘start’ is a release tag. They may fill no purpose in this context, but since cvs requires them they must be present. See Chapter 13 [Tracking sources], page 79, for more information about them.

You can now verify that it worked, and remove your original source directory.

     $ cd ..
     $ cvs checkout yoyodyne/rdir       # Explanation below
     $ diff -r wdir yoyodyne/rdir
     $ rm -r wdir

Erasing the original sources is a good idea, to make sure that you do not accidentally edit them in wdir, bypassing cvs. Of course, it would be wise to make sure that you have a backup of the sources before you remove them.

The checkout command can either take a module name as argument (as it has done in all previous examples) or a path name relative to $CVSROOT, as it did in the example above.

It is a good idea to check that the permissions cvs sets on the directories inside $CVSROOT are reasonable, and that they belong to the proper groups. See Section 2.2.2 [File permissions], page 9.

If some of the files you want to import are binary, you may want to use the wrappers features to specify which files are binary and which are not. See Section C.2 [Wrappers], page 142.

3.1.2 Creating Files From Other Version Control Systems

If you have a project which you are maintaining with another version control system, such as rcs, you may wish to put the files from that project into cvs, and preserve the revision history of the files.

From RCS If you have been using rcs, find the rcs files—usually a file named ‘foo.c’ will have its rcs file in ‘RCS/foo.c,v’ (but it could be other places; consult the rcs documentation for details). Then create the appropriate directories in cvs if they do not already exist. Then copy the files into the appropriate directories in the cvs repository (the name in the repository must be the name of the source file with ‘,v’ added; the files go directly in the appropriate directory of the repository, not in an ‘RCS’ subdirectory). This is one of the few times when it is a good idea to access the cvs repository directly, rather than using cvs commands. Then you are ready to check out a new working directory. The rcs file should not be locked when you move it into cvs; if it is, cvs will have trouble letting you operate on it.

From another version control system Many version control systems have the ability to export rcs files in the standard format. If yours does, export the rcs files and then follow the above instructions. Failing that, probably your best bet is to write a script that will check out the files one revision at a time using the command line interface to the other system, and then check the revisions into cvs. The ‘sccs2rcs’ script mentioned below may be a useful example to follow.

From SCCS There is a script in the ‘contrib’ directory of the cvs source distribution called ‘sccs2rcs’ which converts sccs files to rcs files. Note: you must run it on a machine which has both sccs and rcs installed, and like everything else in contrib it is unsupported (your mileage may vary).

From PVCS There is a script in the ‘contrib’ directory of the cvs source distribution called ‘pvcs_to_rcs’ which converts pvcs archives to rcs files. You must run it on a machine which has both pvcs and rcs installed, and like everything else in contrib it is unsupported (your mileage may vary). See the comments in the script for details.

3.1.3 Creating a directory tree from scratch

For a new project, the easiest thing to do is probably to create an empty directory structure, like this:

$ mkdir tc
$ mkdir tc/man
$ mkdir tc/testing

After that, you use the import command to create the corresponding (empty) directory structure inside the repository:

$ cd tc
$ cvs import -m "Created directory structure" yoyodyne/dir yoyo start

This will add yoyodyne/dir as a directory under $CVSROOT. Use checkout to get the new project. Then, use add to add files (and new directories)

as needed.

$ cd ..
$ cvs co yoyodyne/dir

Check that the permissions cvs sets on the directories inside $CVSROOT are reasonable.

3.2 Defining the module

The next step is to define the module in the ‘modules’ file. This is not strictly necessary, but modules can be convenient in grouping together related files and directories.

In simple cases these steps are sufficient to define a module.

1. Get a working copy of the modules file.

$ cvs checkout CVSROOT/modules
$ cd CVSROOT

2. Edit the file and insert a line that defines the module. See Section 2.4 [Intro administrative files], page 16, for an introduction. See Section C.1 [modules], page 139, for a full description of the modules file. You can use the following line to define the module ‘tc’:

tc   yoyodyne/tc

3. Commit your changes to the modules file.

$ cvs commit -m "Added the tc module." modules

4. Release the modules module.

$ cd ..
$ cvs release -d CVSROOT

32 CVS—Concurrent Versions System v1.11.23

4 Revisions

For many uses of cvs, one doesn’t need to worry too much about revision numbers; cvs assigns numbers such as 1.1, 1.2, and so on, and that is all one needs to know. However, some people prefer to have more knowledge and control concerning how cvs assigns revision numbers.

If one wants to keep track of a set of revisions involving more than one file, such as which revisions went into a particular release, one uses a tag, which is a symbolic revision which can be assigned to a numeric revision in each file.

4.1 Revision numbers

Each version of a file has a unique revision number. Revision numbers look like ‘1.1’, ‘1.2’, ‘1.3.2.2’ or even ‘1.3.2.2.4.5’. A revision number always has an even number of period-separated decimal integers. By default revision 1.1 is the first revision of a file. Each successive revision is given a new number by increasing the rightmost number by one. The following figure displays a few revisions, with newer revisions to the right.

            +-----+    +-----+    +-----+    +-----+    +-----+
            ! 1.1 !----! 1.2 !----! 1.3 !----! 1.4 !----! 1.5 !
            +-----+    +-----+    +-----+    +-----+    +-----+

It is also possible to end up with numbers containing more than one period, for example ‘1.3.2.2’. Such revisions represent revisions on branches (see Chapter 5 [Branching and merging], page 41); such revision numbers are explained in detail in Section 5.4 [Branches and revisions], page 43.

4.2 Versions, revisions and releases

A file can have several versions, as described above. Likewise, a software product can have several versions. A software product is often given a version number such as ‘4.1.1’.

Versions in the first sense are called revisions in this document, and versions in the second sense are called releases. To avoid confusion, the word version is almost never used in this document.

4.3 Assigning revisions

By default, cvs will assign numeric revisions by leaving the first number the same and incrementing the second number. For example, 1.1, 1.2, 1.3, etc.

When adding a new file, the second number will always be one and the first number will equal the highest first number of any file in that directory. For example, the current directory contains files whose highest numbered revisions are 1.7, 3.1, and 4.12, then an added file will be given the numeric revision 4.1. (When using client/server cvs, only files that are actually sent to the server are considered.)

Normally there is no reason to care about the revision numbers—it is easier to treat them as internal numbers that cvs maintains, and tags provide a better way to distinguish between things like release 1 versus release 2 of your product (see Section 4.4 [Tags], page 34). However, if you want to set the numeric revisions, the ‘-r’ option to cvs commit can do that. The ‘-r’ option implies the ‘-f’ option, in the sense that it causes the files to be committed even if they are not modified.

For example, to bring all your files up to revision 3.0 (including those that haven’t changed), you might invoke:

     $ cvs commit -r 3.0

Note that the number you specify with ‘-r’ must be larger than any existing revision number. That is, if revision 3.0 exists, you cannot ‘cvs commit -r 1.3’. If you want to maintain several releases in parallel, you need to use a branch (see Chapter 5 [Branching and merging], page 41).

4.4 Tags–Symbolic revisions

The revision numbers live a life of their own. They need not have anything at all to do with the release numbers of your software product. Depending on how you use cvs the revision numbers might change several times between two releases. As an example, some of the source files that make up rcs 5.6 have the following revision numbers:

     ci.c            5.21
     co.c            5.9
     ident.c         5.3
     rcs.c           5.12
     rcsbase.h       5.11
     rcsdiff.c       5.10
     rcsedit.c       5.11
     rcsfcmp.c       5.9
     rcsgen.c        5.10
     rcslex.c        5.11
     rcsmap.c        5.2
     rcsutil.c       5.10

You can use the tag command to give a symbolic name to a certain revision of a file. You can use the ‘-v’ flag to the status command to see all tags that a file has, and which revision numbers they represent. Tag names must start with an uppercase or lowercase letter and can contain uppercase and lowercase letters, digits, ‘-’, and ‘_’. The two tag names BASE and HEAD are reserved for use by cvs. It is expected that future names which are special to cvs will be specially named, for example by starting with ‘.’, rather than being named analogously to BASE and HEAD, to avoid conflicts with actual tag names.

You’ll want to choose some convention for naming tags, based on information such as the name of the program and the version number of the release. For example, one might take the name of the program, immediately followed by the version number with ‘.’ changed to ‘-’, so that cvs 1.9 would be tagged with the name cvs1-9. If you choose a consistent convention, then you won’t constantly be guessing whether a tag is cvs-1-9 or cvs1_9 or what. You might even want to consider enforcing your convention in the ‘taginfo’ file (see Section C.6 [taginfo], page 150).

The following example shows how you can add a tag to a file. The commands must be issued inside your working directory. That is, you should issue the command in the directory where ‘backend.c’ resides.

     $ cvs tag rel-0-4 backend.c
     T backend.c
     $ cvs status -v backend.c
     ===================================================================
     File: backend.c         Status: Up-to-date
         Version:            1.4     Tue Dec  1 14:39:01 1992
         RCS Version:        1.4     /u/cvsroot/yoyodyne/tc/backend.c,v
         Sticky Tag:         (none)
         Sticky Date:        (none)
         Sticky Options:     (none)
         Existing Tags:
             rel-0-4                     (revision: 1.4)

For a complete summary of the syntax of cvs tag, including the various options, see Appendix B [Invoking CVS], page 127.

There is seldom reason to tag a file in isolation. A more common use is to tag all the files that constitute a module with the same tag at strategic points in the development life-cycle, such as when a release is made.

     $ cvs tag rel-1-0 .
     cvs tag: Tagging .
     T Makefile
     T backend.c
     T driver.c
     T frontend.c
     T parser.c

(When you give cvs a directory as argument, it generally applies the operation to all the files in that directory, and (recursively), to any subdirectories that it may contain. See Chapter 6 [Recursive behavior], page 51.)

The checkout command has a flag, ‘-r’, that lets you check out a certain revision of a module. This flag makes it easy to retrieve the sources that make up release 1.0 of the module ‘tc’ at any time in the future:

     $ cvs checkout -r rel-1-0 tc

This is useful, for instance, if someone claims that there is a bug in that release, but you cannot find the bug in the current working copy.

You can also check out a module as it was at any given date. See Section A.9.1 [checkout options], page 99. When specifying ‘-r’ to any of these commands, you will need beware of sticky tags; see Section 4.9 [Sticky tags], page 38.

When you tag more than one file with the same tag you can think about the tag as "a curve drawn through a matrix of filename vs. revision number." Say we have 5 files with the following revisions:

             file1   file2   file3   file4   file5
             1.1     1.1     1.1     1.1  /--1.1*      <-*-  TAG
             1.2*-   1.2     1.2    -1.2*-
             1.3  \- 1.3*-   1.3   / 1.3
             1.4          \  1.4  /  1.4
                           \-1.5*-   1.5
                             1.6

At some time in the past, the * versions were tagged. You can think of the tag as a handle attached to the curve drawn through the tagged revisions. When you pull on the handle, you get all the tagged revisions. Another way to look at it is that you "sight" through a set of revisions that is "flat" along the tagged revisions, like this:

             file1   file2   file3   file4   file5
                             1.1
                             1.2
                     1.1     1.3                       _
             1.1     1.2     1.4     1.1              /
             1.2*----1.3*----1.5*----1.2*----1.1     (--- <--- Look here
             1.3             1.6     1.3              \_
             1.4                     1.4
                                     1.5

4.5 Specifying what to tag from the working directory

The example in the previous section demonstrates one of the most common ways to choose which revisions to tag. Namely, running the cvs tag command without arguments causes cvs to select the revisions which are checked out in the current working directory. For example, if the copy of ‘backend.c’ in working directory was checked out from revision 1.4, then cvs will tag revision 1.4. Note that the tag is applied immediately to revision 1.4 in the repository; tagging is not like modifying a file, or other operations in which one first modifies the working directory and then runs cvs commit to transfer that modification to the repository.

One potentially surprising aspect of the fact that cvs tag operates on the repository is that you are tagging the checked-in revisions, which may differ from locally modified files in your working directory. If you want to avoid doing this by mistake, specify the ‘-c’ option to cvs tag. If there are any locally modified files, cvs will abort with an error before it tags any files:

     $ cvs tag -c rel-0-4
     cvs tag: backend.c is locally modified
     cvs [tag aborted]: correct the above errors first!

4.6 Specifying what to tag by date or revision

The cvs rtag command tags the repository as of a certain date or time (or can be used to tag the latest revision). rtag works directly on the repository contents (it requires no prior checkout and does not look for a working directory).

The following options specify which date or revision to tag. See Section A.5 [Common options], page 90, for a complete description of them.

-D date Tag the most recent revision no later than date.

-f Only useful with the ‘-D date’ or ‘-r tag’ flags. If no matching revision is found, use the most recent revision (instead of ignoring the file).

-r tag Only tag those files that contain existing tag tag.

The cvs tag command also allows one to specify files by revision or date, using the same ‘-r’, ‘-D’, and ‘-f’ options. However, this feature is probably not what you want. The reason is that cvs tag chooses which files to tag based on the files that exist in the working directory, rather than the files which existed as of the given tag/date. Therefore, you are generally better off using cvs rtag. The exceptions might be cases like:

     cvs tag -r 1.4 backend.c

4.7 Deleting, moving, and renaming tags

Normally one does not modify tags. They exist in order to record the history of the repository and so deleting them or changing their meaning would, generally, not be what you want.

However, there might be cases in which one uses a tag temporarily or accidentally puts one in the wrong place. Therefore, one might delete, move, or rename a tag. WARNING: The commands in this section are dangerous; they permanently discard historical information and it can be difficult or impossible to recover from errors. If you are a cvs administrator, you may consider restricting these commands with the ‘taginfo’ file (see Section C.6 [taginfo], page 150).

To delete a tag, specify the ‘-d’ option to either cvs tag or cvs rtag. For example:

     cvs rtag -d rel-0-4 tc

deletes the non-branch tag rel-0-4 from the module tc. In the event that branch tags are encountered within the repository with the given name, a warning message will be issued and the branch tag will not be deleted. If you are absolutely certain you know what you are doing, the -B option may be specified to allow deletion of branch tags. In that case, any non-branch tags encountered will trigger warnings and will not be deleted. WARNING: Moving branch tags is very dangerous! If you think you need the -B option, think again and ask your cvs administrator about it (if that isn’t you). There is almost certainly another way to accomplish what you want to accomplish.

When we say move a tag, we mean to make the same name point to different revisions. For example, the stable tag may currently point to revision 1.4 of ‘backend.c’ and perhaps we want to make it point to revision 1.6. To move a non-branch tag, specify the ‘-F’ option to either cvs tag or cvs rtag. For example, the task just mentioned might be accomplished as:

     cvs tag -r 1.6 -F stable backend.c

If any branch tags are encountered in the repository with the given name, a warning is issued and the branch tag is not disturbed. If you are absolutely certain you wish to move the branch tag, the -B option may be specified. In that case, non-branch tags encountered with the given name are ignored with a warning message.

WARNING: Moving branch tags is very dangerous! If you think you need the -B option, think again and ask your cvs administrator about it (if that isn’t you). There is almost certainly another way to accomplish what you want to accomplish.

When we say rename a tag, we mean to make a different name point to the same revisions as the old tag. For example, one may have misspelled the tag name and want to correct it (hopefully before others are relying on the old spelling). To rename a tag, first create a new tag using the ‘-r’ option to cvs rtag, and then delete the old name. (Caution: this method will not work with branch tags.) This leaves the new tag on exactly the same files as the old tag. For example:

     cvs rtag -r old-name-0-4 rel-0-4 tc
     cvs rtag -d old-name-0-4 tc

4.8 Tagging and adding and removing files

The subject of exactly how tagging interacts with adding and removing files is somewhat obscure; for the most part cvs will keep track of whether files exist or not without too much fussing. By default, tags are applied to only files which have a revision corresponding to what is being tagged. Files which did not exist yet, or which were already removed, simply omit the tag, and cvs knows to treat the absence of a tag as meaning that the file didn’t exist as of that tag.

However, this can lose a small amount of information. For example, suppose a file was added and then removed. Then, if the tag is missing for that file, there is no way to know whether the tag refers to the time before the file was added, or the time after it was removed. If you specify the ‘-r’ option to cvs rtag, then cvs tags the files which have been removed, and thereby avoids this problem. For example, one might specify -r HEAD to tag the head.

On the subject of adding and removing files, the cvs rtag command has a ‘-a’ option which means to clear the tag from removed files that would not otherwise be tagged. For example, one might specify this option in conjunction with ‘-F’ when moving a tag. If one moved a tag without ‘-a’, then the tag in the removed files might still refer to the old revision, rather than reflecting the fact that the file had been removed. I don’t think this is necessary if ‘-r’ is specified, as noted above.

4.9 Sticky tags

Sometimes a working copy’s revision has extra data associated with it, for example it might be on a branch (see Chapter 5 [Branching and merging], page 41), or restricted to versions prior to a certain date by ‘checkout -D’ or ‘update -D’. Because this data persists – that is, it applies to subsequent commands in the working copy – we refer to it as sticky.

Most of the time, stickiness is an obscure aspect of cvs that you don’t need to think about. However, even if you don’t want to use the feature, you may need to know something about sticky tags (for example, how to avoid them!).

You can use the status command to see if any sticky tags or dates are set:

     $ cvs status driver.c
     ===================================================================
     File: driver.c          Status: Up-to-date

         Version:            1.7.2.1 Sat Dec  5 19:35:03 1992
         RCS Version:        1.7.2.1 /u/cvsroot/yoyodyne/tc/driver.c,v
         Sticky Tag:         rel-1-0-patches (branch: 1.7.2)
         Sticky Date:        (none)
         Sticky Options:     (none)

The sticky tags will remain on your working files until you delete them with ‘cvs update -A’. The ‘-A’ option merges local changes into the version of the file from the head of the trunk, removing any sticky tags, dates, or options (other than sticky ‘-k’ options on locally modified files). See Section A.19 [update], page 122 for more on the operation of cvs

update.

The most common use of sticky tags is to identify which branch one is working on, as described in Section 5.3 [Accessing branches], page 42. However, non-branch sticky tags have uses as well. For example, suppose that you want to avoid updating your working directory, to isolate yourself from possibly destabilizing changes other people are making. You can, of course, just refrain from running cvs update. But if you want to avoid updating only a portion of a larger tree, then sticky tags can help. If you check out a certain revision (such as 1.4) it will become sticky. Subsequent cvs update commands will not retrieve the latest revision until you reset the tag with cvs update -A. Likewise, use of the ‘-D’ option to update or checkout sets a sticky date, which, similarly, causes that date to be used for future retrievals.

People often want to retrieve an old version of a file without setting a sticky tag. This can be done with the ‘-p’ option to checkout or update, which sends the contents of the file to standard output. For example:

     $ cvs update -p -r 1.1 file1 >file1
     ===================================================================
     Checking out file1
     RCS:  /tmp/cvs-sanity/cvsroot/first-dir/Attic/file1,v
     VERS: 1.1
     ***************
     $

However, this isn’t the easiest way, if you are asking how to undo a previous checkin (in this example, put ‘file1’ back to the way it was as of revision 1.1). In that case you are better off using the ‘-j’ option to update; for further discussion see Section 5.8 [Merging two revisions], page 46.

40 CVS—Concurrent Versions System v1.11.23

5 Branching and merging

cvs allows you to isolate changes onto a separate line of development, known as a branch. When you change files on a branch, those changes do not appear on the main trunk or other branches.

Later you can move changes from one branch to another branch (or the main trunk) by merging. Merging involves first running cvs update -j, to merge the changes into the working directory. You can then commit that revision, and thus effectively copy the changes onto another branch.

5.1 What branches are good for

Suppose that release 1.0 of tc has been made. You are continuing to develop tc, planning to create release 1.1 in a couple of months. After a while your customers start to complain about a fatal bug. You check out release 1.0 (see Section 4.4 [Tags], page 34) and find the bug (which turns out to have a trivial fix). However, the current revision of the sources are in a state of flux and are not expected to be stable for at least another month. There is no way to make a bug fix release based on the newest sources.

The thing to do in a situation like this is to create a branch on the revision trees for all the files that make up release 1.0 of tc. You can then make modifications to the branch without disturbing the main trunk. When the modifications are finished you can elect to either incorporate them on the main trunk, or leave them on the branch.

5.2 Creating a branch

You can create a branch with tag -b; for example, assuming you’re in a working copy:

     $ cvs tag -b rel-1-0-patches

This splits off a branch based on the current revisions in the working copy, assigning that branch the name ‘rel-1-0-patches’.

It is important to understand that branches get created in the repository, not in the working copy. Creating a branch based on current revisions, as the above example does, will not automatically switch the working copy to be on the new branch. For information on how to do that, see Section 5.3 [Accessing branches], page 42.

You can also create a branch without reference to any working copy, by using rtag:

     $ cvs rtag -b -r rel-1-0 rel-1-0-patches tc

‘-r rel-1-0’ says that this branch should be rooted at the revision that corresponds to the tag ‘rel-1-0’. It need not be the most recent revision – it’s often useful to split a branch off an old revision (for example, when fixing a bug in a past release otherwise known to be stable).

As with ‘tag’, the ‘-b’ flag tells rtag to create a branch (rather than just a symbolic revision name). Note that the numeric revision number that matches ‘rel-1-0’ will probably be different from file to file.

So, the full effect of the command is to create a new branch – named ‘rel-1-0-patches’ – in module ‘tc’, rooted in the revision tree at the point tagged by ‘rel-1-0’.

5.3 Accessing branches

You can retrieve a branch in one of two ways: by checking it out fresh from the repository, or by switching an existing working copy over to the branch.

To check out a branch from the repository, invoke ‘checkout’ with the ‘-r’ flag, followed by the tag name of the branch (see Section 5.2 [Creating a branch], page 41):

     $ cvs checkout -r rel-1-0-patches tc

Or, if you already have a working copy, you can switch it to a given branch with ‘update

-r’:
     $ cvs update -r rel-1-0-patches tc

or equivalently:

     $ cd tc
     $ cvs update -r rel-1-0-patches

It does not matter if the working copy was originally on the main trunk or on some other branch – the above command will switch it to the named branch. And similarly to a regular ‘update’ command, ‘update -r’ merges any changes you have made, notifying you of conflicts where they occur.

Once you have a working copy tied to a particular branch, it remains there until you tell it otherwise. This means that changes checked in from the working copy will add new revisions on that branch, while leaving the main trunk and other branches unaffected.

To find out what branch a working copy is on, you can use the ‘status’ command. In its output, look for the field named ‘Sticky tag’ (see Section 4.9 [Sticky tags], page 38) – that’s cvs’s way of telling you the branch, if any, of the current working files:

     $ cvs status -v driver.c backend.c
     ===================================================================
     File: driver.c          Status: Up-to-date
         Version:            1.7     Sat Dec  5 18:25:54 1992
         RCS Version:        1.7     /u/cvsroot/yoyodyne/tc/driver.c,v
         Sticky Tag:         rel-1-0-patches (branch: 1.7.2)
         Sticky Date:        (none)
         Sticky Options:     (none)
         Existing Tags:
             rel-1-0-patches             (branch: 1.7.2)
             rel-1-0                     (revision: 1.7)
     ===================================================================
     File: backend.c         Status: Up-to-date
         Version:            1.4     Tue Dec  1 14:39:01 1992
         RCS Version:        1.4     /u/cvsroot/yoyodyne/tc/backend.c,v
         Sticky Tag:         rel-1-0-patches (branch: 1.4.2)
         Sticky Date:        (none)
         Sticky Options:     (none)

         Existing Tags:
             rel-1-0-patches             (branch: 1.4.2)
             rel-1-0                     (revision: 1.4)
             rel-0-4                     (revision: 1.4)

Don’t be confused by the fact that the branch numbers for each file are different (‘1.7.2’ and ‘1.4.2’ respectively). The branch tag is the same, ‘rel-1-0-patches’, and the files are indeed on the same branch. The numbers simply reflect the point in each file’s revision history at which the branch was made. In the above example, one can deduce that ‘driver.c’ had been through more changes than ‘backend.c’ before this branch was created.

See Section 5.4 [Branches and revisions], page 43 for details about how branch numbers are constructed.

5.4 Branches and revisions

Ordinarily, a file’s revision history is a linear series of increments (see Section 4.1 [Revision numbers], page 33):

            +-----+    +-----+    +-----+    +-----+    +-----+
            ! 1.1 !----! 1.2 !----! 1.3 !----! 1.4 !----! 1.5 !
            +-----+    +-----+    +-----+    +-----+    +-----+

However, cvs is not limited to linear development. The revision tree can be split into branches, where each branch is a self-maintained line of development. Changes made on one branch can easily be moved back to the main trunk.

Each branch has a branch number, consisting of an odd number of period-separated decimal integers. The branch number is created by appending an integer to the revision number where the corresponding branch forked off. Having branch numbers allows more than one branch to be forked off from a certain revision.

All revisions on a branch have revision numbers formed by appending an ordinal number to the branch number. The following figure illustrates branching with an example.

                                                           +-------------+
                                Branch 1.2.2.3.2 ->        ! 1.2.2.3.2.1 !
                                                         / +-------------+
                                                        /
                                                       /
                      +---------+    +---------+    +---------+
     Branch 1.2.2 -> _! 1.2.2.1 !----! 1.2.2.2 !----! 1.2.2.3 !
                    / +---------+    +---------+    +---------+
                   /
                  /
     +-----+    +-----+    +-----+    +-----+    +-----+
     ! 1.1 !----! 1.2 !----! 1.3 !----! 1.4 !----! 1.5 !  <- The main trunk
     +-----+    +-----+    +-----+    +-----+    +-----+
                     !
                     !
                     !   +---------+    +---------+    +---------+
     Branch 1.2.4 -> +---! 1.2.4.1 !----! 1.2.4.2 !----! 1.2.4.3 !
                         +---------+    +---------+    +---------+

The exact details of how the branch number is constructed is not something you normally need to be concerned about, but here is how it works: When cvs creates a branch number it picks the first unused even integer, starting with 2. So when you want to create a branch from revision 6.4 it will be numbered 6.4.2. All branch numbers ending in a zero (such as 6.4.0) are used internally by cvs (see Section 5.5 [Magic branch numbers], page 44). The branch 1.1.1 has a special meaning. See Chapter 13 [Tracking sources], page 79.

5.5 Magic branch numbers

This section describes a cvs feature called magic branches. For most purposes, you need not worry about magic branches; cvs handles them for you. However, they are visible to you in certain circumstances, so it may be useful to have some idea of how it works.

Externally, branch numbers consist of an odd number of dot-separated decimal integers. See Section 4.1 [Revision numbers], page 33. That is not the whole truth, however. For efficiency reasons cvs sometimes inserts an extra 0 in the second rightmost position (1.2.4 becomes 1.2.0.4, 8.9.10.11.12 becomes 8.9.10.11.0.12 and so on).

cvs does a pretty good job at hiding these so called magic branches, but in a few places the hiding is incomplete:

• The magic branch number appears in the output from cvs log.

• You cannot specify a symbolic branch name to cvs admin.

You can use the admin command to reassign a symbolic name to a branch the way rcs expects it to be. If R4patches is assigned to the branch 1.4.2 (magic branch number 1.4.0.2) in file ‘numbers.c’ you can do this:

     $ cvs admin -NR4patches:1.4.2 numbers.c

It only works if at least one revision is already committed on the branch. Be very careful so that you do not assign the tag to the wrong number. (There is no way to see how the tag was assigned yesterday).

5.6 Merging an entire branch

You can merge changes made on a branch into your working copy by giving the ‘-j branchname’ flag to the update subcommand. With one ‘-j branchname’ option it merges the changes made between the greatest common ancestor (GCA) of the branch and the destination revision (in the simple case below the GCA is the point where the branch forked) and the newest revision on that branch into your working copy.

The ‘-j’ stands for “join”.

Consider this revision tree:

     +-----+    +-----+    +-----+    +-----+
     ! 1.1 !----! 1.2 !----! 1.3 !----! 1.4 !      <- The main trunk
     +-----+    +-----+    +-----+    +-----+
                     !
                     !
                     !   +---------+    +---------+
     Branch R1fix -> +---! 1.2.2.1 !----! 1.2.2.2 !
                         +---------+    +---------+

The branch 1.2.2 has been given the tag (symbolic name) ‘R1fix’. The following example assumes that the module ‘mod’ contains only one file, ‘m.c’.

$ cvs checkout mod # Retrieve the latest revision, 1.4

$ cvs update -j R1fix m.c # Merge all changes made on the branch, # i.e. the changes between revision 1.2 # and 1.2.2.2, into your working copy # of the file.

     $ cvs commit -m "Included R1fix" # Create revision 1.5.

A conflict can result from a merge operation. If that happens, you should resolve it before committing the new revision. See Section 10.3 [Conflicts example], page 65.

If your source files contain keywords (see Chapter 12 [Keyword substitution], page 75), you might be getting more conflicts than strictly necessary. See Section 5.10 [Merging and keywords], page 47, for information on how to avoid this.

The checkout command also supports the ‘-j branchname’ flag. The same effect as above could be achieved with this:

     $ cvs checkout -j R1fix mod
     $ cvs commit -m "Included R1fix"

It should be noted that update -j tagname will also work but may not produce the desired result. See Section 5.9 [Merging adds and removals], page 47, for more.

5.7 Merging from a branch several times

Continuing our example, the revision tree now looks like this:

     +-----+    +-----+    +-----+    +-----+    +-----+
     ! 1.1 !----! 1.2 !----! 1.3 !----! 1.4 !----! 1.5 !   <- The main trunk
     +-----+    +-----+    +-----+    +-----+    +-----+
                     !                           *
                     !                          *
                     !   +---------+    +---------+
     Branch R1fix -> +---! 1.2.2.1 !----! 1.2.2.2 !
                         +---------+    +---------+

where the starred line represents the merge from the ‘R1fix’ branch to the main trunk, as just discussed.

Now suppose that development continues on the ‘R1fix’ branch:

     +-----+    +-----+    +-----+    +-----+    +-----+
     ! 1.1 !----! 1.2 !----! 1.3 !----! 1.4 !----! 1.5 !   <- The main trunk
     +-----+    +-----+    +-----+    +-----+    +-----+
                     !                           *
                     !                          *
                     !   +---------+    +---------+    +---------+
     Branch R1fix -> +---! 1.2.2.1 !----! 1.2.2.2 !----! 1.2.2.3 !
                         +---------+    +---------+    +---------+

and then you want to merge those new changes onto the main trunk. If you just use the cvs update -j R1fix m.c command again, cvs will attempt to merge again the changes which you have already merged, which can have undesirable side effects.

So instead you need to specify that you only want to merge the changes on the branch which have not yet been merged into the trunk. To do that you specify two ‘-j’ options, and cvs merges the changes from the first revision to the second revision. For example, in this case the simplest way would be

     cvs update -j 1.2.2.2 -j R1fix m.c    # Merge changes from 1.2.2.2 to the

# head of the R1fix branch

The problem with this is that you need to specify the 1.2.2.2 revision manually. A slightly better approach might be to use the date the last merge was done:

     cvs update -j R1fix:yesterday -j R1fix m.c

Better yet, tag the R1fix branch after every merge into the trunk, and then use that tag for subsequent merges:

     cvs update -j merged_from_R1fix_to_trunk -j R1fix m.c

5.8 Merging differences between any two revisions

With two ‘-j revision’ flags, the update (and checkout) command can merge the differences between any two revisions into your working file.

     $ cvs update -j 1.5 -j 1.3 backend.c

will undo all changes made between revision 1.3 and 1.5. Note the order of the revisions!

If you try to use this option when operating on multiple files, remember that the numeric revisions will probably be very different between the various files. You almost always use symbolic tags rather than revision numbers when operating on multiple files.

Specifying two ‘-j’ options can also undo file removals or additions. For example, suppose you have a file named ‘file1’ which existed as revision 1.1, and you then removed it (thus adding a dead revision 1.2). Now suppose you want to add it again, with the same contents it had previously. Here is how to do it:

     $ cvs update -j 1.2 -j 1.1 file1
     U file1
     $ cvs commit -m test
     Checking in file1;
     /tmp/cvs-sanity/cvsroot/first-dir/file1,v  <--  file1
     new revision: 1.3; previous revision: 1.2
     done
     $

5.9 Merging can add or remove files

If the changes which you are merging involve removing or adding some files, update -j will reflect such additions or removals.

For example:

     cvs update -A
     touch a b c
     cvs add a b c ; cvs ci -m "added" a b c
     cvs tag -b branchtag
     cvs update -r branchtag
     touch d ; cvs add d
     rm a ; cvs rm a
     cvs ci -m "added d, removed a"
     cvs update -A
     cvs update -jbranchtag

After these commands are executed and a ‘cvs commit’ is done, file ‘a’ will be removed and file ‘d’ added in the main branch.

Note that using a single static tag (‘-j tagname’) rather than a dynamic tag (‘-j branchname’) to merge changes from a branch will usually not remove files which were removed on the branch since cvs does not automatically add static tags to dead revisions. The exception to this rule occurs when a static tag has been attached to a dead revision manually. Use the branch tag to merge all changes from the branch or use two static tags as merge endpoints to be sure that all intended changes are propagated in the merge.

5.10 Merging and keywords

If you merge files containing keywords (see Chapter 12 [Keyword substitution], page 75), you will normally get numerous conflicts during the merge, because the keywords are expanded differently in the revisions which you are merging.

Therefore, you will often want to specify the ‘-kk’ (see Section 12.4 [Substitution modes], page 77) switch to the merge command line. By substituting just the name of the keyword, not the expanded value of that keyword, this option ensures that the revisions which you are merging will be the same as each other, and avoid spurious conflicts.

For example, suppose you have a file like this:

       +---------+
      _! 1.1.2.1 !   <-  br1
     / +---------+
    /
   /
+-----+    +-----+
! 1.1 !----! 1.2 !
+-----+    +-----+

and your working directory is currently on the trunk (revision 1.2). Then you might get the following results from a merge:

$ cat file1
key $Revision
: 1.2 $
. . .
$ cvs update -j br1
U file1
RCS file: /cvsroot/first-dir/file1,v
retrieving revision 1.1
retrieving revision 1.1.2.1
Merging differences between 1.1 and 1.1.2.1 into file1
rcsmerge: warning: conflicts during merge
$ cat file1
<<<<<<< file1
key $Revision
: 1.2 $
=======
key $Revision
: 1.1.2.1 $
>>>>>>> 1.1.2.1
. . .

What happened was that the merge tried to merge the differences between 1.1 and 1.1.2.1 into your working directory. So, since the keyword changed from Revision: 1.1 to Revision: 1.1.2.1, cvs tried to merge that change into your working directory, which conflicted with the fact that your working directory had contained Revision: 1.2.

Here is what happens if you had used ‘-kk’:

$ cat file1
key $Revision
: 1.2 $
. . .
$ cvs update -kk -j br1
U file1
RCS file: /cvsroot/first-dir/file1,v

     retrieving revision 1.1
     retrieving revision 1.1.2.1
     Merging differences between 1.1 and 1.1.2.1 into file1
     $ cat file1
     key $Revision
     $
     . . .

What is going on here is that revision 1.1 and 1.1.2.1 both expand as plain Revision, and therefore merging the changes between them into the working directory need not change anything. Therefore, there is no conflict.

There is, however, one major caveat with using ‘-kk’ on merges. Namely, it overrides whatever keyword expansion mode cvs would normally have used. In particular, this is a problem if the mode had been ‘-kb’ for a binary file. Therefore, if your repository contains binary files, you will need to deal with the conflicts rather than using ‘-kk’.

As a result of using ‘-kk’ during the merge, each file examined by the update will have ‘-kk’ set as sticky options. Running update -A will clear the sticky options on unmodified files, but it will not clear the sticky options on modified files. To get back to the default keyword substitution for modified files, you must commit the results of the merge and then

run update -A.

50 CVS—Concurrent Versions System v1.11.23

6 Recursive behavior

Almost all of the subcommands of cvs work recursively when you specify a directory as an argument. For instance, consider this directory structure:

$HOME
|
+--tc
|   |
+--CVS
|      (internal cvs files)
+--Makefile
+--backend.c
+--driver.c
+--frontend.c
+--parser.c
+--man
|    |
|    +--CVS
|    |  (internal cvs files)
|    +--tc.1
|
+--testing
     |
     +--CVS
     |  (internal cvs files)
     +--testpgm.t
     +--test2.t

If ‘tc’ is the current working directory, the following is true:

• ‘cvs update testing’ is equivalent to

cvs update testing/testpgm.t testing/test2.t

• ‘cvs update testing man’ updates all files in the subdirectories

• ‘cvs update .’ or just ‘cvs update’ updates all files in the tc directory

If no arguments are given to update it will update all files in the current working directory and all its subdirectories. In other words, ‘.’ is a default argument to update. This is also true for most of the cvs subcommands, not only the update command.

The recursive behavior of the cvs subcommands can be turned off with the ‘-l’ option. Conversely, the ‘-R’ option can be used to force recursion if ‘-l’ is specified in ‘~/.cvsrc’ (see Section A.3 [~/.cvsrc], page 88).

$ cvs update -l # Don’t update files in subdirectories 52 CVS—Concurrent Versions System v1.11.23

7 Adding, removing, and renaming files and directories

In the course of a project, one will often add new files. Likewise with removing or renaming, or with directories. The general concept to keep in mind in all these cases is that instead of making an irreversible change you want cvs to record the fact that a change has taken place, just as with modifying an existing file. The exact mechanisms to do this in cvs vary depending on the situation.

7.1 Adding files to a directory

To add a new file to a directory, follow these steps.

• You must have a working copy of the directory. See Section 1.3.1 [Getting the source], page 4.

• Create the new file inside your working copy of the directory.

• Use ‘cvs add filename’ to tell cvs that you want to version control the file. If the file contains binary data, specify ‘-kb’ (see Chapter 9 [Binary files], page 61).

• Use ‘cvs commit filename’ to actually check in the file into the repository. Other developers cannot see the file until you perform this step.

You can also use the add command to add a new directory. Unlike most other commands, the add command is not recursive. You have to explicitly

name files and directories that you wish to add to the repository. However, each directory will need to be added separately before you will be able to add new files to those directories.

$ mkdir -p foo/bar
$ cp ~/myfile foo/bar/myfile
$ cvs add foo foo/bar
$ cvs add foo/bar/myfile

[Command] cvs add [-k kflag] [-m message] files … Schedule files to be added to the repository. The files or directories specified with add must already exist in the current directory. To add a whole new directory hierarchy to the source repository (for example, files received from a third-party vendor), use the import command instead. See Section A.14 [import], page 115. The added files are not placed in the source repository until you use commit to make the change permanent. Doing an add on a file that was removed with the remove command will undo the effect of the remove, unless a commit command intervened. See Section 7.2 [Removing files], page 54, for an example. The ‘-k’ option specifies the default way that this file will be checked out; for more information see Section 12.4 [Substitution modes], page 77. The ‘-m’ option specifies a description for the file. This description appears in the history log (if it is enabled, see Section C.9 [history file], page 152). It will also be saved in the version history inside the repository when the file is committed. The log command displays this description. The description can be changed using ‘admin -t’. See Section A.7 [admin], page 94. If you omit the ‘-m description’ flag, an empty string will be used. You will not be prompted for a description.

For example, the following commands add the file ‘backend.c’ to the repository:

$ cvs add backend.c
$ cvs commit -m "Early version. Not yet compilable." backend.c

When you add a file it is added only on the branch which you are working on (see Chapter 5 [Branching and merging], page 41). You can later merge the additions to another branch if you want (see Section 5.9 [Merging adds and removals], page 47).

7.2 Removing files

Directories change. New files are added, and old files disappear. Still, you want to be able to retrieve an exact copy of old releases.

Here is what you can do to remove a file, but remain able to retrieve old revisions:

• Make sure that you have not made any uncommitted modifications to the file. See Section 1.3.4 [Viewing differences], page 5, for one way to do that. You can also use the status or update command. If you remove the file without committing your changes, you will of course not be able to retrieve the file as it was immediately before you deleted it.

• Remove the file from your working copy of the directory. You can for instance use rm.

• Use ‘cvs remove filename’ to tell cvs that you really want to delete the file.

• Use ‘cvs commit filename’ to actually perform the removal of the file from the repository.

When you commit the removal of the file, cvs records the fact that the file no longer exists. It is possible for a file to exist on only some branches and not on others, or to re-add another file with the same name later. cvs will correctly create or not create the file, based on the ‘-r’ and ‘-D’ options specified to checkout or update.

[Command] cvs remove [options] files …

Schedule file(s) to be removed from the repository (files which have not already been removed from the working directory are not processed). This command does not actually remove the file from the repository until you commit the removal. For a full list of options, see Appendix B [Invoking CVS], page 127.

Here is an example of removing several files:

$ cd test
$ rm *.c
$ cvs remove
cvs remove: Removing .
cvs remove: scheduling a.c for removal
cvs remove: scheduling b.c for removal
cvs remove: use 'cvs commit' to remove these files permanently
$ cvs ci -m "Removed unneeded files"
cvs commit: Examining .
cvs commit: Committing .

As a convenience you can remove the file and cvs remove it in one step, by specifying the ‘-f’ option. For example, the above example could also be done like this:

$ cd test
$ cvs remove -f *.c
cvs remove: scheduling a.c for removal
cvs remove: scheduling b.c for removal
cvs remove: use 'cvs commit' to remove these files permanently
$ cvs ci -m "Removed unneeded files"
cvs commit: Examining .
cvs commit: Committing .

If you execute remove for a file, and then change your mind before you commit, you can undo the remove with an add command.

$ ls
CVS   ja.h  oj.c
$ rm oj.c
$ cvs remove oj.c
cvs remove: scheduling oj.c for removal
cvs remove: use 'cvs commit' to remove this file permanently
$ cvs add oj.c
U oj.c
cvs add: oj.c, version 1.1.1.1, resurrected

If you realize your mistake before you run the remove command you can use update to resurrect the file:

$ rm oj.c
$ cvs update oj.c
cvs update: warning: oj.c was lost
U oj.c

When you remove a file it is removed only on the branch which you are working on (see Chapter 5 [Branching and merging], page 41). You can later merge the removals to another branch if you want (see Section 5.9 [Merging adds and removals], page 47).

7.3 Removing directories

In concept, removing directories is somewhat similar to removing files—you want the directory to not exist in your current working directories, but you also want to be able to retrieve old releases in which the directory existed.

The way that you remove a directory is to remove all the files in it. You don’t remove the directory itself; there is no way to do that. Instead you specify the ‘-P’ option to cvs update or cvs checkout, which will cause cvs to remove empty directories from working directories. (Note that cvs export always removes empty directories.) Probably the best way to do this is to always specify ‘-P’; if you want an empty directory then put a dummy file (for example ‘.keepme’) in it to prevent ‘-P’ from removing it.

Note that ‘-P’ is implied by the ‘-r’ or ‘-D’ options of checkout. This way, cvs will be able to correctly create the directory or not depending on whether the particular version you are checking out contains any files in that directory.

7.4 Moving and renaming files

Moving files to a different directory or renaming them is not difficult, but some of the ways in which this works may be non-obvious. (Moving or renaming a directory is even harder. See Section 7.5 [Moving directories], page 57.).

The examples below assume that the file old is renamed to new.

7.4.1 The Normal way to Rename

The normal way to move a file is to copy old to new, and then issue the normal cvs commands to remove old from the repository, and add new to it.

     $ mv old new
     $ cvs remove old
     $ cvs add new
     $ cvs commit -m "Renamed old to new" old new

This is the simplest way to move a file, it is not error-prone, and it preserves the history of what was done. Note that to access the history of the file you must specify the old or the new name, depending on what portion of the history you are accessing. For example, cvs log old will give the log up until the time of the rename.

When new is committed its revision numbers will start again, usually at 1.1, so if that bothers you, use the ‘-r rev’ option to commit. For more information see Section 4.3 [Assigning revisions], page 33.

7.4.2 Moving the history file

This method is more dangerous, since it involves moving files inside the repository. Read this entire section before trying it out!

     $ cd $CVSROOT/dir
     $ mv old,v new,v

Advantages:

• The log of changes is maintained intact.

• The revision numbers are not affected.

Disadvantages:

• Old releases cannot easily be fetched from the repository. (The file will show up as new even in revisions from the time before it was renamed).

• There is no log information of when the file was renamed.

• Nasty things might happen if someone accesses the history file while you are moving it. Make sure no one else runs any of the cvs commands while you move it.

7.4.3 Copying the history file

This way also involves direct modifications to the repository. It is safe, but not without drawbacks.

# Copy the rcs file inside the repository

     $ cd $CVSROOT/dir
     $ cp old,v new,v

# Remove the old file

     $ cd ~/dir
     $ rm old
     $ cvs remove old
     $ cvs commit old

# Remove all tags from new

     $ cvs update new

$ cvs log new # Remember the non-branch tag names

     $ cvs tag -d tag1 new
     $ cvs tag -d tag2 new
     ...

By removing the tags you will be able to check out old revisions. Advantages:

• Checking out old revisions works correctly, as long as you use ‘-rtag’ and not ‘-Ddate’ to retrieve the revisions.

• The log of changes is maintained intact.

• The revision numbers are not affected.

Disadvantages:

• You cannot easily see the history of the file across the rename.

7.5 Moving and renaming directories

The normal way to rename or move a directory is to rename or move each file within it as described in Section 7.4.1 [Outside], page 56. Then check out with the ‘-P’ option, as described in Section 7.3 [Removing directories], page 55.

If you really want to hack the repository to rename or delete a directory in the repository, you can do it like this:

1. Inform everyone who has a checked out copy of the directory that the directory will be renamed. They should commit all their changes in all their copies of the project containing the directory to be removed, and remove all their working copies of said project, before you take the steps below.

2. Rename the directory inside the repository.

         $ cd $CVSROOT/parent-dir
         $ mv old-dir new-dir

3. Fix the cvs administrative files, if necessary (for instance if you renamed an entire module).

4. Tell everyone that they can check out again and continue working.

If someone had a working copy the cvs commands will cease to work for him, until he removes the directory that disappeared inside the repository.

It is almost always better to move the files in the directory instead of moving the directory. If you move the directory you are unlikely to be able to retrieve old releases correctly, since they probably depend on the name of the directories.

58 CVS—Concurrent Versions System v1.11.23

8 History browsing

Once you have used cvs to store a version control history—what files have changed when, how, and by whom, there are a variety of mechanisms for looking through the history.

8.1 Log messages

Whenever you commit a file you specify a log message.

To look through the log messages which have been specified for every revision which has been committed, use the cvs log command (see Section A.15 [log], page 116).

8.2 The history database

You can use the history file (see Section C.9 [history file], page 152) to log various cvs actions. To retrieve the information from the history file, use the cvs history command (see Section A.13 [history], page 113).

Note: you can control what is logged to this file by using the ‘LogHistory’ keyword in the ‘CVSROOT/config’ file (see Section C.11 [config], page 153).

8.3 User-defined logging

You can customize cvs to log various kinds of actions, in whatever manner you choose. These mechanisms operate by executing a script at various times. The script might append a message to a file listing the information and the programmer who created it, or send mail to a group of developers, or, perhaps, post a message to a particular newsgroup. To log commits, use the ‘loginfo’ file (see Section C.4.4 [loginfo], page 148). To log tags, use the ‘taginfo’ file (see Section C.6 [taginfo], page 150). To log checkouts, exports, and tags, respectively, you can also use the ‘-o’, ‘-e’, and ‘-t’ options in the modules file. For a more flexible way of giving notifications to various users, which requires less in the way of keeping centralized scripts up to date, use the cvs watch add command (see Section 10.6.2 [Getting Notified], page 69); this command is useful even if you are not using cvs watch on.

60 CVS—Concurrent Versions System v1.11.23

9 Handling binary files

The most common use for cvs is to store text files. With text files, cvs can merge revisions, display the differences between revisions in a human-visible fashion, and other such operations. However, if you are willing to give up a few of these abilities, cvs can store binary files. For example, one might store a web site in cvs including both text files and binary images.

9.1 The issues with binary files

While the need to manage binary files may seem obvious if the files that you customarily work with are binary, putting them into version control does present some additional issues.

One basic function of version control is to show the differences between two revisions. For example, if someone else checked in a new version of a file, you may wish to look at what they changed and determine whether their changes are good. For text files, cvs provides this functionality via the cvs diff command. For binary files, it may be possible to extract the two revisions and then compare them with a tool external to cvs (for example, word processing software often has such a feature). If there is no such tool, one must track changes via other mechanisms, such as urging people to write good log messages, and hoping that the changes they actually made were the changes that they intended to make.

Another ability of a version control system is the ability to merge two revisions. For cvs this happens in two contexts. The first is when users make changes in separate working directories (see Chapter 10 [Multiple developers], page 63). The second is when one merges explicitly with the ‘update -j’ command (see Chapter 5 [Branching and merging], page 41).

In the case of text files, cvs can merge changes made independently, and signal a conflict if the changes conflict. With binary files, the best that cvs can do is present the two different copies of the file, and leave it to the user to resolve the conflict. The user may choose one copy or the other, or may run an external merge tool which knows about that particular file format, if one exists. Note that having the user merge relies primarily on the user to not accidentally omit some changes, and thus is potentially error prone.

If this process is thought to be undesirable, the best choice may be to avoid merging. To avoid the merges that result from separate working directories, see the discussion of reserved checkouts (file locking) in Chapter 10 [Multiple developers], page 63. To avoid the merges resulting from branches, restrict use of branches.

9.2 How to store binary files

There are two issues with using cvs to store binary files. The first is that cvs by default converts line endings between the canonical form in which they are stored in the repository (linefeed only), and the form appropriate to the operating system in use on the client (for example, carriage return followed by line feed for Windows NT).

The second is that a binary file might happen to contain data which looks like a keyword (see Chapter 12 [Keyword substitution], page 75), so keyword expansion must be turned off.

The ‘-kb’ option available with some cvs commands insures that neither line ending conversion nor keyword expansion will be done.

Here is an example of how you can create a new file using the ‘-kb’ flag:

     $ echo '$Id
     $' > kotest
     $ cvs add -kb -m"A test file" kotest
     $ cvs ci -m"First checkin; contains a keyword" kotest

If a file accidentally gets added without ‘-kb’, one can use the cvs admin command to recover. For example:

     $ echo '$Id
     $' > kotest
     $ cvs add -m"A test file" kotest
     $ cvs ci -m"First checkin; contains a keyword" kotest
     $ cvs admin -kb kotest
     $ cvs update -A kotest

# For non-unix systems:

# Copy in a good copy of the file from outside CVS

     $ cvs commit -m "make it binary" kotest

When you check in the file ‘kotest’ the file is not preserved as a binary file, because you did not check it in as a binary file. The cvs admin -kb command sets the default keyword substitution method for this file, but it does not alter the working copy of the file that you have. If you need to cope with line endings (that is, you are using cvs on a non-unix system), then you need to check in a new copy of the file, as shown by the cvs commit command above. On unix, the cvs update -A command suffices. (Note that you can use cvs log to determine the default keyword substitution method for a file and cvs status to determine the keyword substitution method for a working copy.)

However, in using cvs admin -k to change the keyword expansion, be aware that the keyword expansion mode is not version controlled. This means that, for example, that if you have a text file in old releases, and a binary file with the same name in new releases, cvs provides no way to check out the file in text or binary mode depending on what version you are checking out. There is no good workaround for this problem.

You can also set a default for whether cvs add and cvs import treat a file as binary based on its name; for example you could say that files who names end in ‘.exe’ are binary. See Section C.2 [Wrappers], page 142. There is currently no way to have cvs detect whether a file is binary based on its contents. The main difficulty with designing such a feature is that it is not clear how to distinguish between binary and non-binary files, and the rules to apply would vary considerably with the operating system.

10 Multiple developers

When more than one person works on a software project things often get complicated. Often, two people try to edit the same file simultaneously. One solution, known as file locking or reserved checkouts, is to allow only one person to edit each file at a time. This is the only solution with some version control systems, including rcs and sccs. Currently the usual way to get reserved checkouts with cvs is the cvs admin -l command (see Section A.7.1 [admin options], page 94). This is not as nicely integrated into cvs as the watch features, described below, but it seems that most people with a need for reserved checkouts find it adequate. It also may be possible to use the watches features described below, together with suitable procedures (not enforced by software), to avoid having two people edit at the same time.

The default model with cvs is known as unreserved checkouts. In this model, developers can edit their own working copy of a file simultaneously. The first person that commits his changes has no automatic way of knowing that another has started to edit it. Others will get an error message when they try to commit the file. They must then use cvs commands to bring their working copy up to date with the repository revision. This process is almost automatic.

cvs also supports mechanisms which facilitate various kinds of communication, without actually enforcing rules like reserved checkouts do.

The rest of this chapter describes how these various models work, and some of the issues involved in choosing between them.

10.1 File status

Based on what operations you have performed on a checked out file, and what operations others have performed to that file in the repository, one can classify a file in a number of states. The states, as reported by the status command, are:

Up-to-date The file is identical with the latest revision in the repository for the branch in use.

Locally Modified You have edited the file, and not yet committed your changes.

Locally Added You have added the file with add, and not yet committed your changes.

Locally Removed You have removed the file with remove, and not yet committed your changes.

Needs Checkout Someone else has committed a newer revision to the repository. The name is slightly misleading; you will ordinarily use update rather than checkout to get that newer revision.

Needs Patch Like Needs Checkout, but the cvs server will send a patch rather than the entire file. Sending a patch or sending an entire file accomplishes the same thing.

Needs Merge Someone else has committed a newer revision to the repository, and you have also made modifications to the file.

Unresolved Conflict A file with the same name as this new file has been added to the repository from a second workspace. This file will need to be moved out of the way to allow an update to complete.

File had conflicts on merge This is like Locally Modified, except that a previous update command gave a conflict. If you have not already done so, you need to resolve the conflict as described in Section 10.3 [Conflicts example], page 65.

Unknown cvs doesn’t know anything about this file. For example, you have created a new file and have not run add.

To help clarify the file status, status also reports the Working revision which is the revision that the file in the working directory derives from, and the Repository revision which is the latest revision in the repository for the branch in use.

The options to status are listed in Appendix B [Invoking CVS], page 127. For information on its Sticky tag and Sticky date output, see Section 4.9 [Sticky tags], page 38. For information on its Sticky options output, see the ‘-k’ option in Section A.19.1 [update options], page 123.

You can think of the status and update commands as somewhat complementary. You use update to bring your files up to date, and you can use status to give you some idea of what an update would do (of course, the state of the repository might change before you actually run update). In fact, if you want a command to display file status in a more brief format than is displayed by the status command, you can invoke

     $ cvs -n -q update

The ‘-n’ option means to not actually do the update, but merely to display statuses; the ‘-q’ option avoids printing the name of each directory. For more information on the update command, and these options, see Appendix B [Invoking CVS], page 127.

10.2 Bringing a file up to date

When you want to update or merge a file, use the cvs update -d command. For files that are not up to date this is roughly equivalent to a checkout command: the newest revision of the file is extracted from the repository and put in your working directory. The -d option, not necessary with checkout, tells cvs that you wish it to create directories added by other developers.

Your modifications to a file are never lost when you use update. If no newer revision exists, running update has no effect. If you have edited the file, and a newer revision is available, cvs will merge all changes into your working copy.

For instance, imagine that you checked out revision 1.4 and started editing it. In the meantime someone else committed revision 1.5, and shortly after that revision 1.6. If you run update on the file now, cvs will incorporate all changes between revision 1.4 and 1.6 into your file.

If any of the changes between 1.4 and 1.6 were made too close to any of the changes you have made, an overlap occurs. In such cases a warning is printed, and the resulting file includes both versions of the lines that overlap, delimited by special markers. See Section A.19 [update], page 122, for a complete description of the update command.

10.3 Conflicts example

Suppose revision 1.4 of ‘driver.c’ contains this:

#include <stdio.h>
void main()
{
parse();
if (nerr == 0)
    gencode();
else
    fprintf(stderr, "No code generated.\n");
exit(nerr == 0 ? 0 : 1);
}

Revision 1.6 of ‘driver.c’ contains this:

#include <stdio.h>
int main(int argc,
     char **argv)
{
parse();
if (argc != 1)
{
    fprintf(stderr, "tc: No args expected.\n");
    exit(1);
}
if (nerr == 0)
    gencode();
else
    fprintf(stderr, "No code generated.\n");
exit(!!nerr);
}

Your working copy of ‘driver.c’, based on revision 1.4, contains this before you run ‘cvs

update’:
#include <stdlib.h>
#include <stdio.h>
void main()
{
init_scanner();
parse();

    if (nerr == 0)
        gencode();
    else
        fprintf(stderr, "No code generated.\n");
    exit(nerr == 0 ? EXIT_SUCCESS : EXIT_FAILURE);
}
You run ‘cvs update’:
$ cvs update driver.c
RCS file: /usr/local/cvsroot/yoyodyne/tc/driver.c,v
retrieving revision 1.4
retrieving revision 1.6
Merging differences between 1.4 and 1.6 into driver.c
rcsmerge warning: overlaps during merge
cvs update: conflicts found in driver.c
C driver.c

cvs tells you that there were some conflicts. Your original working file is saved unmodified in ‘.#driver.c.1.4’. The new version of ‘driver.c’ contains this:

#include <stdlib.h>
#include <stdio.h>
int main(int argc,
         char **argv)
{
    init_scanner();
    parse();
    if (argc != 1)
    {
        fprintf(stderr, "tc: No args expected.\n");
        exit(1);
    }
    if (nerr == 0)
        gencode();
    else
        fprintf(stderr, "No code generated.\n");
<<<<<<< driver.c
    exit(nerr == 0 ? EXIT_SUCCESS : EXIT_FAILURE);
=======
    exit(!!nerr);
>>>>>>> 1.6
}

Note how all non-overlapping modifications are incorporated in your working copy, and that the overlapping section is clearly marked with ‘<<<<<<<’, ‘=======’ and ‘>>>>>>>’.

You resolve the conflict by editing the file, removing the markers and the erroneous line. Suppose you end up with this file:

     #include <stdlib.h>
     #include <stdio.h>
     int main(int argc,
              char **argv)
     {
         init_scanner();
         parse();
         if (argc != 1)
         {
             fprintf(stderr, "tc: No args expected.\n");
             exit(1);
         }
         if (nerr == 0)
             gencode();
         else
             fprintf(stderr, "No code generated.\n");
         exit(nerr == 0 ? EXIT_SUCCESS : EXIT_FAILURE);
     }

You can now go ahead and commit this as revision 1.7.

     $ cvs commit -m "Initialize scanner. Use symbolic exit values." driver.c
     Checking in driver.c;
     /usr/local/cvsroot/yoyodyne/tc/driver.c,v  <--  driver.c
     new revision: 1.7; previous revision: 1.6
     done

For your protection, cvs will refuse to check in a file if a conflict occurred and you have not resolved the conflict. Currently to resolve a conflict, you must change the timestamp on the file. In previous versions of cvs, you also needed to insure that the file contains no conflict markers. Because your file may legitimately contain conflict markers (that is, occurrences of ‘>>>>>>> ’ at the start of a line that don’t mark a conflict), the current version of cvs will print a warning and proceed to check in the file.

If you use release 1.04 or later of pcl-cvs (a gnu Emacs front-end for cvs) you can use an Emacs package called emerge to help you resolve conflicts. See the documentation for pcl-cvs.

10.4 Informing others about commits

It is often useful to inform others when you commit a new revision of a file. The ‘loginfo’ file can be used to automate this process. See Section C.4.4 [loginfo], page 148. You can use these features of cvs to, for instance, instruct cvs to mail a message to all developers, or post a message to a local newsgroup.

10.5 Several developers simultaneously attempting to run CVS

If several developers try to run cvs at the same time, one may get the following message:

     [11:43:23] waiting for bach's lock in /usr/local/cvsroot/foo

cvs will try again every 30 seconds, and either continue with the operation or print the message again, if it still needs to wait. If a lock seems to stick around for an undue amount of time, find the person holding the lock and ask them about the cvs command they are running. If they aren’t running a cvs command, look in the repository directory mentioned in the message and remove files which they own whose names start with ‘#cvs.rfl’,

‘#cvs.wfl’, or ‘#cvs.lock’.

Note that these locks are to protect cvs’s internal data structures and have no relationship to the word lock in the sense used by rcs—which refers to reserved checkouts (see Chapter 10 [Multiple developers], page 63).

Any number of people can be reading from a given repository at a time; only when someone is writing do the locks prevent other people from reading or writing.

One might hope for the following property:

If someone commits some changes in one cvs command, then an update by someone else will either get all the changes, or none of them.

but cvs does not have this property. For example, given the files

     a/one.c
     a/two.c
     b/three.c
     b/four.c

if someone runs

     cvs ci a/two.c b/three.c

and someone else runs cvs update at the same time, the person running update might get only the change to ‘b/three.c’ and not the change to ‘a/two.c’.

10.6 Mechanisms to track who is editing files

For many groups, use of cvs in its default mode is perfectly satisfactory. Users may sometimes go to check in a modification only to find that another modification has intervened, but they deal with it and proceed with their check in. Other groups prefer to be able to know who is editing what files, so that if two people try to edit the same file they can choose to talk about who is doing what when rather than be surprised at check in time. The features in this section allow such coordination, while retaining the ability of two developers to edit the same file at the same time.

For maximum benefit developers should use cvs edit (not chmod) to make files readwrite to edit them, and cvs release (not rm) to discard a working directory which is no longer in use, but cvs is not able to enforce this behavior.

10.6.1 Telling CVS to watch certain files

To enable the watch features, you first specify that certain files are to be watched.

[Command] cvs watch on [-lR] [files]. . .

Specify that developers should run cvs edit before editing files. cvs will create working copies of files read-only, to remind developers to run the cvs edit command before working on them.

If files includes the name of a directory, cvs arranges to watch all files added to the corresponding repository directory, and sets a default for files added in the future; this allows the user to set notification policies on a per-directory basis. The contents of the directory are processed recursively, unless the -l option is given. The -R option can be used to force recursion if the -l option is set in ‘~/.cvsrc’ (see Section A.3 [~/.cvsrc], page 88).

If files is omitted, it defaults to the current directory.

[Command] cvs watch off [-lR] [files]. . .

Do not create files read-only on checkout; thus, developers will not be reminded to

use cvs edit and cvs unedit.

The files and options are processed as for cvs watch on.

10.6.2 Telling CVS to notify you

You can tell cvs that you want to receive notifications about various actions taken on a file. You can do this without using cvs watch on for the file, but generally you will want to use cvs watch on, to remind developers to use the cvs edit command.

[Command] cvs watch add [-lR] [-a action]. . . [files]. . .

Add the current user to the list of people to receive notification of work done on files.

The -a option specifies what kinds of events cvs should notify the user about. action is one of the following:

edit Another user has applied the cvs edit command (described below) to a watched file.

commit Another user has committed changes to one of the named files.

unedit Another user has abandoned editing a file (other than by committing

changes). They can do this in several ways, by: • applying the cvs unedit command (described below) to the file • applying the cvs release command (see Section A.17 [release], page 120) to the file’s parent directory (or recursively to a directory more than one level up)

• deleting the file and allowing cvs update to recreate it

all All of the above.

none None of the above. (This is useful with cvs edit, described below.)

The -a option may appear more than once, or not at all. If omitted, the action defaults to all.

The files and options are processed as for cvs watch on.

[Command] cvs watch remove [-lR] [-a action]. . . [files]. . .

Remove a notification request established using cvs watch add; the arguments are the same. If the -a option is present, only watches for the specified actions are removed.

When the conditions exist for notification, cvs calls the ‘notify’ administrative file. Edit ‘notify’ as one edits the other administrative files (see Section 2.4 [Intro administrative files], page 16). This file follows the usual conventions for administrative files (see Section C.3.1 [syntax], page 143), where each line is a regular expression followed by a command to execute. The command should contain a single occurrence of ‘%s’ which will be replaced by the user to notify; the rest of the information regarding the notification will be supplied to the command on standard input. The standard thing to put in the notify file is the single line:

     ALL mail %s -s "CVS notification"

This causes users to be notified by electronic mail.

Note that if you set this up in the straightforward way, users receive notifications on the server machine. One could of course write a ‘notify’ script which directed notifications elsewhere, but to make this easy, cvs allows you to associate a notification address for each user. To do so create a file ‘users’ in ‘CVSROOT’ with a line for each user in the format user:value. Then instead of passing the name of the user to be notified to ‘notify’, cvs will pass the value (normally an email address on some other machine).

cvs does not notify you for your own changes. Currently this check is done based on whether the user name of the person taking the action which triggers notification matches the user name of the person getting notification. In fact, in general, the watches features only track one edit by each user. It probably would be more useful if watches tracked each working directory separately, so this behavior might be worth changing.

10.6.3 How to edit a file which is being watched

Since a file which is being watched is checked out read-only, you cannot simply edit it. To make it read-write, and inform others that you are planning to edit it, use the cvs edit command. Some systems call this a checkout, but cvs uses that term for obtaining a copy of the sources (see Section 1.3.1 [Getting the source], page 4), an operation which those systems call a get or a fetch.

[Command] cvs edit [-lR] [-a action]. . . [files]. . .

Prepare to edit the working files files. cvs makes the files read-write, and notifies users who have requested edit notification for any of files. The cvs edit command accepts the same options as the cvs watch add command, and establishes a temporary watch for the user on files; cvs will remove the watch when files are unedited or committed. If the user does not wish to receive notifications, she should specify -a none. The files and the options are processed as for the cvs watch commands.

Normally when you are done with a set of changes, you use the cvs commit command, which checks in your changes and returns the watched files to their usual read-only state. But if you instead decide to abandon your changes, or not to make any changes, you can use the cvs unedit command.

[Command] cvs unedit [-lR] [files]. . .

Abandon work on the working files files, and revert them to the repository versions on which they are based. cvs makes those files read-only for which users have requested notification using cvs watch on. cvs notifies users who have requested unedit notification for any of files. The files and options are processed as for the cvs watch commands. If watches are not in use, the unedit command probably does not work, and the way to revert to the repository version is with the command cvs update -C file (see Section A.19 [update], page 122). The meaning is not precisely the same; the latter may also bring in some changes which have been made in the repository since the last time you updated.

When using client/server cvs, you can use the cvs edit and cvs unedit commands even if cvs is unable to successfully communicate with the server; the notifications will be sent upon the next successful cvs command.

10.6.4 Information about who is watching and editing

[Command] cvs watchers [-lR] [files]. . .

List the users currently watching changes to files. The report includes the files being watched, and the mail address of each watcher. The files and options are processed as for the cvs watch commands.

[Command] cvs editors [-lR] [files]. . .

List the users currently working on files. The report includes the mail address of each user, the time when the user began working with the file, and the host and path of the working directory containing the file. The files and options are processed as for the cvs watch commands.

10.6.5 Using watches with old versions of CVS

If you use the watch features on a repository, it creates ‘CVS’ directories in the repository and stores the information about watches in that directory. If you attempt to use cvs 1.6 or earlier with the repository, you get an error message such as the following (all on one line):

     cvs update: cannot open CVS/Entries for reading:
     No such file or directory

and your operation will likely be aborted. To use the watch features, you must upgrade all copies of cvs which use that repository in local or server mode. If you cannot upgrade, use the watch off and watch remove commands to remove all watches, and that will restore the repository to a state which cvs 1.6 can cope with.

10.7 Choosing between reserved or unreserved checkouts

Reserved and unreserved checkouts each have pros and cons. Let it be said that a lot of this is a matter of opinion or what works given different groups’ working styles, but here is a brief description of some of the issues. There are many ways to organize a team of developers. cvs does not try to enforce a certain organization. It is a tool that can be used in several ways.

Reserved checkouts can be very counter-productive. If two persons want to edit different parts of a file, there may be no reason to prevent either of them from doing so. Also, it is common for someone to take out a lock on a file, because they are planning to edit it, but then forget to release the lock.

People, especially people who are familiar with reserved checkouts, often wonder how often conflicts occur if unreserved checkouts are used, and how difficult they are to resolve. The experience with many groups is that they occur rarely and usually are relatively straightforward to resolve.

The rarity of serious conflicts may be surprising, until one realizes that they occur only when two developers disagree on the proper design for a given section of code; such a disagreement suggests that the team has not been communicating properly in the first place. In order to collaborate under any source management regimen, developers must agree on the general design of the system; given this agreement, overlapping changes are usually straightforward to merge.

In some cases unreserved checkouts are clearly inappropriate. If no merge tool exists for the kind of file you are managing (for example word processor files or files edited by Computer Aided Design programs), and it is not desirable to change to a program which uses a mergeable data format, then resolving conflicts is going to be unpleasant enough that you generally will be better off to simply avoid the conflicts instead, by using reserved checkouts.

The watches features described above in Section 10.6 [Watches], page 68 can be considered to be an intermediate model between reserved checkouts and unreserved checkouts. When you go to edit a file, it is possible to find out who else is editing it. And rather than having the system simply forbid both people editing the file, it can tell you what the situation is and let you figure out whether it is a problem in that particular case or not. Therefore, for some groups it can be considered the best of both the reserved checkout and unreserved checkout worlds.

11 Revision management

If you have read this far, you probably have a pretty good grasp on what cvs can do for you. This chapter talks a little about things that you still have to decide.

If you are doing development on your own using cvs you could probably skip this chapter. The questions this chapter takes up become more important when more than one person is working in a repository.

11.1 When to commit?

Your group should decide which policy to use regarding commits. Several policies are possible, and as your experience with cvs grows you will probably find out what works for you.

If you commit files too quickly you might commit files that do not even compile. If your partner updates his working sources to include your buggy file, he will be unable to compile the code. On the other hand, other persons will not be able to benefit from the improvements you make to the code if you commit very seldom, and conflicts will probably be more common.

It is common to only commit files after making sure that they can be compiled. Some sites require that the files pass a test suite. Policies like this can be enforced using the commitinfo file (see Section C.4.1 [commitinfo], page 144), but you should think twice before you enforce such a convention. By making the development environment too controlled it might become too regimented and thus counter-productive to the real goal, which is to get software written.

74 CVS—Concurrent Versions System v1.11.23

12 Keyword substitution

As long as you edit source files inside a working directory you can always find out the state of your files via ‘cvs status’ and ‘cvs log’. But as soon as you export the files from your development environment it becomes harder to identify which revisions they are.

cvs can use a mechanism known as keyword substitution (or keyword expansion) to help identifying the files. Embedded strings of the form $keyword$ and $keyword:...$ in a file are replaced with strings of the form $keyword:value$ whenever you obtain a new revision of the file.

12.1 Keyword List

This is a list of the keywords:

$Author$ The login name of the user who checked in the revision.

$Date$ The date and time (UTC) the revision was checked in.

$Header$ A standard header containing the full pathname of the rcs file, the revision number, the date (UTC), the author, the state, and the locker (if locked). Files will normally never be locked when you use cvs.

$Id$ Same as $Header$, except that the rcs filename is without a path.

$Name$ Tag name used to check out this file. The keyword is expanded only if one checks out with an explicit tag name. For example, when running the command cvs co -r first, the keyword expands to ‘Name: first’.

$Locker$ The login name of the user who locked the revision (empty if not locked, which is the normal case unless cvs admin -l is in use).

$Log$ The log message supplied during commit, preceded by a header containing the rcs filename, the revision number, the author, and the date (UTC). Existing log messages are not replaced. Instead, the new log message is inserted after $Log:...$. Each new line is prefixed with the same string which precedes the $Log keyword. For example, if the file contains:

                 /* Here is what people have been up to:
                  *
                  * $Log
               : frob.c,v $
                  * Revision 1.1  1997/01/03 14:23:51  joe
                  * Add the superfrobnicate option
                  *
                  */

then additional lines which are added when expanding the $Log keyword will be preceded by ‘ * ’. Unlike previous versions of cvs and rcs, the comment leader from the rcs file is not used. The $Log keyword is useful for accumulating a complete change log in a source file, but for several reasons it can be problematic. See Section 12.5 [Log keyword], page 77.

$RCSfile$

The name of the RCS file without a path.

$Revision$

The revision number assigned to the revision.

$Source$ The full pathname of the RCS file.

$State$ The state assigned to the revision. States can be assigned with cvs admin -s—see Section A.7.1 [admin options], page 94.

12.2 Using keywords

To include a keyword string you simply include the relevant text string, such as $Id$, inside the file, and commit the file. cvs will automatically (Or, more accurately, as part of the update run that automatically happens after a commit.) expand the string as part of the commit operation.

It is common to embed the $Id$ string in the source files so that it gets passed through to generated files. For example, if you are managing computer program source code, you might include a variable which is initialized to contain that string. Or some C compilers may provide a #pragma ident directive. Or a document management system might provide a way to pass a string through to generated files.

The ident command (which is part of the rcs package) can be used to extract keywords and their values from a file. This can be handy for text files, but it is even more useful for extracting keywords from binary files.

     $ ident samp.c
     samp.c:
          $Id
     : samp.c,v 1.5 1993/10/19 14:57:32 ceder Exp $
     $ gcc samp.c
     $ ident a.out
     a.out:
          $Id
     : samp.c,v 1.5 1993/10/19 14:57:32 ceder Exp $

Sccs is another popular revision control system. It has a command, what, which is very similar to ident and used for the same purpose. Many sites without rcs have sccs. Since what looks for the character sequence @(#) it is easy to include keywords that are detected by either command. Simply prefix the keyword with the magic sccs phrase, like this:

     static char *id="@(#) $Id
     : ab.c,v 1.5 1993/10/19 14:57:32 ceder Exp $";

12.3 Avoiding substitution

Keyword substitution has its disadvantages. Sometimes you might want the literal text string ‘$Author$’ to appear inside a file without cvs interpreting it as a keyword and expanding it into something like ‘$Author: ceder $’.

There is unfortunately no way to selectively turn off keyword substitution. You can use ‘-ko’ (see Section 12.4 [Substitution modes], page 77) to turn off keyword substitution entirely.

In many cases you can avoid using keywords in the source, even though they appear in the final product. For example, the source for this manual contains ‘$@asis{}Author$’ whenever the text ‘$Author$’ should appear. In nroff and troff you can embed the null-character \& inside the keyword for a similar effect.

12.4 Substitution modes

Each file has a stored default substitution mode, and each working directory copy of a file also has a substitution mode. The former is set by the ‘-k’ option to cvs add and cvs admin; the latter is set by the ‘-k’ or ‘-A’ options to cvs checkout or cvs update. cvs diff and cvs rdiff also have ‘-k’ options. For some examples, see Chapter 9 [Binary files], page 61, and Section 5.10 [Merging and keywords], page 47.

The modes available are:

‘-kkv’ Generate keyword strings using the default form, e.g. $Revision: 5.7 $ for the Revision keyword.

‘-kkvl’ Like ‘-kkv’, except that a locker’s name is always inserted if the given revision is currently locked. The locker’s name is only relevant if cvs admin -l is in use.

‘-kk’ Generate only keyword names in keyword strings; omit their values. For example, for the Revision keyword, generate the string $Revision$ instead of $Revision: 5.7 $. This option is useful to ignore differences due to keyword substitution when comparing different revisions of a file (see Section 5.10 [Merging and keywords], page 47).

‘-ko’ Generate the old keyword string, present in the working file just before it was checked in. For example, for the Revision keyword, generate the string $Revision: 1.1 $ instead of $Revision: 5.7 $ if that is how the string appeared when the file was checked in.

‘-kb’ Like ‘-ko’, but also inhibit conversion of line endings between the canonical form in which they are stored in the repository (linefeed only), and the form appropriate to the operating system in use on the client. For systems, like unix, which use linefeed only to terminate lines, this is the same as ‘-ko’. For more information on binary files, see Chapter 9 [Binary files], page 61.

‘-kv’ Generate only keyword values for keyword strings. For example, for the Revision keyword, generate the string 5.7 instead of $Revision: 5.7 $. This can help generate files in programming languages where it is hard to strip keyword delimiters like $Revision: $ from a string. However, further keyword substitution cannot be performed once the keyword names are removed, so this option should be used with care. One often would like to use ‘-kv’ with cvs export—see Section A.12 [export], page 112. But be aware that doesn’t handle an export containing binary files correctly.

12.5 Problems with the $Log$ keyword.

The $Log$ keyword is somewhat controversial. As long as you are working on your development system the information is easily accessible even if you do not use the $Log$ keyword—just do a cvs log. Once you export the file the history information might be useless anyhow.

A more serious concern is that cvs is not good at handling $Log$ entries when a branch is merged onto the main trunk. Conflicts often result from the merging operation.

People also tend to "fix" the log entries in the file (correcting spelling mistakes and maybe even factual errors). If that is done the information from cvs log will not be consistent with the information inside the file. This may or may not be a problem in real life.

It has been suggested that the $Log$ keyword should be inserted last in the file, and not in the files header, if it is to be used at all. That way the long list of change messages will not interfere with everyday source file browsing.

13 Tracking third-party sources

If you modify a program to better fit your site, you probably want to include your modifications when the next release of the program arrives. cvs can help you with this task.

In the terminology used in cvs, the supplier of the program is called a vendor. The unmodified distribution from the vendor is checked in on its own branch, the vendor branch. cvs reserves branch 1.1.1 for this use.

When you modify the source and commit it, your revision will end up on the main trunk. When a new release is made by the vendor, you commit it on the vendor branch and copy the modifications onto the main trunk.

Use the import command to create and update the vendor branch. When you import a new file, the vendor branch is made the ‘head’ revision, so anyone that checks out a copy of the file gets that revision. When a local modification is committed it is placed on the main trunk, and made the ‘head’ revision.

13.1 Importing for the first time

Use the import command to check in the sources for the first time. When you use the import command to track third-party sources, the vendor tag and release tags are useful. The vendor tag is a symbolic name for the branch (which is always 1.1.1, unless you use the ‘-b branch’ flag—see Section 13.6 [Multiple vendor branches], page 81.). The release tags are symbolic names for a particular release, such as ‘FSF_0_04’.

Note that import does not change the directory in which you invoke it. In particular, it does not set up that directory as a cvs working directory; if you want to work with the sources import them first and then check them out into a different directory (see Section 1.3.1 [Getting the source], page 4).

Suppose you have the sources to a program called wdiff in a directory ‘wdiff-0.04’, and are going to make private modifications that you want to be able to use even when new releases are made in the future. You start by importing the source to your repository:

     $ cd wdiff-0.04
     $ cvs import -m "Import of FSF v. 0.04" fsf/wdiff FSF_DIST WDIFF_0_04

The vendor tag is named ‘FSF_DIST’ in the above example, and the only release tag assigned is ‘WDIFF_0_04’.

13.2 Updating with the import command

When a new release of the source arrives, you import it into the repository with the same import command that you used to set up the repository in the first place. The only difference is that you specify a different release tag this time:

     $ tar xfz wdiff-0.05.tar.gz
     $ cd wdiff-0.05
     $ cvs import -m "Import of FSF v. 0.05" fsf/wdiff FSF_DIST WDIFF_0_05

WARNING: If you use a release tag that already exists in one of the repository archives, files removed by an import may not be detected.

For files that have not been modified locally, the newly created revision becomes the head revision. If you have made local changes, import will warn you that you must merge the changes into the main trunk, and tell you to use ‘checkout -j’ to do so:

     $ cvs checkout -jFSF_DIST:yesterday -jFSF_DIST wdiff

The above command will check out the latest revision of ‘wdiff’, merging the changes made on the vendor branch ‘FSF_DIST’ since yesterday into the working copy. If any conflicts arise during the merge they should be resolved in the normal way (see Section 10.3 [Conflicts example], page 65). Then, the modified files may be committed.

However, it is much better to use the two release tags rather than using a date on the branch as suggested above:

     $ cvs checkout -jWDIFF_0_04 -jWDIFF_0_05 wdiff

The reason this is better is that using a date, as suggested above, assumes that you do not import more than one release of a product per day. More importantly, using the release tags allows cvs to detect files that were removed between the two vendor releases and mark them for removal. Since import has no way to detect removed files, you should do a merge like this even if import doesn’t tell you to.

13.3 Reverting to the latest vendor release

You can also revert local changes completely and return to the latest vendor release by changing the ‘head’ revision back to the vendor branch on all files. For example, if you have a checked-out copy of the sources in ‘~/work.d/wdiff’, and you want to revert to the vendor’s version for all the files in that directory, you would type:

     $ cd ~/work.d/wdiff
     $ cvs admin -bFSF_DIST .

You must specify the ‘-bFSF_DIST’ without any space after the ‘-b’. See Section A.7.1 [admin options], page 94.

13.4 How to handle binary files with cvs import

Use the ‘-k’ wrapper option to tell import which files are binary. See Section C.2 [Wrappers], page 142.

13.5 How to handle keyword substitution with cvs import

The sources which you are importing may contain keywords (see Chapter 12 [Keyword substitution], page 75). For example, the vendor may use cvs or some other system which uses similar keyword expansion syntax. If you just import the files in the default fashion, then the keyword expansions supplied by the vendor will be replaced by keyword expansions supplied by your own copy of cvs. It may be more convenient to maintain the expansions supplied by the vendor, so that this information can supply information about the sources that you imported from the vendor.

To maintain the keyword expansions supplied by the vendor, supply the ‘-ko’ option to cvs import the first time you import the file. This will turn off keyword expansion for that file entirely, so if you want to be more selective you’ll have to think about what you want and use the ‘-k’ option to cvs update or cvs admin as appropriate.

13.6 Multiple vendor branches

All the examples so far assume that there is only one vendor from which you are getting sources. In some situations you might get sources from a variety of places. For example, suppose that you are dealing with a project where many different people and teams are modifying the software. There are a variety of ways to handle this, but in some cases you have a bunch of source trees lying around and what you want to do more than anything else is just to all put them in cvs so that you at least have them in one place.

For handling situations in which there may be more than one vendor, you may specify the ‘-b’ option to cvs import. It takes as an argument the vendor branch to import to. The default is ‘-b 1.1.1’.

For example, suppose that there are two teams, the red team and the blue team, that are sending you sources. You want to import the red team’s efforts to branch 1.1.1 and use the vendor tag RED. You want to import the blue team’s efforts to branch 1.1.3 and use the vendor tag BLUE. So the commands you might use are:

     $ cvs import dir RED RED_1-0
     $ cvs import -b 1.1.3 dir BLUE BLUE_1-5

Note that if your vendor tag does not match your ‘-b’ option, cvs will not detect this case! For example,

     $ cvs import -b 1.1.3 dir RED RED_1-0

Be careful; this kind of mismatch is sure to sow confusion or worse. I can’t think of a useful purpose for the ability to specify a mismatch here, but if you discover such a use, don’t. cvs is likely to make this an error in some future release.

82 CVS—Concurrent Versions System v1.11.23

14 How your build system interacts with CVS

As mentioned in the introduction, cvs does not contain software for building your software from source code. This section describes how various aspects of your build system might interact with cvs.

One common question, especially from people who are accustomed to rcs, is how to make their build get an up to date copy of the sources. The answer to this with cvs is two-fold. First of all, since cvs itself can recurse through directories, there is no need to modify your ‘Makefile’ (or whatever configuration file your build tool uses) to make sure each file is up to date. Instead, just use two commands, first cvs -q update and then make or whatever the command is to invoke your build tool. Secondly, you do not necessarily want to get a copy of a change someone else made until you have finished your own work. One suggested approach is to first update your sources, then implement, build and test the change you were thinking of, and then commit your sources (updating first if necessary). By periodically (in between changes, using the approach just described) updating your entire tree, you ensure that your sources are sufficiently up to date.

One common need is to record which versions of which source files went into a particular build. This kind of functionality is sometimes called bill of materials or something similar. The best way to do this with cvs is to use the tag command to record which versions went into a given build (see Section 4.4 [Tags], page 34).

Using cvs in the most straightforward manner possible, each developer will have a copy of the entire source tree which is used in a particular build. If the source tree is small, or if developers are geographically dispersed, this is the preferred solution. In fact one approach for larger projects is to break a project down into smaller separately-compiled subsystems, and arrange a way of releasing them internally so that each developer need check out only those subsystems which they are actively working on.

Another approach is to set up a structure which allows developers to have their own copies of some files, and for other files to access source files from a central location. Many people have come up with some such a system using features such as the symbolic link feature found in many operating systems, or the VPATH feature found in many versions of make. One build tool which is designed to help with this kind of thing is Odin (see

ftp://ftp.cs.colorado.edu/pub/distribs/odin).

84 CVS—Concurrent Versions System v1.11.23

15 Special Files

In normal circumstances, cvs works only with regular files. Every file in a project is assumed to be persistent; it must be possible to open, read and close them; and so on. cvs also ignores file permissions and ownerships, leaving such issues to be resolved by the developer at installation time. In other words, it is not possible to "check in" a device into a repository; if the device file cannot be opened, cvs will refuse to handle it. Files also lose their ownerships and permissions during repository transactions.

86 CVS—Concurrent Versions System v1.11.23

Appendix A Guide to CVS commands

This appendix describes the overall structure of cvs commands, and describes some commands in detail (others are described elsewhere; for a quick reference to cvs commands, see Appendix B [Invoking CVS], page 127).

A.1 Overall structure of CVS commands

The overall format of all cvs commands is:

     cvs [ cvs_options ] cvs_command [ command_options ] [ command_args ]

cvs The name of the cvs program.

cvs_options

Some options that affect all sub-commands of cvs. These are described below.

cvs_command

One of several different sub-commands. Some of the commands have aliases that can be used instead; those aliases are noted in the reference manual for that command. There are only two situations where you may omit ‘cvs_command’: ‘cvs -H’ elicits a list of available commands, and ‘cvs -v’ displays version information on cvs itself.

command_options

Options that are specific for the command.

command_args

Arguments to the commands.

There is unfortunately some confusion between cvs_options and command_options. When given as a cvs_option, some options only affect some of the commands. When given as a command_option it may have a different meaning, and be accepted by more commands. In other words, do not take the above categorization too seriously. Look at the documentation instead.

A.2 CVS’s exit status

cvs can indicate to the calling environment whether it succeeded or failed by setting its exit status. The exact way of testing the exit status will vary from one operating system to another. For example in a unix shell script the ‘$?’ variable will be 0 if the last command returned a successful exit status, or greater than 0 if the exit status indicated failure.

If cvs is successful, it returns a successful status; if there is an error, it prints an error message and returns a failure status. The one exception to this is the cvs diff command. It will return a successful status if it found no differences, or a failure status if there were differences or if there was an error. Because this behavior provides no good way to detect errors, in the future it is possible that cvs diff will be changed to behave like the other cvs commands.

A.3 Default options and the ~/.cvsrc file

There are some command_options that are used so often that you might have set up an alias or some other means to make sure you always specify that option. One example (the one that drove the implementation of the ‘.cvsrc’ support, actually) is that many people find the default output of the ‘diff’ command to be very hard to read, and that either context diffs or unidiffs are much easier to understand.

The ‘~/.cvsrc’ file is a way that you can add default options to cvs_commands within cvs, instead of relying on aliases or other shell scripts.

The format of the ‘~/.cvsrc’ file is simple. The file is searched for a line that begins with the same name as the cvs_command being executed. If a match is found, then the remainder of the line is split up (at whitespace characters) into separate options and added to the command arguments before any options from the command line.

If a command has two names (e.g., checkout and co), the official name, not necessarily the one used on the command line, will be used to match against the file. So if this is the contents of the user’s ‘~/.cvsrc’ file:

     log -N
     diff -uN
     rdiff -u
     update -Pd
     checkout -P
     release -d

the command ‘cvs checkout foo’ would have the ‘-P’ option added to the arguments, as well as ‘cvs co foo’.

With the example file above, the output from ‘cvs diff foobar’ will be in unidiff format. ‘cvs diff -c foobar’ will provide context diffs, as usual. Getting "old" format diffs would be slightly more complicated, because diff doesn’t have an option to specify use of the "old" format, so you would need ‘cvs -f diff foobar’.

In place of the command name you can use cvs to specify global options (see Section A.4 [Global options], page 88). For example the following line in ‘.cvsrc’

     cvs -z6

causes cvs to use compression level 6.

A.4 Global options

The available ‘cvs_options’ (that are given to the left of ‘cvs_command’) are:

--allow-root=rootdir

Specify legal cvsroot directory. See Section 2.9.3.1 [Password authentication server], page 21.

-a Authenticate all communication between the client and the server. Only has an effect on the cvs client. As of this writing, this is only implemented when using a GSSAPI connection (see Section 2.9.4 [GSSAPI authenticated], page 26). Authentication prevents certain sorts of attacks involving hijacking the active tcp connection. Enabling authentication does not enable encryption.

-b bindir

In cvs 1.9.18 and older, this specified that rcs programs are in the bindir directory. Current versions of cvs do not run rcs programs; for compatibility this option is accepted, but it does nothing.

-T tempdir

Use tempdir as the directory where temporary files are located. Overrides the setting of the $TMPDIR environment variable and any precompiled directory. This parameter should be specified as an absolute pathname. (When running client/server, ‘-T’ affects only the local process; specifying ‘-T’ for the client has no effect on the server and vice versa.)

-d cvs_root_directory

Use cvs root directory as the root directory pathname of the repository. Overrides the setting of the $CVSROOT environment variable. See Chapter 2 [Repository], page 7.

-e editor

Use editor to enter revision log information. Overrides the setting of the $CVSEDITOR and $EDITOR environment variables. For more information, see Section 1.3.2 [Committing your changes], page 4.

-f Do not read the ‘~/.cvsrc’ file. This option is most often used because of the non-orthogonality of the cvs option set. For example, the ‘cvs log’ option ‘-N’ (turn off display of tag names) does not have a corresponding option to turn the display on. So if you have ‘-N’ in the ‘~/.cvsrc’ entry for ‘log’, you may need to use ‘-f’ to show the tag names.

-H

--help Display usage information about the specified ‘cvs_command’ (but do not actually execute the command). If you don’t specify a command name, ‘cvs -H’ displays overall help for cvs, including a list of other help options.

-n Do not change any files. Attempt to execute the ‘cvs_command’, but only to issue reports; do not remove, update, or merge any existing files, or create any new files. Note that cvs will not necessarily produce exactly the same output as without ‘-n’. In some cases the output will be the same, but in other cases cvs will skip some of the processing that would have been required to produce the exact same output.

-Q Cause the command to be really quiet; the command will only generate output for serious problems.

-q Cause the command to be somewhat quiet; informational messages, such as reports of recursion through subdirectories, are suppressed.

-r Make new working files read-only. Same effect as if the $CVSREAD environment variable is set (see Appendix D [Environment variables], page 155). The default is to make working files writable, unless watches are on (see Section 10.6 [Watches], page 68).

-s variable=value

Set a user variable (see Section C.10 [Variables], page 152).

-t Trace program execution; display messages showing the steps of cvs activity. Particularly useful with ‘-n’ to explore the potential impact of an unfamiliar command.

-v
--version

Display version and copyright information for cvs.

-w Make new working files read-write. Overrides the setting of the $CVSREAD environment variable. Files are created read-write by default, unless $CVSREAD is set or ‘-r’ is given.

-x Encrypt all communication between the client and the server. Only has an effect on the cvs client. As of this writing, this is only implemented when using a GSSAPI connection (see Section 2.9.4 [GSSAPI authenticated], page 26) or a Kerberos connection (see Section 2.9.5 [Kerberos authenticated], page 26). Enabling encryption implies that message traffic is also authenticated. Encryption support is not available by default; it must be enabled using a special configure option, ‘--enable-encryption’, when you build cvs.

-z gzip-level

Set the compression level. Valid levels are 1 (high speed, low compression) to 9 (low speed, high compression), or 0 to disable compression (the default). Only has an effect on the cvs client.

A.5 Common command options

This section describes the ‘command_options’ that are available across several cvs commands. These options are always given to the right of ‘cvs_command’. Not all commands support all of these options; each option is only supported for commands where it makes sense. However, when a command has one of these options you can almost always count on the same behavior of the option as in other commands. (Other command options, which are listed with the individual commands, may have different behavior from one cvs command to the other).

The ‘history’ command is an exception; it supports many options that conflict even with these standard options.

-D date_spec

Use the most recent revision no later than date spec. date spec is a single argument, a date description specifying a date in the past. The specification is sticky when you use it to make a private copy of a source file; that is, when you get a working file using ‘-D’, cvs records the date you specified, so that further updates in the same directory will use the same date (for more information on sticky tags/dates, see Section 4.9 [Sticky tags], page 38).

‘-D’ is available with the annotate, checkout, diff, export, history, rdiff,

rtag, and update commands. (The history command uses this option in a slightly different way; see Section A.13.1 [history options], page 113).

A wide variety of date formats are supported by cvs. The most standard ones are ISO8601 (from the International Standards Organization) and the Internet e-mail standard (specified in RFC822 as amended by RFC1123). ISO8601 dates have many variants but a few examples are:

     1972-09-24
     1972-09-24 20:05

There are a lot more ISO8601 date formats, and cvs accepts many of them, but you probably don’t want to hear the whole long story :-). In addition to the dates allowed in Internet e-mail itself, cvs also allows some of the fields to be omitted. For example:

     24 Sep 1972 20:05
     24 Sep

The date is interpreted as being in the local timezone, unless a specific timezone is specified. These two date formats are preferred. However, cvs currently accepts a wide variety of other date formats. They are intentionally not documented here in any detail, and future versions of cvs might not accept all of them. One such format is month/day/year. This may confuse people who are accustomed to having the month and day in the other order; ‘1/4/96’ is January 4, not April 1. Remember to quote the argument to the ‘-D’ flag so that your shell doesn’t interpret spaces as argument separators. A command using the ‘-D’ flag can look like this:

     $ cvs diff -D "1 hour ago" cvs.texinfo

-f When you specify a particular date or tag to cvs commands, they normally ignore files that do not contain the tag (or did not exist prior to the date) that you specified. Use the ‘-f’ option if you want files retrieved even when there is no match for the tag or date. (The most recent revision of the file will be used). Note that even with ‘-f’, a tag that you specify must exist (that is, in some file, not necessary in every file). This is so that cvs will continue to give an error if you mistype a tag name. ‘-f’ is available with these commands: annotate, checkout, export, rdiff,

rtag, and update.

WARNING: The commit and remove commands also have a ‘-f’ option, but it has a different behavior for those commands. See Section A.10.1 [commit options], page 102, and Section 7.2 [Removing files], page 54.

-k kflag Alter the default processing of keywords. See Chapter 12 [Keyword substitution], page 75, for the meaning of kflag. Your kflag specification is sticky when you use it to create a private copy of a source file; that is, when you use this option with the checkout or update commands, cvs associates your selected kflag with the file, and continues to use it with future update commands on the same file until you specify otherwise.

The ‘-k’ option is available with the add, checkout, diff, rdiff, import and update commands.

-l Local; run only in current working directory, rather than recursing through subdirectories. Available with the following commands: annotate, checkout, commit, diff,

edit, editors, export, log, rdiff, remove, rtag, status, tag, unedit,
update, watch, and watchers.
-m message

Use message as log information, instead of invoking an editor. Available with the following commands: add, commit and import.

-n Do not run any tag program. (A program can be specified to run in the modules database (see Section C.1 [modules], page 139); this option bypasses it). This is not the same as the ‘cvs -n’ program option, which you can specify to the left of a cvs command! Available with the checkout, export, and rtag commands.

-P Prune empty directories. See Section 7.3 [Removing directories], page 55.

-p Pipe the files retrieved from the repository to standard output, rather than writing them in the current directory. Available with the checkout and update commands.

-R Process directories recursively. This is on by default.

Available with the following commands: annotate, checkout, commit, diff,

edit, editors, export, rdiff, remove, rtag, status, tag, unedit, update,
watch, and watchers.

-r tag Use the revision specified by the tag argument instead of the default head revision. As well as arbitrary tags defined with the tag or rtag command, two special tags are always available: ‘HEAD’ refers to the most recent version available in the repository, and ‘BASE’ refers to the revision you last checked out into the current working directory. The tag specification is sticky when you use this with checkout or update to make your own copy of a file: cvs remembers the tag and continues to use it on future update commands, until you specify otherwise (for more information on sticky tags/dates, see Section 4.9 [Sticky tags], page 38). The tag can be either a symbolic or numeric tag, as described in Section 4.4 [Tags], page 34, or the name of a branch, as described in Chapter 5 [Branching and merging], page 41. When a command expects a specific revision, the name of a branch is interpreted as the most recent revision on that branch. Specifying the ‘-q’ global option along with the ‘-r’ command option is often useful, to suppress the warning messages when the rcs file does not contain the specified tag. This is not the same as the overall ‘cvs -r’ option, which you can specify to the left of a cvs command!

‘-r’ is available with the annotate, checkout, commit, diff, history, export,
rdiff, rtag, and update commands.

-W Specify file names that should be filtered. You can use this option repeatedly. The spec can be a file name pattern of the same type that you can specify in the ‘.cvswrappers’ file. Available with the following commands: import, and

          update.

A.6 add—Add files and directories to the repository

• Synopsis: add [-k rcs-kflag] [-m message] files...

• Requires: repository, working directory.

• Changes: repository, working directory.

The add command is used to present new files and directories for addition into the cvs repository. When add is used on a directory, a new directory is created in the repository immediately. When used on a file, only the working directory is updated. Changes to the repository are not made until the commit command is used on the newly added file.

The add command also resurrects files that have been previously removed. This can be done before or after the commit command is used to finalize the removal of files. Resurrected files are restored into the working directory at the time the add command is executed.

A.6.1 add options

These standard options are supported by add (see Section A.5 [Common options], page 90, for a complete description of them):

-k kflag Process keywords according to kflag. See Chapter 12 [Keyword substitution], page 75. This option is sticky; future updates of this file in this working directory will use the same kflag. The status command can be viewed to see the sticky options. For more information on the status command, See Appendix B [Invoking CVS], page 127.

-m message

Use message as the log message, instead of invoking an editor.

A.6.2 add examples

A.6.2.1 Adding a directory

     $ mkdir doc
     $ cvs add doc
     Directory /path/to/repository/doc added to the repository

A.6.2.2 Adding a file

     $ >TODO
     $ cvs add TODO
     cvs add: scheduling file `TODO' for addition
     cvs add: use 'cvs commit' to add this file permanently

A.6.2.3 Undoing a remove command

     $ rm -f makefile

     $ cvs remove makefile
     cvs remove: scheduling `makefile' for removal
     cvs remove: use 'cvs commit' to remove this file permanently
     $ cvs add makefile
     U makefile
     cvs add: makefile, version 1.2, resurrected

A.7 admin—Administration

• Requires: repository, working directory.

• Changes: repository.

• Synonym: rcs

This is the cvs interface to assorted administrative facilities. Some of them have questionable usefulness for cvs but exist for historical purposes. Some of the questionable options are likely to disappear in the future. This command does work recursively, so extreme care should be used.

On unix, if there is a group named cvsadmin, only members of that group can run cvs admin (except for the cvs admin -k command, which can be run by anybody). This group should exist on the server, or any system running the non-client/server cvs. To disallow cvs admin for all users, create a group with no users in it. On NT, the cvsadmin feature does not exist and all users can run cvs admin.

A.7.1 admin options

Some of these options have questionable usefulness for cvs but exist for historical purposes. Some even make it impossible to use cvs until you undo the effect!

-Aoldfile

Might not work together with cvs. Append the access list of oldfile to the access list of the rcs file.

-alogins Might not work together with cvs. Append the login names appearing in the comma-separated list logins to the access list of the rcs file.

-b[rev] Set the default branch to rev. In cvs, you normally do not manipulate default branches; sticky tags (see Section 4.9 [Sticky tags], page 38) are a better way to decide which branch you want to work on. There is one reason to run cvs admin -b: to revert to the vendor’s version when using vendor branches (see Section 13.3 [Reverting local changes], page 80). There can be no space between ‘-b’ and its argument.

-cstring Sets the comment leader to string. The comment leader is not used by current versions of cvs or rcs 5.7. Therefore, you can almost surely not worry about it. See Chapter 12 [Keyword substitution], page 75.

-e[logins]

Might not work together with cvs. Erase the login names appearing in the comma-separated list logins from the access list of the RCS file. If logins is omitted, erase the entire access list. There can be no space between ‘-e’ and its argument.

-I Run interactively, even if the standard input is not a terminal. This option does not work with the client/server cvs and is likely to disappear in a future release of cvs.

-i Useless with cvs. This creates and initializes a new rcs file, without depositing a revision. With cvs, add files with the cvs add command (see Section 7.1 [Adding files], page 53).

-ksubst Set the default keyword substitution to subst. See Chapter 12 [Keyword substitution], page 75. Giving an explicit ‘-k’ option to cvs update, cvs export, or cvs checkout overrides this default.

-l[rev] Lock the revision with number rev. If a branch is given, lock the latest revision on that branch. If rev is omitted, lock the latest revision on the default branch. There can be no space between ‘-l’ and its argument. This can be used in conjunction with the ‘rcslock.pl’ script in the ‘contrib’ directory of the cvs source distribution to provide reserved checkouts (where only one user can be editing a given file at a time). See the comments in that file for details (and see the ‘README’ file in that directory for disclaimers about the unsupported nature of contrib). According to comments in that file, locking must set to strict (which is the default).

-L Set locking to strict. Strict locking means that the owner of an RCS file is not exempt from locking for checkin. For use with cvs, strict locking must be set; see the discussion under the ‘-l’ option above.

-mrev:msg

Replace the log message of revision rev with msg.

-Nname[:[rev]]

Act like ‘-n’, except override any previous assignment of name. For use with magic branches, see Section 5.5 [Magic branch numbers], page 44.

-nname[:[rev]]

Associate the symbolic name name with the branch or revision rev. It is normally better to use ‘cvs tag’ or ‘cvs rtag’ instead. Delete the symbolic name if both ‘:’ and rev are omitted; otherwise, print an error message if name is already associated with another number. If rev is symbolic, it is expanded before association. A rev consisting of a branch number followed by a ‘.’ stands for the current latest revision in the branch. A ‘:’ with an empty rev stands for the current latest revision on the default branch, normally the trunk. For example, ‘cvs admin -nname:’ associates name with the current latest revision of all the RCS files; this contrasts with ‘cvs admin -nname:$’ which associates name with the revision numbers extracted from keyword strings in the corresponding working files.

-orange Deletes (outdates) the revisions given by range.

Note that this command can be quite dangerous unless you know exactly what you are doing (for example see the warnings below about how the rev1:rev2 syntax is confusing).

If you are short on disc this option might help you. But think twice before using it—there is no way short of restoring the latest backup to undo this command! If you delete different revisions than you planned, either due to carelessness or (heaven forbid) a cvs bug, there is no opportunity to correct the error before the revisions are deleted. It probably would be a good idea to experiment on a copy of the repository first. Specify range in one of the following ways:

rev1::rev2

Collapse all revisions between rev1 and rev2, so that cvs only stores the differences associated with going from rev1 to rev2, not intermediate steps. For example, after ‘-o 1.3::1.5’ one can retrieve revision 1.3, revision 1.5, or the differences to get from 1.3 to 1.5, but not the revision 1.4, or the differences between 1.3 and 1.4. Other examples: ‘-o 1.3::1.4’ and ‘-o 1.3::1.3’ have no effect, because there are no intermediate revisions to remove.

::rev Collapse revisions between the beginning of the branch containing rev and rev itself. The branchpoint and rev are left intact. For example, ‘-o ::1.3.2.6’ deletes revision 1.3.2.1, revision 1.3.2.5, and everything in between, but leaves 1.3 and 1.3.2.6 intact.

rev:: Collapse revisions between rev and the end of the branch containing rev. Revision rev is left intact but the head revision is deleted.

rev Delete the revision rev. For example, ‘-o 1.3’ is equivalent to ‘-o

1.2::1.4’.
rev1:rev2

Delete the revisions from rev1 to rev2, inclusive, on the same branch. One will not be able to retrieve rev1 or rev2 or any of the revisions in between. For example, the command ‘cvs admin -oR_1_01:R_1_02 .’ is rarely useful. It means to delete revisions up to, and including, the tag R 1 02. But beware! If there are files that have not changed between R 1 02 and R 1 03 the file will have the same numerical revision number assigned to the tags R 1 02 and R 1 03. So not only will it be impossible to retrieve R 1 02; R 1 03 will also have to be restored from the tapes! In most cases you want to specify rev1::rev2 instead.

:rev Delete revisions from the beginning of the branch containing rev up to and including rev.

rev: Delete revisions from revision rev, including rev itself, to the end of the branch containing rev.

None of the revisions to be deleted may have branches or locks. If any of the revisions to be deleted have symbolic names, and one specifies one of the ‘::’ syntaxes, then cvs will give an error and not delete any revisions. If you really want to delete both the symbolic names and the revisions, first delete the symbolic names with cvs tag -d, then run cvs admin -o. If one specifies the non-‘::’ syntaxes, then cvs will delete the revisions but leave the symbolic names pointing to nonexistent revisions. This behavior is preserved for compatibility with previous versions of cvs, but because it isn’t very useful, in the future it may change to be like the ‘::’ case. Due to the way cvs handles branches rev cannot be specified symbolically if it is a branch. See Section 5.5 [Magic branch numbers], page 44 for an explanation. Make sure that no-one has checked out a copy of the revision you outdate. Strange things will happen if he starts to edit it and tries to check it back in. For this reason, this option is not a good way to take back a bogus commit; commit a new revision undoing the bogus change instead (see Section 5.8 [Merging two revisions], page 46).

-q Run quietly; do not print diagnostics.

-sstate[:rev]

Useful with cvs. Set the state attribute of the revision rev to state. If rev is a branch number, assume the latest revision on that branch. If rev is omitted, assume the latest revision on the default branch. Any identifier is acceptable for state. A useful set of states is ‘Exp’ (for experimental), ‘Stab’ (for stable), and ‘Rel’ (for released). By default, the state of a new revision is set to ‘Exp’ when it is created. The state is visible in the output from cvs log (see Section A.15 [log], page 116), and in the ‘$Log$’ and ‘$State$’ keywords (see Chapter 12 [Keyword substitution], page 75). Note that cvs uses the dead state for its own purposes (see Section 2.2.4 [Attic], page 11); to take a file to or from the dead state use commands like cvs remove and cvs add (see Chapter 7 [Adding and removing], page 53), not cvs admin -s.

-t[file] Useful with cvs. Write descriptive text from the contents of the named file into the RCS file, deleting the existing text. The file pathname may not begin with ‘-’. The descriptive text can be seen in the output from ‘cvs log’ (see Section A.15 [log], page 116). There can be no space between ‘-t’ and its argument. If file is omitted, obtain the text from standard input, terminated by end-of-file or by a line containing ‘.’ by itself. Prompt for the text if interaction is possible; see ‘-I’.

-t-string

Similar to ‘-tfile’. Write descriptive text from the string into the rcs file, deleting the existing text. There can be no space between ‘-t’ and its argument.

-U Set locking to non-strict. Non-strict locking means that the owner of a file need not lock a revision for checkin. For use with cvs, strict locking must be set; see the discussion under the ‘-l’ option above.

-u[rev] See the option ‘-l’ above, for a discussion of using this option with cvs. Unlock the revision with number rev. If a branch is given, unlock the latest revision on that branch. If rev is omitted, remove the latest lock held by the caller. Normally, only the locker of a revision may unlock it; somebody else unlocking a revision breaks the lock. This causes the original locker to be sent a commit notification (see Section 10.6.2 [Getting Notified], page 69). There can be no space between ‘-u’ and its argument.

-Vn In previous versions of cvs, this option meant to write an rcs file which would be acceptable to rcs version n, but it is now obsolete and specifying it will produce an error.

-xsuffixes

In previous versions of cvs, this was documented as a way of specifying the names of the rcs files. However, cvs has always required that the rcs files used by cvs end in ‘,v’, so this option has never done anything useful.

A.8 annotate—What revision modified each line of a file?

• Synopsis: annotate [options] files…

• Requires: repository.

• Synonym: blame

• Changes: nothing.

For each file in files, print the head revision of the trunk, together with information on the last modification for each line.

A.8.1 annotate options

These standard options are supported by annotate (see Section A.5 [Common options], page 90 for a complete description of them):

-l Local directory only, no recursion.

-R Process directories recursively.

-f Use head revision if tag/date not found.

-F Annotate binary files.

-r revision

Annotate file as of specified revision/tag.

-D date Annotate file as of specified date.

A.8.2 annotate example

For example:

     $ cvs annotate ssfile
     Annotations for ssfile
     ***************
     1.1          (mary     27-Mar-96): ssfile line 1
     1.2          (joe      28-Mar-96): ssfile line 2

The file ‘ssfile’ currently contains two lines. The ssfile line 1 line was checked in by mary on March 27. Then, on March 28, joe added a line ssfile line 2, without modifying the ssfile line 1 line. This report doesn’t tell you anything about lines which have been deleted or replaced; you need to use cvs diff for that (see Section A.11 [diff], page 103).

The options to cvs annotate are listed in Appendix B [Invoking CVS], page 127, and can be used to select the files and revisions to annotate. The options are described in more detail there and in Section A.5 [Common options], page 90.

A.9 checkout—Check out sources for editing

• Synopsis: checkout [options] modules…

• Requires: repository.

• Changes: working directory.

• Synonyms: co, get

Create or update a working directory containing copies of the source files specified by modules. You must execute checkout before using most of the other cvs commands, since most of them operate on your working directory.

The modules are either symbolic names for some collection of source directories and files, or paths to directories or files in the repository. The symbolic names are defined in the ‘modules’ file. See Section C.1 [modules], page 139.

Depending on the modules you specify, checkout may recursively create directories and populate them with the appropriate source files. You can then edit these source files at any time (regardless of whether other software developers are editing their own copies of the sources); update them to include new changes applied by others to the source repository; or commit your work as a permanent change to the source repository.

Note that checkout is used to create directories. The top-level directory created is always added to the directory where checkout is invoked, and usually has the same name as the specified module. In the case of a module alias, the created sub-directory may have a different name, but you can be sure that it will be a sub-directory, and that checkout will show the relative path leading to each file as it is extracted into your private work area (unless you specify the ‘-Q’ global option).

The files created by checkout are created read-write, unless the ‘-r’ option to cvs (see Section A.4 [Global options], page 88) is specified, the CVSREAD environment variable is specified (see Appendix D [Environment variables], page 155), or a watch is in effect for that file (see Section 10.6 [Watches], page 68).

Note that running checkout on a directory that was already built by a prior checkout is also permitted. This is similar to specifying the ‘-d’ option to the update command in the sense that new directories that have been created in the repository will appear in your work area. However, checkout takes a module name whereas update takes a directory name. Also to use checkout this way it must be run from the top level directory (where you originally ran checkout from), so before you run checkout to update an existing directory, don’t forget to change your directory to the top level directory.

For the output produced by the checkout command, See Section A.19.2 [update output], page 124.

A.9.1 checkout options

These standard options are supported by checkout (see Section A.5 [Common options], page 90 for a complete description of them):

-D date Use the most recent revision no later than date. This option is sticky, and implies ‘-P’. See Section 4.9 [Sticky tags], page 38 for more information on sticky tags/dates.

-f Only useful with the ‘-D date’ or ‘-r tag’ flags. If no matching revision is found, retrieve the most recent revision (instead of ignoring the file).

-k kflag Process keywords according to kflag. See Chapter 12 [Keyword substitution], page 75. This option is sticky; future updates of this file in this working directory will use the same kflag. The status command can be viewed to see the sticky options. See Appendix B [Invoking CVS], page 127 for more information on the status command.

-l Local; run only in current working directory.

-n Do not run any checkout program (as specified with the ‘-o’ option in the modules file; see Section C.1 [modules], page 139).

-P Prune empty directories. See Section 7.5 [Moving directories], page 57.

-p Pipe files to the standard output.

-R Checkout directories recursively. This option is on by default.

-r tag Use revision tag. This option is sticky, and implies ‘-P’. See Section 4.9 [Sticky tags], page 38, for more information on sticky tags/dates.

In addition to those, you can use these special command options with checkout:

-A Reset any sticky tags, dates, or ‘-k’ options. Does not reset sticky ‘-k’ options on modified files. See Section 4.9 [Sticky tags], page 38 for more information on sticky tags/dates.

-c Copy the module file, sorted, to the standard output, instead of creating or modifying any files or directories in your working directory.

-d dir Create a directory called dir for the working files, instead of using the module name. In general, using this flag is equivalent to using ‘mkdir dir; cd dir’ followed by the checkout command without the ‘-d’ flag. There is an important exception, however. It is very convenient when checking out a single item to have the output appear in a directory that doesn’t contain empty intermediate directories. In this case only, cvs tries to “shorten” pathnames to avoid those empty directories. For example, given a module ‘foo’ that contains the file ‘bar.c’, the command ‘cvs co -d dir foo’ will create directory ‘dir’ and place ‘bar.c’ inside. Similarly, given a module ‘bar’ which has subdirectory ‘baz’ wherein there is a file ‘quux.c’, the command ‘cvs co -d dir bar/baz’ will create directory ‘dir’ and place ‘quux.c’ inside. Using the ‘-N’ flag will defeat this behavior. Given the same module definitions above, ‘cvs co -N -d dir foo’ will create directories ‘dir/foo’ and place ‘bar.c’ inside, while ‘cvs co -N -d dir bar/baz’ will create directories ‘dir/bar/baz’ and place ‘quux.c’ inside.

-j tag With two ‘-j’ options, merge changes from the revision specified with the first ‘-j’ option to the revision specified with the second ‘j’ option, into the working directory. With one ‘-j’ option, merge changes from the ancestor revision to the revision specified with the ‘-j’ option, into the working directory. The ancestor revision is the common ancestor of the revision which the working directory is based on, and the revision specified in the ‘-j’ option. In addition, each -j option can contain an optional date specification which, when used with branches, can limit the chosen revision to one within a specific date. An optional date is specified by adding a colon (:) to the tag:

‘-jSymbolic_Tag:Date_Specifier’.

See Chapter 5 [Branching and merging], page 41.

-N Only useful together with ‘-d dir’. With this option, cvs will not “shorten” module paths in your working directory when you check out a single module. See the ‘-d’ flag for examples and a discussion.

-s Like ‘-c’, but include the status of all modules, and sort it by the status string. See Section C.1 [modules], page 139, for info about the ‘-s’ option that is used inside the modules file to set the module status.

A.9.2 checkout examples

Get a copy of the module ‘tc’:

$ cvs checkout tc

Get a copy of the module ‘tc’ as it looked one day ago:

$ cvs checkout -D yesterday tc

A.10 commit—Check files into the repository

• Synopsis: commit [-lRf] [-m ’log message’ | -F file] [-r revision] [files…]

• Requires: working directory, repository.

• Changes: repository.

• Synonym: ci

Use commit when you want to incorporate changes from your working source files into the source repository.

If you don’t specify particular files to commit, all of the files in your working current directory are examined. commit is careful to change in the repository only those files that you have really changed. By default (or if you explicitly specify the ‘-R’ option), files in subdirectories are also examined and committed if they have changed; you can use the ‘-l’ option to limit commit to the current directory only.

commit verifies that the selected files are up to date with the current revisions in the source repository; it will notify you, and exit without committing, if any of the specified files must be made current first with update (see Section A.19 [update], page 122). commit does not call the update command for you, but rather leaves that for you to do when the time is right.

When all is well, an editor is invoked to allow you to enter a log message that will be written to one or more logging programs (see Section C.1 [modules], page 139, and see Section C.4.4 [loginfo], page 148) and placed in the rcs file inside the repository. This log message can be retrieved with the log command; See Section A.15 [log], page 116. You can specify the log message on the command line with the ‘-m message’ option, and thus avoid the editor invocation, or use the ‘-F file’ option to specify that the argument file contains the log message.

A.10.1 commit options

These standard options are supported by commit (see Section A.5 [Common options], page 90 for a complete description of them):

-l Local; run only in current working directory.

-R Commit directories recursively. This is on by default.

-r revision

Commit to revision. revision must be either a branch, or a revision on the main trunk that is higher than any existing revision number (see Section 4.3 [Assigning revisions], page 33). You cannot commit to a specific revision on a branch.

commit also supports these options:

-F file Read the log message from file, instead of invoking an editor.

-f Note that this is not the standard behavior of the ‘-f’ option as defined in Section A.5 [Common options], page 90. Force cvs to commit a new revision even if you haven’t made any changes to the file. If the current revision of file is 1.7, then the following two commands are equivalent:

               $ cvs commit -f file
               $ cvs commit -r 1.8 file

The ‘-f’ option disables recursion (i.e., it implies ‘-l’). To force cvs to commit a new revision for all files in all subdirectories, you must use ‘-f -R’.

-m message

Use message as the log message, instead of invoking an editor.

A.10.2 commit examples

A.10.2.1 Committing to a branch

You can commit to a branch revision (one that has an even number of dots) with the ‘-r’ option. To create a branch revision, use the ‘-b’ option of the rtag or tag commands (see Chapter 5 [Branching and merging], page 41). Then, either checkout or update can be used to base your sources on the newly created branch. From that point on, all commit changes made within these working sources will be automatically added to a branch revision, thereby not disturbing main-line development in any way. For example, if you had to create a patch to the 1.2 version of the product, even though the 2.0 version is already under development, you might do:

     $ cvs rtag -b -r FCS1_2 FCS1_2_Patch product_module
     $ cvs checkout -r FCS1_2_Patch product_module
     $ cd product_module
     [[ hack away ]]
     $ cvs commit

This works automatically since the ‘-r’ option is sticky.

A.10.2.2 Creating the branch after editing

Say you have been working on some extremely experimental software, based on whatever revision you happened to checkout last week. If others in your group would like to work on this software with you, but without disturbing main-line development, you could commit your change to a new branch. Others can then checkout your experimental stuff and utilize the full benefit of cvs conflict resolution. The scenario might look like:

     [[ hacked sources are present ]]
     $ cvs tag -b EXPR1
     $ cvs update -r EXPR1
     $ cvs commit

The update command will make the ‘-r EXPR1’ option sticky on all files. Note that your changes to the files will never be removed by the update command. The commit will automatically commit to the correct branch, because the ‘-r’ is sticky. You could also do like this:

     [[ hacked sources are present ]]
     $ cvs tag -b EXPR1
     $ cvs commit -r EXPR1

but then, only those files that were changed by you will have the ‘-r EXPR1’ sticky flag. If you hack away, and commit without specifying the ‘-r EXPR1’ flag, some files may accidentally end up on the main trunk.

To work with you on the experimental change, others would simply do

     $ cvs checkout -r EXPR1 whatever_module

A.11 diff—Show differences between revisions

• Synopsis: diff [-lR] [-k kflag] [format options] [[-r rev1 | -D date1] [-r rev2 | -D date2]] [files…]

• Requires: working directory, repository.

• Changes: nothing.

The diff command is used to compare different revisions of files. The default action is to compare your working files with the revisions they were based on, and report any differences that are found.

If any file names are given, only those files are compared. If any directories are given, all files under them will be compared.

The exit status for diff is different than for other cvs commands; for details See Section A.2 [Exit status], page 87.

A.11.1 diff options

These standard options are supported by diff (see Section A.5 [Common options], page 90 for a complete description of them):

-D date Use the most recent revision no later than date. See ‘-r’ for how this affects the comparison.

-k kflag Process keywords according to kflag. See Chapter 12 [Keyword substitution], page 75.

-l Local; run only in current working directory.

-R Examine directories recursively. This option is on by default.

-r tag Compare with revision tag. Zero, one or two ‘-r’ options can be present. With no ‘-r’ option, the working file will be compared with the revision it was based on. With one ‘-r’, that revision will be compared to your current working file. With two ‘-r’ options those two revisions will be compared (and your working file will not affect the outcome in any way). One or both ‘-r’ options can be replaced by a ‘-D date’ option, described above.

The following options specify the format of the output. They have the same meaning as in GNU diff. Most options have two equivalent names, one of which is a single letter preceded by ‘-’, and the other of which is a long name preceded by ‘--’.

‘-lines’ Show lines (an integer) lines of context. This option does not specify an output format by itself; it has no effect unless it is combined with ‘-c’ or ‘-u’. This option is obsolete. For proper operation, patch typically needs at least two lines of context.

‘-a’ Treat all files as text and compare them line-by-line, even if they do not seem to be text.

‘-b’ Ignore trailing white space and consider all other sequences of one or more white space characters to be equivalent.

‘-B’ Ignore changes that just insert or delete blank lines.

‘--binary’

Read and write data in binary mode.

‘--brief’ Report only whether the files differ, not the details of the differences.

‘-c’ Use the context output format.

‘-C lines’
‘--context[=lines]’

Use the context output format, showing lines (an integer) lines of context, or three if lines is not given. For proper operation, patch typically needs at least two lines of context.

‘--changed-group-format=format’

Use format to output a line group containing differing lines from both files in if-then-else format. See Section A.11.1.1 [Line group formats], page 107.

‘-d’ Change the algorithm to perhaps find a smaller set of changes. This makes diff slower (sometimes much slower).

‘-e’

‘--ed’ Make output that is a valid ed script.

‘--expand-tabs’

Expand tabs to spaces in the output, to preserve the alignment of tabs in the input files.

‘-f’ Make output that looks vaguely like an ed script but has changes in the order they appear in the file.

‘-F regexp’

In context and unified format, for each hunk of differences, show some of the last preceding line that matches regexp.

‘--forward-ed’

Make output that looks vaguely like an ed script but has changes in the order they appear in the file.

‘-H’ Use heuristics to speed handling of large files that have numerous scattered small changes.

‘--horizon-lines=lines’

Do not discard the last lines lines of the common prefix and the first lines lines of the common suffix.

‘-i’ Ignore changes in case; consider upper- and lower-case letters equivalent.

‘-I regexp’

Ignore changes that just insert or delete lines that match regexp.

‘--ifdef=name’

Make merged if-then-else output using name.

‘--ignore-all-space’

Ignore white space when comparing lines.

‘--ignore-blank-lines’

Ignore changes that just insert or delete blank lines.

‘--ignore-case’

Ignore changes in case; consider upper- and lower-case to be the same.

‘--ignore-matching-lines=regexp’

Ignore changes that just insert or delete lines that match regexp.

‘--ignore-space-change’

Ignore trailing white space and consider all other sequences of one or more white space characters to be equivalent.

‘--initial-tab’

Output a tab rather than a space before the text of a line in normal or context format. This causes the alignment of tabs in the line to look normal.

‘-L label’

Use label instead of the file name in the context format and unified format headers.

‘--label=label’

Use label instead of the file name in the context format and unified format headers.

‘--left-column’

Print only the left column of two common lines in side by side format.

‘--line-format=format’

Use format to output all input lines in if-then-else format. See Section A.11.1.2 [Line formats], page 110.

‘--minimal’

Change the algorithm to perhaps find a smaller set of changes. This makes diff slower (sometimes much slower).

‘-n’ Output RCS-format diffs; like ‘-f’ except that each command specifies the number of lines affected.

‘-N’
‘--new-file’

In directory comparison, if a file is found in only one directory, treat it as present but empty in the other directory.

‘--new-group-format=format’

Use format to output a group of lines taken from just the second file in if-then-else format. See Section A.11.1.1 [Line group formats], page 107.

‘--new-line-format=format’

Use format to output a line taken from just the second file in if-then-else format. See Section A.11.1.2 [Line formats], page 110.

‘--old-group-format=format’

Use format to output a group of lines taken from just the first file in if-then-else format. See Section A.11.1.1 [Line group formats], page 107.

‘--old-line-format=format’

Use format to output a line taken from just the first file in if-then-else format. See Section A.11.1.2 [Line formats], page 110.

‘-p’ Show which C function each change is in.

‘--rcs’ Output RCS-format diffs; like ‘-f’ except that each command specifies the number of lines affected.

‘--report-identical-files’

‘-s’ Report when two files are the same.

‘--show-c-function’

Show which C function each change is in.

‘--show-function-line=regexp’

In context and unified format, for each hunk of differences, show some of the last preceding line that matches regexp.

‘--side-by-side’

Use the side by side output format.

‘--speed-large-files’

Use heuristics to speed handling of large files that have numerous scattered small changes.

‘--suppress-common-lines’

Do not print common lines in side by side format.

‘-t’ Expand tabs to spaces in the output, to preserve the alignment of tabs in the input files.

‘-T’ Output a tab rather than a space before the text of a line in normal or context format. This causes the alignment of tabs in the line to look normal.

‘--text’ Treat all files as text and compare them line-by-line, even if they do not appear to be text.

‘-u’ Use the unified output format.

‘--unchanged-group-format=format’

Use format to output a group of common lines taken from both files in if-then-else format. See Section A.11.1.1 [Line group formats], page 107.

‘--unchanged-line-format=format’

Use format to output a line common to both files in if-then-else format. See Section A.11.1.2 [Line formats], page 110.

‘-U lines’
‘--unified[=lines]’

Use the unified output format, showing lines (an integer) lines of context, or three if lines is not given. For proper operation, patch typically needs at least two lines of context.

‘-w’ Ignore white space when comparing lines.

‘-W columns’
‘--width=columns’

Use an output width of columns in side by side format.

‘-y’ Use the side by side output format.

A.11.1.1 Line group formats

Line group formats let you specify formats suitable for many applications that allow if-then-else input, including programming languages and text formatting languages. A line group format specifies the output format for a contiguous group of similar lines.

For example, the following command compares the TeX file ‘myfile’ with the original version from the repository, and outputs a merged file in which old regions are surrounded by ‘\begin{em}’-‘\end{em}’ lines, and new regions are surrounded by ‘\begin{bf}’-‘\end{bf}’ lines.

cvs diff \
   --old-group-format='\begin{em}
%<\end{em}
' \
   --new-group-format='\begin{bf}
%>\end{bf}
' \
   myfile

The following command is equivalent to the above example, but it is a little more verbose, because it spells out the default line group formats.

cvs diff \
   --old-group-format='\begin{em}
%<\end{em}
' \
   --new-group-format='\begin{bf}
%>\end{bf}
' \
   --unchanged-group-format='%=' \
   --changed-group-format='\begin{em}
%<\end{em}
\begin{bf}
%>\end{bf}
' \
   myfile

Here is a more advanced example, which outputs a diff listing with headers containing line numbers in a “plain English” style.

cvs diff \
   --unchanged-group-format='' \
   --old-group-format='-------- %dn line%(n=1?:s) deleted at %df:
%<' \
   --new-group-format='-------- %dN line%(N=1?:s) added after %de:
%>' \
   --changed-group-format='-------- %dn line%(n=1?:s) changed at %df:
%<-------- to:
%>' \
   myfile

To specify a line group format, use one of the options listed below. You can specify up to four line group formats, one for each kind of line group. You should quote format, because it typically contains shell metacharacters.

‘--old-group-format=format’

These line groups are hunks containing only lines from the first file. The default old group format is the same as the changed group format if it is specified; otherwise it is a format that outputs the line group as-is.

‘--new-group-format=format’

These line groups are hunks containing only lines from the second file. The default new group format is same as the changed group format if it is specified; otherwise it is a format that outputs the line group as-is.

‘--changed-group-format=format’

These line groups are hunks containing lines from both files. The default changed group format is the concatenation of the old and new group formats.

‘--unchanged-group-format=format’

These line groups contain lines common to both files. The default unchanged group format is a format that outputs the line group as-is.

In a line group format, ordinary characters represent themselves; conversion specifications start with ‘%’ and have one of the following forms.

‘%<’ stands for the lines from the first file, including the trailing newline. Each line is formatted according to the old line format (see Section A.11.1.2 [Line formats], page 110).

‘%>’ stands for the lines from the second file, including the trailing newline. Each line is formatted according to the new line format.

‘%=’ stands for the lines common to both files, including the trailing newline. Each line is formatted according to the unchanged line format.

‘%%’ stands for ‘%’.

‘%c'C'’ where C is a single character, stands for C. C may not be a backslash or an apostrophe. For example, ‘%c':'’ stands for a colon, even inside the then-part of an if-then-else format, which a colon would normally terminate.

‘%c'\O'’ where O is a string of 1, 2, or 3 octal digits, stands for the character with octal code O. For example, ‘%c'\0'’ stands for a null character.

‘Fn’ where F is a printf conversion specification and n is one of the following letters, stands for n’s value formatted with F.

‘e’ The line number of the line just before the group in the old file.

‘f’ The line number of the first line in the group in the old file; equals e + 1.

‘l’ The line number of the last line in the group in the old file.

‘m’ The line number of the line just after the group in the old file; equals l + 1.

‘n’ The number of lines in the group in the old file; equals l - f + 1.

‘E, F, L, M, N’

Likewise, for lines in the new file.

The printf conversion specification can be ‘%d’, ‘%o’, ‘%x’, or ‘%X’, specifying decimal, octal, lower case hexadecimal, or upper case hexadecimal output respectively. After the ‘%’ the following options can appear in sequence: a ‘-’ specifying left-justification; an integer specifying the minimum field width; and a period followed by an optional integer specifying the minimum number of digits. For example, ‘%5dN’ prints the number of new lines in the group in a field of width 5 characters, using the printf format "%5d".

‘(A=B?T:E)’

If A equals B then T else E. A and B are each either a decimal constant or a single letter interpreted as above. This format spec is equivalent to T if A’s value equals B’s; otherwise it is equivalent to E.

For example, ‘%(N=0?no:%dN) line%(N=1?:s)’ is equivalent to ‘no lines’ if N

(the number of lines in the group in the new file) is 0, to ‘1 line’ if N is 1, and to ‘%dN lines’ otherwise.

A.11.1.2 Line formats

Line formats control how each line taken from an input file is output as part of a line group in if-then-else format.

For example, the following command outputs text with a one-column change indicator to the left of the text. The first column of output is ‘-’ for deleted lines, ‘|’ for added lines, and a space for unchanged lines. The formats contain newline characters where newlines are desired on output.

cvs diff \
--old-line-format='-%l
' \
--new-line-format='|%l
' \
--unchanged-line-format=' %l
' \
myfile

To specify a line format, use one of the following options. You should quote format, since it often contains shell metacharacters.

‘--old-line-format=format’

formats lines just from the first file.

‘--new-line-format=format’

formats lines just from the second file.

‘--unchanged-line-format=format’

formats lines common to both files.

‘--line-format=format’

formats all lines; in effect, it sets all three above options simultaneously.

In a line format, ordinary characters represent themselves; conversion specifications start with ‘%’ and have one of the following forms.

‘%l’ stands for the contents of the line, not counting its trailing newline (if any). This format ignores whether the line is incomplete.

‘%L’ stands for the contents of the line, including its trailing newline (if any). If a line is incomplete, this format preserves its incompleteness.

‘%%’ stands for ‘%’.

‘%c'C'’ where C is a single character, stands for C. C may not be a backslash or an apostrophe. For example, ‘%c':'’ stands for a colon.

‘%c'\O'’ where O is a string of 1, 2, or 3 octal digits, stands for the character with octal code O. For example, ‘%c'\0'’ stands for a null character.

‘Fn’ where F is a printf conversion specification, stands for the line number formatted with F. For example, ‘%.5dn’ prints the line number using the printf format "%.5d". See Section A.11.1.1 [Line group formats], page 107, for more about printf conversion specifications.

The default line format is ‘%l’ followed by a newline character.

If the input contains tab characters and it is important that they line up on output, you should ensure that ‘%l’ or ‘%L’ in a line format is just after a tab stop (e.g. by preceding ‘%l’ or ‘%L’ with a tab character), or you should use the ‘-t’ or ‘--expand-tabs’ option.

Taken together, the line and line group formats let you specify many different formats. For example, the following command uses a format similar to diff’s normal format. You can tailor this command to get fine control over diff’s output.

     cvs diff \
        --old-line-format='< %l
     ' \
        --new-line-format='> %l
     ' \
        --old-group-format='%df%(f=l?:,%dl)d%dE
     %<' \
        --new-group-format='%dea%dF%(F=L?:,%dL)
     %>' \
        --changed-group-format='%df%(f=l?:,%dl)c%dF%(F=L?:,%dL)
     %<---
     %>' \
        --unchanged-group-format='' \
        myfile

A.11.2 diff examples

The following line produces a Unidiff (‘-u’ flag) between revision 1.14 and 1.19 of ‘backend.c’. Due to the ‘-kk’ flag no keywords are substituted, so differences that only depend on keyword substitution are ignored.

     $ cvs diff -kk -u -r 1.14 -r 1.19 backend.c

Suppose the experimental branch EXPR1 was based on a set of files tagged RELEASE 1 0. To see what has happened on that branch, the following can be used:

     $ cvs diff -r RELEASE_1_0 -r EXPR1

A command like this can be used to produce a context diff between two releases:

     $ cvs diff -c -r RELEASE_1_0 -r RELEASE_1_1 > diffs

If you are maintaining ChangeLogs, a command like the following just before you commit your changes may help you write the ChangeLog entry. All local modifications that have not yet been committed will be printed.

     $ cvs diff -u | less

A.12 export—Export sources from CVS, similar to checkout

• Synopsis: export [-flNnR] [-r rev|-D date] [-k subst] [-d dir] module…

• Requires: repository.

• Changes: current directory.

This command is a variant of checkout; use it when you want a copy of the source for module without the cvs administrative directories. For example, you might use export to prepare source for shipment off-site. This command requires that you specify a date or tag (with ‘-D’ or ‘-r’), so that you can count on reproducing the source you ship to others (and thus it always prunes empty directories).

One often would like to use ‘-kv’ with cvs export. This causes any keywords to be expanded such that an import done at some other site will not lose the keyword revision information. But be aware that doesn’t handle an export containing binary files correctly. Also be aware that after having used ‘-kv’, one can no longer use the ident command (which is part of the rcs suite—see ident(1)) which looks for keyword strings. If you want to be able to use ident you must not use ‘-kv’.

A.12.1 export options

These standard options are supported by export (see Section A.5 [Common options], page 90, for a complete description of them):

-D date Use the most recent revision no later than date.

-f If no matching revision is found, retrieve the most recent revision (instead of ignoring the file).

-l Local; run only in current working directory.

-n Do not run any checkout program.

-R Export directories recursively. This is on by default.

-r tag Use revision tag.

In addition, these options (that are common to checkout and export) are also supported:

-d dir Create a directory called dir for the working files, instead of using the module name. See Section A.9.1 [checkout options], page 99 for complete details on how cvs handles this flag.

-k subst Set keyword expansion mode (see Section 12.4 [Substitution modes], page 77).

-N Only useful together with ‘-d dir’. See Section A.9.1 [checkout options], page 99 for complete details on how cvs handles this flag.

A.13 history—Show status of files and users

• Synopsis: history [-report] [-flags] [-options args] [files…]

• Requires: the file ‘$CVSROOT/CVSROOT/history’

• Changes: nothing. cvs can keep a history file that tracks each use of the checkout, commit, rtag, update,

and release commands. You can use history to display this information in various formats.

Logging must be enabled by creating the file ‘$CVSROOT/CVSROOT/history’. history uses ‘-f’, ‘-l’, ‘-n’, and ‘-p’ in ways that conflict with the normal use inside

cvs (see Section A.5 [Common options], page 90).

A.13.1 history options

Several options (shown above as ‘-report’) control what kind of report is generated: -c Report on each time commit was used (i.e., each time the repository was mod-

ified).

-e Everything (all record types). Equivalent to specifying ‘-x’ with all record types. Of course, ‘-e’ will also include record types which are added in a future version of cvs; if you are writing a script which can only handle certain record types, you’ll want to specify ‘-x’.

-m module

Report on a particular module. (You can meaningfully use ‘-m’ more than once on the command line.)

-o Report on checked-out modules. This is the default report type. -T Report on all tags. -x type Extract a particular set of record types type from the cvs history. The types

are indicated by single letters, which you may specify in combination. Certain commands have a single record type: F release O checkout E export T rtag One of five record types may result from an update: C A merge was necessary but collisions were detected (requiring man-

ual merging).

G A merge was necessary and it succeeded. U A working file was copied from the repository. P A working file was patched to match the repository.

W The working copy of a file was deleted during update (because it was gone from the repository).

One of three record types results from commit:

A A file was added for the first time.

M A file was modified.

R A file was removed.

The options shown as ‘-flags’ constrain or expand the report without requiring option arguments:

-a Show data for all users (the default is to show data only for the user executing

history).

-l Show last modification only.

-w Show only the records for modifications done from the same working directory where history is executing.

The options shown as ‘-options args’ constrain the report based on an argument:

-b str Show data back to a record containing the string str in either the module name, the file name, or the repository path.

-D date Show data since date. This is slightly different from the normal use of ‘-D date’, which selects the newest revision older than date.

-f file Show data for a particular file (you can specify several ‘-f’ options on the same command line). This is equivalent to specifying the file on the command line.

-n module

Show data for a particular module (you can specify several ‘-n’ options on the same command line).

-p repository

Show data for a particular source repository (you can specify several ‘-p’ options on the same command line).

-r rev Show records referring to revisions since the revision or tag named rev appears in individual rcs files. Each rcs file is searched for the revision or tag.

-t tag Show records since tag tag was last added to the history file. This differs from the ‘-r’ flag above in that it reads only the history file, not the rcs files, and is much faster.

-u name Show records for user name.

-z timezone

Show times in the selected records using the specified time zone instead of UTC.

A.14 import—Import sources into CVS, using vendor branches

• Synopsis: import [-options] repository vendortag releasetag…

• Requires: Repository, source distribution directory.

• Changes: repository.

Use import to incorporate an entire source distribution from an outside source (e.g., a source vendor) into your source repository directory. You can use this command both for initial creation of a repository, and for wholesale updates to the module from the outside source. See Chapter 13 [Tracking sources], page 79 for a discussion on this subject.

The repository argument gives a directory name (or a path to a directory) under the cvs root directory for repositories; if the directory did not exist, import creates it.

When you use import for updates to source that has been modified in your source repository (since a prior import), it will notify you of any files that conflict in the two branches of development; use ‘checkout -j’ to reconcile the differences, as import instructs you to do.

If cvs decides a file should be ignored (see Section C.7 [cvsignore], page 150), it does not import it and prints ‘I ’ followed by the filename (see Section A.14.2 [import output], page 116 for a complete description of the output).

If the file ‘$CVSROOT/CVSROOT/cvswrappers’ exists, any file whose names match the specifications in that file will be treated as packages and the appropriate filtering will be performed on the file/directory before being imported. See Section C.2 [Wrappers], page 142.

The outside source is saved in a first-level branch, by default 1.1.1. Updates are leaves of this branch; for example, files from the first imported collection of source will be revision 1.1.1.1, then files from the first imported update will be revision 1.1.1.2, and so on.

At least three arguments are required. repository is needed to identify the collection of source. vendortag is a tag for the entire branch (e.g., for 1.1.1). You must also specify at least one releasetag to uniquely identify the files at the leaves created each time you execute import. The releasetag should be new, not previously existing in the repository file, and uniquely identify the imported release,

Note that import does not change the directory in which you invoke it. In particular, it does not set up that directory as a cvs working directory; if you want to work with the sources import them first and then check them out into a different directory (see Section 1.3.1 [Getting the source], page 4).

A.14.1 import options

This standard option is supported by import (see Section A.5 [Common options], page 90 for a complete description):

-m message

Use message as log information, instead of invoking an editor.

There are the following additional special options.

-b branch

See Section 13.6 [Multiple vendor branches], page 81.

-d Use each file’s modification time as the time of import rather than the current time.

-k subst Indicate the keyword expansion mode desired. This setting will apply to all files created during the import, but not to any files that previously existed in the repository. See Section 12.4 [Substitution modes], page 77 for a list of valid ‘-k’ settings.

-I name Specify file names that should be ignored during import. You can use this option repeatedly. To avoid ignoring any files at all (even those ignored by default), specify ‘-I !’. name can be a file name pattern of the same type that you can specify in the ‘.cvsignore’ file. See Section C.7 [cvsignore], page 150.

-W spec Specify file names that should be filtered during import. You can use this option repeatedly. spec can be a file name pattern of the same type that you can specify in the ‘.cvswrappers’ file. See Section C.2 [Wrappers], page 142.

A.14.2 import output

import keeps you informed of its progress by printing a line for each file, preceded by one character indicating the status of the file:

U file The file already exists in the repository and has not been locally modified; a new revision has been created (if necessary).

N file The file is a new file which has been added to the repository.

C file The file already exists in the repository but has been locally modified; you will have to merge the changes.

I file The file is being ignored (see Section C.7 [cvsignore], page 150).

L file The file is a symbolic link; cvs import ignores symbolic links. People periodically suggest that this behavior should be changed, but if there is a consensus on what it should be changed to, it doesn’t seem to be apparent. (Various options in the ‘modules’ file can be used to recreate symbolic links on checkout, update, etc.; see Section C.1 [modules], page 139.)

A.14.3 import examples

See Chapter 13 [Tracking sources], page 79, and Section 3.1.1 [From files], page 29.

A.15 log—Print out log information for files

• Synopsis: log [options] [files…]

• Requires: repository, working directory.

• Changes: nothing.

Display log information for files. log used to call the rcs utility rlog. Although this is no longer true in the current sources, this history determines the format of the output and the options, which are not quite in the style of the other cvs commands.

The output includes the location of the rcs file, the head revision (the latest revision on the trunk), all symbolic names (tags) and some other things. For each revision, the revision number, the author, the number of lines added/deleted and the log message are printed. All times are displayed in Coordinated Universal Time (UTC). (Other parts of cvs print times in the local timezone).

log uses ‘-R’ in a way that conflicts with the normal use inside cvs (see Section A.5 [Common options], page 90).

A.15.1 log options

By default, log prints all information that is available. All other options restrict the output. Note that the revision selection options (-b, -d, -r, -s, and -w) have no effect, other than possibly causing a search for files in Attic directories, when used in conjunction with the options that restrict the output to only log header fields (-h, -R, and -t) unless the -S option is also specified.

-b Print information about the revisions on the default branch, normally the highest branch on the trunk.

-d dates Print information about revisions with a checkin date/time in the range given by the semicolon-separated list of dates. The date formats accepted are those accepted by the ‘-D’ option to many other cvs commands (see Section A.5 [Common options], page 90). Dates can be combined into ranges as follows:

d1<d2

d2>d1 Select the revisions that were deposited between d1 and d2.

<d

d> Select all revisions dated d or earlier.

d<

>d Select all revisions dated d or later.

d Select the single, latest revision dated d or earlier.

The ‘>’ or ‘<’ characters may be followed by ‘=’ to indicate an inclusive range rather than an exclusive one. Note that the separator is a semicolon (;).

-h Print only the name of the rcs file, name of the file in the working directory, head, default branch, access list, locks, symbolic names, and suffix.

-l Local; run only in current working directory. (Default is to run recursively).

-N Do not print the list of tags for this file. This option can be very useful when your site uses a lot of tags, so rather than "more"’ing over 3 pages of tag information, the log information is presented without tags at all.

-n Print the list of tags for this file. This option can be very useful when your ‘.cvsrc’ file has a ‘log -N’ entry as a way to get a full list of all of the tags.

-R Print only the name of the rcs file.

-rrevisions

Print information about revisions given in the comma-separated list revisions of revisions and ranges. The following table explains the available range formats:

rev1:rev2

Revisions rev1 to rev2 (which must be on the same branch).

rev1::rev2

The same, but excluding rev1.

:rev

::rev Revisions from the beginning of the branch up to and including rev.

rev: Revisions starting with rev to the end of the branch containing rev.

rev:: Revisions starting just after rev to the end of the branch containing rev.

branch An argument that is a branch means all revisions on that branch.

branch1:branch2
branch1::branch2

A range of branches means all revisions on the branches in that range.

branch. The latest revision in branch.

A bare ‘-r’ with no revisions means the latest revision on the default branch, normally the trunk. There can be no space between the ‘-r’ option and its argument.

-S Suppress the header if no revisions are selected.

-s states

Print information about revisions whose state attributes match one of the states given in the comma-separated list states. Individual states may be any text string, though cvs commonly only uses two states, ‘Exp’ and ‘dead’. See Section A.7.1 [admin options], page 94 for more information.

-t Print the same as ‘-h’, plus the descriptive text.

-wlogins Print information about revisions checked in by users with login names appearing in the comma-separated list logins. If logins is omitted, the user’s login is assumed. There can be no space between the ‘-w’ option and its argument.

log prints the intersection of the revisions selected with the options ‘-d’, ‘-s’, and ‘-w’, intersected with the union of the revisions selected by ‘-b’ and ‘-r’.

A.15.2 log examples

Contributed examples are gratefully accepted.

A.16 rdiff—’patch’ format diffs between releases

• rdiff [-flags] [-V vn] [-r t|-D d [-r t2|-D d2]] modules…

• Requires: repository.

• Changes: nothing.

• Synonym: patch

Builds a Larry Wall format patch(1) file between two releases, that can be fed directly into the patch program to bring an old release up-to-date with the new release. (This is one of the few cvs commands that operates directly from the repository, and doesn’t require a prior checkout.) The diff output is sent to the standard output device.

You can specify (using the standard ‘-r’ and ‘-D’ options) any combination of one or two revisions or dates. If only one revision or date is specified, the patch file reflects differences between that revision or date and the current head revisions in the rcs file.

Note that if the software release affected is contained in more than one directory, then it may be necessary to specify the ‘-p’ option to the patch command when patching the old sources, so that patch is able to find the files that are located in other directories.

A.16.1 rdiff options

These standard options are supported by rdiff (see Section A.5 [Common options], page 90 for a complete description of them):

-D date Use the most recent revision no later than date.

-f If no matching revision is found, retrieve the most recent revision (instead of ignoring the file).

-k kflag Process keywords according to kflag. See Chapter 12 [Keyword substitution], page 75.

-l Local; don’t descend subdirectories.

-R Examine directories recursively. This option is on by default.

-r tag Use revision tag.

In addition to the above, these options are available:

-c Use the context diff format. This is the default format.

-s Create a summary change report instead of a patch. The summary includes information about files that were changed or added between the releases. It is sent to the standard output device. This is useful for finding out, for example, which files have changed between two dates or revisions.

-t A diff of the top two revisions is sent to the standard output device. This is most useful for seeing what the last change to a file was.

-u Use the unidiff format for the context diffs. Remember that old versions of the patch program can’t handle the unidiff format, so if you plan to post this patch to the net you should probably not use ‘-u’.

-V vn Expand keywords according to the rules current in rcs version vn (the expansion format changed with rcs version 5). Note that this option is no longer accepted. cvs will always expand keywords the way that rcs version 5 does.

A.16.2 rdiff examples

Suppose you receive mail from foo@example.net asking for an update from release 1.2 to 1.4 of the tc compiler. You have no such patches on hand, but with cvs that can easily be fixed with a command such as this:

     $ cvs rdiff -c -r FOO1_2 -r FOO1_4 tc | \
     > Mail -s 'The patches you asked for' foo@example.net

Suppose you have made release 1.3, and forked a branch called ‘R_1_3fix’ for bug fixes. ‘R_1_3_1’ corresponds to release 1.3.1, which was made some time ago. Now, you want to see how much development has been done on the branch. This command can be used:

     $ cvs patch -s -r R_1_3_1 -r R_1_3fix module-name
     cvs rdiff: Diffing module-name
     File ChangeLog,v changed from revision 1.52.2.5 to 1.52.2.6
     File foo.c,v changed from revision 1.52.2.3 to 1.52.2.4
     File bar.h,v changed from revision 1.29.2.1 to 1.2

A.17 release—Indicate that a Module is no longer in use

• release [-d] directories…

• Requires: Working directory.

• Changes: Working directory, history log.

This command is meant to safely cancel the effect of ‘cvs checkout’. Since cvs doesn’t lock files, it isn’t strictly necessary to use this command. You can always simply delete your working directory, if you like; but you risk losing changes you may have forgotten, and you leave no trace in the cvs history file (see Section C.9 [history file], page 152) that you’ve abandoned your checkout.

Use ‘cvs release’ to avoid these problems. This command checks that no uncommitted changes are present; that you are executing it from immediately above a cvs working directory; and that the repository recorded for your files is the same as the repository defined in the module database.

If all these conditions are true, ‘cvs release’ leaves a record of its execution (attesting to your intentionally abandoning your checkout) in the cvs history log.

A.17.1 release options

The release command supports one command option:

-d Delete your working copy of the file if the release succeeds. If this flag is not given your files will remain in your working directory. WARNING: The release command deletes all directories and files recursively. This has the very serious side-effect that any directory created inside checkedout sources, and not added to the repository (using the add command; see Section 7.1 [Adding files], page 53) will be silently deleted—even if it is nonempty!

A.17.2 release output

Before release releases your sources it will print a one-line message for any file that is not up-to-date.

U file

P file There exists a newer revision of this file in the repository, and you have not modified your local copy of the file (‘U’ and ‘P’ mean the same thing).

A file The file has been added to your private copy of the sources, but has not yet been committed to the repository. If you delete your copy of the sources this file will be lost.

R file The file has been removed from your private copy of the sources, but has not yet been removed from the repository, since you have not yet committed the removal. See Section A.10 [commit], page 101.

M file The file is modified in your working directory. There might also be a newer revision inside the repository.

? file file is in your working directory, but does not correspond to anything in the source repository, and is not in the list of files for cvs to ignore (see the description of the ‘-I’ option, and see Section C.7 [cvsignore], page 150). If you remove your working sources, this file will be lost.

A.17.3 release examples

Release the ‘tc’ directory, and delete your local working copy of the files.

$ cd .. # You must stand immediately above the

# sources when you issue ‘cvs release’.

     $ cvs release -d tc
     You have [0] altered files in this repository.
     Are you sure you want to release (and delete) directory `tc': y
     $

A.18 remove—Remove files from active use

• Synopsis: remove [-flR] [files...]

• Requires: repository, working directory.

• Changes: working directory.

The remove command is used to remove unwanted files from active use. The user normally deletes the files from the working directory prior to invocation of the remove command. Only the working directory is updated. Changes to the repository are not made until the commit command is run.

The remove command does not delete files from from the repository. cvs keeps all historical data in the repository so that it is possible to reconstruct previous states of the projects under revision control.

To undo cvs remove or to resurrect files that were previously removed, See Section A.6 [add], page 93.

A.18.1 remove options

These standard options are supported by remove (see Section A.5 [Common options], page 90 for a complete description of them):

-l Local; run only in current working directory. See Chapter 6 [Recursive behavior], page 51.

-R Process directories recursively. See Chapter 6 [Recursive behavior], page 51.

In addition, these options are also supported:

-f Note that this is not the standard behavior of the ‘-f’ option as defined in Section A.5 [Common options], page 90. Delete files before removing them. Entire directory hierarchies are easily removed using ‘-f’, but take note that it is not as easy to resurrect directory hierarchies as it is to remove them.

A.18.2 remove examples

A.18.2.1 Removing a file

$ cvs remove remove.me
cvs remove: file `remove.me' still in working directory
cvs remove: 1 file exists; remove it first
$ rm -f remove.me
$ cvs remove remove.me
cvs remove: scheduling `remove.me' for removal
cvs remove: use 'cvs commit' to remove this file permanently
$ ls remove.it
remove.it
$ cvs remove -f remove.it
cvs remove: scheduling `remove.it' for removal
cvs remove: use 'cvs commit' to remove this file permanently

A.18.2.2 Removing entire directories

$ tree -d a
a
|-- CVS
`-- b
    `-- CVS
3 directories
$ cvs remove -f a
cvs remove: Removing a
cvs remove: Removing a/b
cvs remove: scheduling `a/b/c' for removal
cvs remove: use 'cvs commit' to remove this file permanently

A.19 update—Bring work tree in sync with repository

• update [-ACdflPpR] [-I name] [-j rev [-j rev]] [-k kflag] [-r tag|-D date] [-W spec] files… • Requires: repository, working directory.

• Changes: working directory.

After you’ve run checkout to create your private copy of source from the common repository, other developers will continue changing the central source. From time to time, when it is convenient in your development process, you can use the update command from within your working directory to reconcile your work with any revisions applied to the source repository since your last checkout or update.

A.19.1 update options

These standard options are available with update (see Section A.5 [Common options], page 90 for a complete description of them):

-D date Use the most recent revision no later than date. This option is sticky, and implies ‘-P’. See Section 4.9 [Sticky tags], page 38 for more information on sticky tags/dates.

-f Only useful with the ‘-D date’ or ‘-r tag’ flags. If no matching revision is found, retrieve the most recent revision (instead of ignoring the file).

-k kflag Process keywords according to kflag. See Chapter 12 [Keyword substitution], page 75. This option is sticky; future updates of this file in this working directory will use the same kflag. The status command can be viewed to see the sticky options. See Appendix B [Invoking CVS], page 127 for more information on the status command.

-l Local; run only in current working directory. See Chapter 6 [Recursive behavior], page 51.

-P Prune empty directories. See Section 7.5 [Moving directories], page 57.

-p Pipe files to the standard output.

-R Update directories recursively (default). See Chapter 6 [Recursive behavior], page 51.

-r rev Retrieve revision/tag rev. This option is sticky, and implies ‘-P’. See Section 4.9 [Sticky tags], page 38, for more information on sticky tags/dates.

These special options are also available with update.

-A Reset any sticky tags, dates, or ‘-k’ options. Does not reset sticky ‘-k’ options on modified files. See Section 4.9 [Sticky tags], page 38 for more information on sticky tags/dates.

-C Overwrite locally modified files with clean copies from the repository (the modified file is saved in ‘.#file.revision’, however).

-d Create any directories that exist in the repository if they’re missing from the working directory. Normally, update acts only on directories and files that were already enrolled in your working directory. This is useful for updating directories that were created in the repository since the initial checkout; but it has an unfortunate side effect. If you deliberately avoided certain directories in the repository when you created your working directory (either through use of a module name or by listing explicitly the files and directories you wanted on the command line), then updating with ‘-d’ will create those directories, which may not be what you want.

-I name Ignore files whose names match name (in your working directory) during the update. You can specify ‘-I’ more than once on the command line to specify several files to ignore. Use ‘-I !’ to avoid ignoring any files at all. See Section C.7 [cvsignore], page 150 for other ways to make cvs ignore some files.

-Wspec Specify file names that should be filtered during update. You can use this option repeatedly. spec can be a file name pattern of the same type that you can specify in the ‘.cvswrappers’ file. See Section C.2 [Wrappers], page 142.

-jrevision

With two ‘-j’ options, merge changes from the revision specified with the first ‘-j’ option to the revision specified with the second ‘j’ option, into the working directory. With one ‘-j’ option, merge changes from the ancestor revision to the revision specified with the ‘-j’ option, into the working directory. The ancestor revision is the common ancestor of the revision which the working directory is based on, and the revision specified in the ‘-j’ option. Note that using a single ‘-j tagname’ option rather than ‘-j branchname’ to merge changes from a branch will often not remove files which were removed on the branch. See Section 5.9 [Merging adds and removals], page 47 for more information. In addition, each ‘-j’ option can contain an optional date specification which, when used with branches, can limit the chosen revision to one within a specific date. An optional date is specified by adding a colon (:) to the tag:

‘-jSymbolic_Tag:Date_Specifier’.

See Chapter 5 [Branching and merging], page 41.

A.19.2 update output

update and checkout keep you informed of their progress by printing a line for each file, preceded by one character indicating the status of the file:

U file The file was brought up to date with respect to the repository. This is done for any file that exists in the repository but not in your working directory, and for files that you haven’t changed but are not the most recent versions available in the repository.

P file Like ‘U’, but the cvs server sends a patch instead of an entire file. This accomplishes the same thing as ‘U’ using less bandwidth.

A file The file has been added to your private copy of the sources, and will be added to the source repository when you run commit on the file. This is a reminder to you that the file needs to be committed.

R file The file has been removed from your private copy of the sources, and will be removed from the source repository when you run commit on the file. This is a reminder to you that the file needs to be committed.

M file The file is modified in your working directory.

‘M’ can indicate one of two states for a file you’re working on: either there were no modifications to the same file in the repository, so that your file remains as you last saw it; or there were modifications in the repository as well as in your copy, but they were merged successfully, without conflict, in your working directory. cvs will print some messages if it merges your work, and a backup copy of your working file (as it looked before you ran update) will be made. The exact name of that file is printed while update runs.

C file A conflict was detected while trying to merge your changes to file with changes from the source repository. file (the copy in your working directory) is now the result of attempting to merge the two revisions; an unmodified copy of your file is also in your working directory, with the name ‘.#file.revision’ where revision is the revision that your modified file started from. Resolve the conflict as described in Section 10.3 [Conflicts example], page 65. (Note that some systems automatically purge files that begin with ‘.#’ if they have not been accessed for a few days. If you intend to keep a copy of your original file, it is a very good idea to rename it.) Under vms, the file name starts with ‘__’ rather than ‘.#’.

? file file is in your working directory, but does not correspond to anything in the source repository, and is not in the list of files for cvs to ignore (see the description of the ‘-I’ option, and see Section C.7 [cvsignore], page 150).

126 CVS—Concurrent Versions System v1.11.23

Appendix B Quick reference to CVS commands

This appendix describes how to invoke cvs, with references to where each command or feature is described in detail. For other references run the cvs --help command, or see [Index], page 173.

A cvs command looks like:

     cvs [ global_options ] command [ command_options ] [ command_args ]

Global options:

--allow-root=rootdir

Specify legal cvsroot directory (server only) (not in cvs 1.9 and older). See Section 2.9.3.1 [Password authentication server], page 21.

-a Authenticate all communication (client only) (not in cvs 1.9 and older). See

Section A.4 [Global options], page 88.

-b Specify RCS location (cvs 1.9 and older). See Section A.4 [Global options],

page 88.

-d root Specify the cvsroot. See Chapter 2 [Repository], page 7.

-e editor

Edit messages with editor. See Section 1.3.2 [Committing your changes], page 4.

-f Do not read the ‘~/.cvsrc’ file. See Section A.4 [Global options], page 88.

-H

--help Print a help message. See Section A.4 [Global options], page 88. -n Do not change any files. See Section A.4 [Global options], page 88. -Q Be really quiet. See Section A.4 [Global options], page 88. -q Be somewhat quiet. See Section A.4 [Global options], page 88. -r Make new working files read-only. See Section A.4 [Global options], page 88.

-s variable=value

Set a user variable. See Section C.10 [Variables], page 152.

-T tempdir

Put temporary files in tempdir. See Section A.4 [Global options], page 88.

-t Trace cvs execution. See Section A.4 [Global options], page 88.

-v
--version

Display version and copyright information for cvs.

-w Make new working files read-write. See Section A.4 [Global options], page 88. -x Encrypt all communication (client only). See Section A.4 [Global options],

page 88.

-z gzip-level

Set the compression level (client only). See Section A.4 [Global options], page 88.

Keyword expansion modes (see Section 12.4 [Substitution modes], page 77):

-kkv  $Id
: file1,v 1.1 1993/12/09 03:21:13 joe Exp $
-kkvl $Id
: file1,v 1.1 1993/12/09 03:21:13 joe Exp harry $
-kk   $Id
$
-kv   file1,v 1.1 1993/12/09 03:21:13 joe Exp
-ko   no expansion
-kb   no expansion, file is binary

Keywords (see Section 12.1 [Keyword list], page 75):

$Author
: joe $
$Date
: 1993/12/09 03:21:13 $
$Header
: /home/files/file1,v 1.1 1993/12/09 03:21:13 joe Exp harry $
$Id
: file1,v 1.1 1993/12/09 03:21:13 joe Exp harry $
$Locker
: harry $
$Name
: snapshot_1_14 $
$RCSfile
: file1,v $
$Revision
: 1.1 $
$Source
: /home/files/file1,v $
$State
: Exp $
$Log
: file1,v $
Revision 1.1  1993/12/09 03:30:17  joe
Initial revision

Commands, command options, and command arguments:

add [options] [files...]

Add a new file/directory. See Section 7.1 [Adding files], page 53.

-k kflag Set keyword expansion.

-m msg Set file description.

admin [options] [files...]

Administration of history files in the repository. See Section A.7 [admin], page 94.

-b[rev] Set default branch. See Section 13.3 [Reverting local changes],

page 80.

-cstring Set comment leader. -ksubst Set keyword substitution. See Chapter 12 [Keyword substitution],

page 75.

-l[rev] Lock revision rev, or latest revision.

-mrev:msg

Replace the log message of revision rev with msg.

-orange Delete revisions from the repository. See Section A.7.1 [admin op-

tions], page 94.

-q Run quietly; do not print diagnostics.

-sstate[:rev]

Set the state. See Section A.7.1 [admin options], page 94 for more information on possible states.

-t Set file description from standard input. -tfile Set file description from file.

-t-string

Set file description to string.

-u[rev] Unlock revision rev, or latest revision.

annotate [options] [files...]

Show last revision where each line was modified. See Section A.8 [annotate], page 98. -D date Annotate the most recent revision no later than date. See Sec-

tion A.5 [Common options], page 90.

-F Force annotation of binary files. (Without this option, binary files

are skipped with a message.)

-f Use head revision if tag/date not found. See Section A.5 [Common

options], page 90.

-l Local; run only in current working directory. See Chapter 6 [Re-

cursive behavior], page 51.

-R Operate recursively (default). See Chapter 6 [Recursive behavior],

page 51.

-r tag Annotate revision tag. See Section A.5 [Common options], page 90.

checkout [options] modules...

Get a copy of the sources. See Section A.9 [checkout], page 99. -A Reset any sticky tags/date/options. See Section 4.9 [Sticky tags],

page 38 and Chapter 12 [Keyword substitution], page 75.

-c Output the module database. See Section A.9.1 [checkout options],

page 99.

-D date Check out revisions as of date (is sticky). See Section A.5 [Common

options], page 90.

-d dir Check out into dir. See Section A.9.1 [checkout options], page 99. -f Use head revision if tag/date not found. See Section A.5 [Common

options], page 90.

-j rev Merge in changes. See Section A.9.1 [checkout options], page 99. -k kflag Use kflag keyword expansion. See Section 12.4 [Substitution

modes], page 77.

-l Local; run only in current working directory. See Chapter 6 [Re-

cursive behavior], page 51.

-N Don’t “shorten” module paths if -d specified. See Section A.9.1

[checkout options], page 99.

-n Do not run module program (if any). See Section A.9.1 [checkout

options], page 99.

-P Prune empty directories. See Section 7.5 [Moving directories],

page 57.

-p Check out files to standard output (avoids stickiness). See Sec-

tion A.9.1 [checkout options], page 99.

-R Operate recursively (default). See Chapter 6 [Recursive behavior],

page 51.

-r tag Checkout revision tag (is sticky). See Section A.5 [Common op-

tions], page 90.

-s Like -c, but include module status. See Section A.9.1 [checkout

options], page 99.

commit [options] [files...]

Check changes into the repository. See Section A.10 [commit], page 101. -F file Read log message from file. See Section A.10.1 [commit options],

page 102.

-f Force the file to be committed; disables recursion. See

Section A.10.1 [commit options], page 102.

-l Local; run only in current working directory. See Chapter 6 [Re-

cursive behavior], page 51.

-m msg Use msg as log message. See Section A.10.1 [commit options],

page 102.

-n Do not run module program (if any). See Section A.10.1 [commit options], page 102.

-R Operate recursively (default). See Chapter 6 [Recursive behavior], page 51.

-r rev Commit to rev. See Section A.10.1 [commit options], page 102.

diff [options] [files...]

Show differences between revisions. See Section A.11 [diff], page 103. In addition to the options shown below, accepts a wide variety of options to control output style, for example ‘-c’ for context diffs.

-D date1 Diff revision for date against working file. See Section A.11.1 [diff options], page 104.

-D date2 Diff rev1/date1 against date2. See Section A.11.1 [diff options], page 104.

-l Local; run only in current working directory. See Chapter 6 [Recursive behavior], page 51.

-N Include diffs for added and removed files. See Section A.11.1 [diff options], page 104.

-R Operate recursively (default). See Chapter 6 [Recursive behavior], page 51.

-r rev1 Diff revision for rev1 against working file. See Section A.11.1 [diff options], page 104.

-r rev2 Diff rev1/date1 against rev2. See Section A.11.1 [diff options], page 104.

edit [options] [files...]

Get ready to edit a watched file. See Section 10.6.3 [Editing files], page 70.

-a actions

Specify actions for temporary watch, where actions is edit, unedit, commit, all, or none. See Section 10.6.3 [Editing files], page 70.

-l Local; run only in current working directory. See Chapter 6 [Recursive behavior], page 51.

-R Operate recursively (default). See Chapter 6 [Recursive behavior], page 51.

editors [options] [files...]

See who is editing a watched file. See Section 10.6.4 [Watch information], page 71.

-l Local; run only in current working directory. See Chapter 6 [Recursive behavior], page 51.

-R Operate recursively (default). See Chapter 6 [Recursive behavior], page 51.

export [options] modules...

Export files from cvs. See Section A.12 [export], page 112.

-D date Check out revisions as of date. See Section A.5 [Common options],

page 90.

-d dir Check out into dir. See Section A.12.1 [export options], page 112. -f Use head revision if tag/date not found. See Section A.5 [Common

options], page 90.

-k kflag Use kflag keyword expansion. See Section 12.4 [Substitution

modes], page 77.

-l Local; run only in current working directory. See Chapter 6 [Re-

cursive behavior], page 51.

-N Don’t “shorten” module paths if -d specified. See Section A.12.1

[export options], page 112.

-n Do not run module program (if any). See Section A.12.1 [export

options], page 112.

-R Operate recursively (default). See Chapter 6 [Recursive behavior],

page 51.

-r tag Checkout revision tag. See Section A.5 [Common options], page 90.

history [options] [files...]

Show repository access history. See Section A.13 [history], page 113. -a All users (default is self). See Section A.13.1 [history options],

page 113.

-b str Back to record with str in module/file/repos field. See

Section A.13.1 [history options], page 113.

-c Report on committed (modified) files. See Section A.13.1 [history

options], page 113.

-D date Since date. See Section A.13.1 [history options], page 113. -e Report on all record types. See Section A.13.1 [history options],

page 113.

-l Last modified (committed or modified report). See Section A.13.1

[history options], page 113.

-m module

Report on module (repeatable). See Section A.13.1 [history options], page 113.

-n module

In module. See Section A.13.1 [history options], page 113.

-o Report on checked out modules. See Section A.13.1 [history op-

tions], page 113.

-p repository

In repository. See Section A.13.1 [history options], page 113.

-r rev Since revision rev. See Section A.13.1 [history options], page 113. -T Produce report on all TAGs. See Section A.13.1 [history options],

page 113.

-t tag Since tag record placed in history file (by anyone). See

Section A.13.1 [history options], page 113.

-u user For user user (repeatable). See Section A.13.1 [history options],

page 113.

-w Working directory must match. See Section A.13.1 [history op-

tions], page 113.

-x types Report on types, one or more of TOEFWUPCGMAR. See Section A.13.1

[history options], page 113.

-z zone Output for time zone zone. See Section A.13.1 [history options],

page 113.

import [options] repository vendor-tag release-tags...

Import files into cvs, using vendor branches. See Section A.14 [import], page 115. -b bra Import to vendor branch bra. See Section 13.6 [Multiple vendor

branches], page 81.

-d Use the file’s modification time as the time of import. See Sec-

tion A.14.1 [import options], page 115.

-k kflag Set default keyword substitution mode. See Section A.14.1 [import

options], page 115.

-m msg Use msg for log message. See Section A.14.1 [import options],

page 115.

-I ign More files to ignore (! to reset). See Section A.14.1 [import options],

page 115.

-W spec More wrappers. See Section A.14.1 [import options], page 115.

init Create a cvs repository if it doesn’t exist. See Section 2.6 [Creating a reposi-

tory], page 18.

kserver Kerberos authenticated server. See Section 2.9.5 [Kerberos authenticated],

page 26.

log [options] [files...]

Print out history information for files. See Section A.15 [log], page 116. -b Only list revisions on the default branch. See Section A.15.1 [log

options], page 117.

-d dates Specify dates (d1<d2 for range, d for latest before). See Sec-

tion A.15.1 [log options], page 117.

-h Only print header. See Section A.15.1 [log options], page 117.

-l Local; run only in current working directory. See Chapter 6 [Recursive behavior], page 51.

-N Do not list tags. See Section A.15.1 [log options], page 117.

-R Only print name of RCS file. See Section A.15.1 [log options], page 117.

-rrevs Only list revisions revs. See Section A.15.1 [log options], page 117.

-s states

Only list revisions with specified states. See Section A.15.1 [log options], page 117.

-t Only print header and descriptive text. See Section A.15.1 [log options], page 117.

-wlogins Only list revisions checked in by specified logins. See Section A.15.1 [log options], page 117.

login Prompt for password for authenticating server. See Section 2.9.3.2 [Password authentication client], page 24.

logout Remove stored password for authenticating server. See Section 2.9.3.2 [Password authentication client], page 24.

pserver Password authenticated server. See Section 2.9.3.1 [Password authentication server], page 21.

rannotate [options] [modules...]

Show last revision where each line was modified. See Section A.8 [annotate], page 98.

-D date Annotate the most recent revision no later than date. See Section A.5 [Common options], page 90.

-F Force annotation of binary files. (Without this option, binary files are skipped with a message.)

-f Use head revision if tag/date not found. See Section A.5 [Common options], page 90.

-l Local; run only in current working directory. See Chapter 6 [Recursive behavior], page 51.

-R Operate recursively (default). See Chapter 6 [Recursive behavior], page 51.

-r tag Annotate revision tag. See Section A.5 [Common options], page 90.

rdiff [options] modules...

Show differences between releases. See Section A.16 [rdiff], page 118.

-c Context diff output format (default). See Section A.16.1 [rdiff options], page 119.

-D date Select revisions based on date. See Section A.5 [Common options], page 90.

-f Use head revision if tag/date not found. See Section A.5 [Common

options], page 90.

-l Local; run only in current working directory. See Chapter 6 [Re-

cursive behavior], page 51.

-R Operate recursively (default). See Chapter 6 [Recursive behavior],

page 51.

-r rev Select revisions based on rev. See Section A.5 [Common options],

page 90.

-s Short patch - one liner per file. See Section A.16.1 [rdiff options],

page 119.

-t Top two diffs - last change made to the file. See Section A.11.1 [diff

options], page 104.

-u Unidiff output format. See Section A.16.1 [rdiff options], page 119. -V vers Use RCS Version vers for keyword expansion (obsolete). See Sec-

tion A.16.1 [rdiff options], page 119.

release [options] directory

Indicate that a directory is no longer in use. See Section A.17 [release], page 120. -d Delete the given directory. See Section A.17.1 [release options],

page 120.

remove [options] [files...]

Remove an entry from the repository. See Section 7.2 [Removing files], page 54. -f Delete the file before removing it. See Section 7.2 [Removing files],

page 54.

-l Local; run only in current working directory. See Chapter 6 [Re-

cursive behavior], page 51.

-R Operate recursively (default). See Chapter 6 [Recursive behavior],

page 51.

rlog [options] [files...]

Print out history information for modules. See Section A.15 [log], page 116. -b Only list revisions on the default branch. See Section A.15.1 [log

options], page 117.

-d dates Specify dates (d1<d2 for range, d for latest before). See Sec-

tion A.15.1 [log options], page 117.

-h Only print header. See Section A.15.1 [log options], page 117. -l Local; run only in current working directory. See Chapter 6 [Re-

cursive behavior], page 51.

-N Do not list tags. See Section A.15.1 [log options], page 117. -R Only print name of RCS file. See Section A.15.1 [log options],

page 117.

-rrevs Only list revisions revs. See Section A.15.1 [log options], page 117.

-s states

Only list revisions with specified states. See Section A.15.1 [log options], page 117.

-t Only print header and descriptive text. See Section A.15.1 [log

options], page 117.

-wlogins Only list revisions checked in by specified logins. See Section A.15.1

[log options], page 117.

rtag [options] tag modules...

Add a symbolic tag to a module. See Chapter 4 [Revisions], page 33 and Chapter 5 [Branching and merging], page 41. -a Clear tag from removed files that would not otherwise be tagged.

See Section 4.8 [Tagging add/remove], page 38.

-b Create a branch named tag. See Chapter 5 [Branching and merg-

ing], page 41.

-B Used in conjunction with -F or -d, enables movement and deletion

of branch tags. Use with extreme caution.

-D date Tag revisions as of date. See Section 4.6 [Tagging by date/tag],

page 36.

-d Delete tag. See Section 4.7 [Modifying tags], page 37. -F Move tag if it already exists. See Section 4.7 [Modifying tags],

page 37.

-f Force a head revision match if tag/date not found. See Section 4.6

[Tagging by date/tag], page 36.

-l Local; run only in current working directory. See Chapter 6 [Re-

cursive behavior], page 51.

-n No execution of tag program. See Section A.5 [Common options],

page 90.

-R Operate recursively (default). See Chapter 6 [Recursive behavior],

page 51.

-r rev Tag existing tag rev. See Section 4.6 [Tagging by date/tag], page 36.

server Rsh server. See Section 2.9.2 [Connecting via rsh], page 20.

status [options] files...

Display status information in a working directory. See Section 10.1 [File status], page 63. -l Local; run only in current working directory. See Chapter 6 [Re-

cursive behavior], page 51.

-R Operate recursively (default). See Chapter 6 [Recursive behavior], page 51.

-v Include tag information for file. See Section 4.4 [Tags], page 34.

tag [options] tag [files...]

Add a symbolic tag to checked out version of files. See Chapter 4 [Revisions], page 33 and Chapter 5 [Branching and merging], page 41. -b Create a branch named tag. See Chapter 5 [Branching and merg-

ing], page 41.

-c Check that working files are unmodified. See Section 4.5 [Tagging

the working directory], page 36.

-D date Tag revisions as of date. See Section 4.6 [Tagging by date/tag],

page 36.

-d Delete tag. See Section 4.7 [Modifying tags], page 37. -F Move tag if it already exists. See Section 4.7 [Modifying tags],

page 37.

-f Force a head revision match if tag/date not found. See Section 4.6

[Tagging by date/tag], page 36.

-l Local; run only in current working directory. See Chapter 6 [Re-

cursive behavior], page 51.

-R Operate recursively (default). See Chapter 6 [Recursive behavior],

page 51.

-r rev Tag existing tag rev. See Section 4.6 [Tagging by date/tag], page 36.

unedit [options] [files...]

Undo an edit command. See Section 10.6.3 [Editing files], page 70. -l Local; run only in current working directory. See Chapter 6 [Re-

cursive behavior], page 51.

-R Operate recursively (default). See Chapter 6 [Recursive behavior],

page 51.

update [options] [files...]

Bring work tree in sync with repository. See Section A.19 [update], page 122. -A Reset any sticky tags/date/options. See Section 4.9 [Sticky tags],

page 38 and Chapter 12 [Keyword substitution], page 75.

-C Overwrite locally modified files with clean copies from the reposi-

tory (the modified file is saved in ‘.#file.revision’, however).

-D date Check out revisions as of date (is sticky). See Section A.5 [Common

options], page 90.

-d Create directories. See Section A.19.1 [update options], page 123. -f Use head revision if tag/date not found. See Section A.5 [Common

options], page 90.

-I ign More files to ignore (! to reset). See Section A.14.1 [import options], page 115.

-j rev Merge in changes. See Section A.19.1 [update options], page 123.

-k kflag Use kflag keyword expansion. See Section 12.4 [Substitution modes], page 77.

-l Local; run only in current working directory. See Chapter 6 [Recursive behavior], page 51.

-P Prune empty directories. See Section 7.5 [Moving directories], page 57.

-p Check out files to standard output (avoids stickiness). See Section A.19.1 [update options], page 123.

-R Operate recursively (default). See Chapter 6 [Recursive behavior], page 51.

-r tag Checkout revision tag (is sticky). See Section A.5 [Common options], page 90.

-W spec More wrappers. See Section A.14.1 [import options], page 115.

version

Display the version of cvs being used. If the repository is remote, display both the client and server versions.

watch [on|off|add|remove] [options] [files...]

on/off: turn on/off read-only checkouts of files. See Section 10.6.1 [Setting a watch], page 68. add/remove: add or remove notification on actions. See Section 10.6.2 [Getting Notified], page 69.

-a actions

Specify actions for temporary watch, where actions is edit, unedit, commit, all, or none. See Section 10.6.3 [Editing files], page 70.

-l Local; run only in current working directory. See Chapter 6 [Recursive behavior], page 51.

-R Operate recursively (default). See Chapter 6 [Recursive behavior], page 51.

watchers [options] [files...]

See who is watching a file. See Section 10.6.4 [Watch information], page 71.

-l Local; run only in current working directory. See Chapter 6 [Recursive behavior], page 51.

-R Operate recursively (default). See Chapter 6 [Recursive behavior], page 51.

Appendix C Reference manual for Administrative files

Inside the repository, in the directory ‘$CVSROOT/CVSROOT’, there are a number of supportive files for cvs. You can use cvs in a limited fashion without any of them, but if they are set up properly they can help make life easier. For a discussion of how to edit them, see Section 2.4 [Intro administrative files], page 16.

The most important of these files is the ‘modules’ file, which defines the modules inside the repository.

C.1 The modules file

The ‘modules’ file records your definitions of names for collections of source code. cvs will use these definitions if you use cvs to update the modules file (use normal commands like

add, commit, etc).

The ‘modules’ file may contain blank lines and comments (lines beginning with ‘#’) as well as module definitions. Long lines can be continued on the next line by specifying a backslash (‘\’) as the last character on the line.

There are three basic types of modules: alias modules, regular modules, and ampersand modules. The difference between them is the way that they map files in the repository to files in the working directory. In all of the following examples, the top-level repository contains a directory called ‘first-dir’, which contains two files, ‘file1’ and ‘file2’, and a directory ‘sdir’. ‘first-dir/sdir’ contains a file ‘sfile’.

C.1.1 Alias modules

Alias modules are the simplest kind of module:

mname -a aliases...

This represents the simplest way of defining a module mname. The ‘-a’ flags the definition as a simple alias: cvs will treat any use of mname (as a command argument) as if the list of names aliases had been specified instead. aliases may contain either other module names or paths. When you use paths in aliases, checkout creates all intermediate directories in the working directory, just as if the path had been specified explicitly in the cvs arguments.

For example, if the modules file contains:

     amodule -a first-dir

then the following two commands are equivalent:

     $ cvs co amodule
     $ cvs co first-dir

and they each would provide output such as:

     cvs checkout: Updating first-dir
     U first-dir/file1
     U first-dir/file2
     cvs checkout: Updating first-dir/sdir
     U first-dir/sdir/sfile

C.1.2 Regular modules

mname [ options ] dir [ files... ]

In the simplest case, this form of module definition reduces to ‘mname dir’. This defines all the files in directory dir as module mname. dir is a relative path (from $CVSROOT) to a directory of source in the source repository. In this case, on checkout, a single directory called mname is created as a working directory; no intermediate directory levels are used by default, even if dir was a path involving several directory levels.

For example, if a module is defined by:

regmodule first-dir

then regmodule will contain the files from first-dir:

$ cvs co regmodule
cvs checkout: Updating regmodule
U regmodule/file1
U regmodule/file2
cvs checkout: Updating regmodule/sdir
U regmodule/sdir/sfile
$

By explicitly specifying files in the module definition after dir, you can select particular files from directory dir. Here is an example:

regfiles first-dir/sdir sfile

With this definition, getting the regfiles module will create a single working directory ‘regfiles’ containing the file listed, which comes from a directory deeper in the cvs source repository:

$ cvs co regfiles
U regfiles/sfile
$

C.1.3 Ampersand modules

A module definition can refer to other modules by including ‘&module’ in its definition.

mname [ options ] &module...

Then getting the module creates a subdirectory for each such module, in the directory containing the module. For example, if modules contains

ampermod &first-dir

then a checkout will create an ampermod directory which contains a directory called firstdir, which in turns contains all the directories and files which live there. For example, the command

$ cvs co ampermod

will create the following files:

ampermod/first-dir/file1
ampermod/first-dir/file2
ampermod/first-dir/sdir/sfile

There is one quirk/bug: the messages that cvs prints omit the ‘ampermod’, and thus do not correctly display the location to which it is checking out the files:

     $ cvs co ampermod
     cvs checkout: Updating first-dir
     U first-dir/file1
     U first-dir/file2
     cvs checkout: Updating first-dir/sdir
     U first-dir/sdir/sfile
     $

Do not rely on this buggy behavior; it may get fixed in a future release of cvs.

C.1.4 Excluding directories

An alias module may exclude particular directories from other modules by using an exclamation mark (‘!’) before the name of each directory to be excluded.

For example, if the modules file contains:

     exmodule -a !first-dir/sdir first-dir

then checking out the module ‘exmodule’ will check out everything in ‘first-dir’ except any files in the subdirectory ‘first-dir/sdir’.

C.1.5 Module options

Either regular modules or ampersand modules can contain options, which supply additional information concerning the module.

-d name Name the working directory something other than the module name.

-e prog Specify a program prog to run whenever files in a module are exported. prog runs with a single argument, the module name.

-o prog Specify a program prog to run whenever files in a module are checked out. prog runs with a single argument, the module name. See Section C.1.6 [Module program options], page 142 for information on how prog is called.

-s status

Assign a status to the module. When the module file is printed with ‘cvs checkout -s’ the modules are sorted according to primarily module status, and secondarily according to the module name. This option has no other meaning. You can use this option for several things besides status: for instance, list the person that is responsible for this module.

-t prog Specify a program prog to run whenever files in a module are tagged with rtag. prog runs with two arguments: the module name and the symbolic tag specified to rtag. It is not run when tag is executed. Generally you will find that the ‘taginfo’ file is a better solution (see Section C.6 [taginfo], page 150).

You should also see see Section C.1.6 [Module program options], page 142 about how the “program options” programs are run.

C.1.6 How the modules file “program options” programs are run

For checkout, rtag, and export, the program is server-based, and as such the following applies:-

If using remote access methods (pserver, ext, etc.), cvs will execute this program on the server from a temporary directory. The path is searched for this program.

If using “local access” (on a local or remote NFS file system, i.e., repository set just to a path), the program will be executed from the newly checked-out tree, if found there, or alternatively searched for in the path if not.

The programs are all run after the operation has effectively completed.

C.2 The cvswrappers file

Wrappers refers to a cvs feature which lets you control certain settings based on the name of the file which is being operated on. The settings are ‘-k’ for binary files, and ‘-m’ for nonmergeable text files.

The ‘-m’ option specifies the merge methodology that should be used when a non-binary file is updated. MERGE means the usual cvs behavior: try to merge the files. COPY means that cvs update will refuse to merge files, as it also does for files specified as binary with ‘-kb’ (but if the file is specified as binary, there is no need to specify ‘-m 'COPY'’). cvs will provide the user with the two versions of the files, and require the user using mechanisms outside cvs, to insert any necessary changes.

WARNING: Do not use COPY with cvs 1.9 or earlier - such versions of cvs will copy one version of your file over the other, wiping out the previous contents. The ‘-m’ wrapper option only affects behavior when merging is done on update; it does not affect how files are stored. See Chapter 9 [Binary files], page 61, for more on binary files.

The basic format of the file ‘cvswrappers’ is:

     wildcard     [option value][option value]...
     where option is one of
     -m           update methodology      value: MERGE or COPY
     -k           keyword expansion       value: expansion mode
     and value is a single-quote delimited value.

For example, the following command imports a directory, treating files whose name ends in ‘.exe’ as binary:

     cvs import -I ! -W "*.exe -k 'b'" first-dir vendortag reltag

C.3 The Trigger Scripts

Several of the administrative files support triggers, or the launching external scripts or programs at specific times before or after particular events. The individual files are discussed in the later sections, Section C.4 [commit files], page 143 and Section C.6 [taginfo], page 150, but some of the common elements are discussed here.

All the trigger scripts are launched in a copy of the user sandbox being committed, on the server, in client-server mode. In local mode, the scripts are actually launched directly from the user sandbox directory being committed. For most intents and purposes, the same scripts can be run in both locations without alteration.

C.3.1 The common syntax

The administrative files such as ‘commitinfo’, ‘loginfo’, ‘rcsinfo’, ‘verifymsg’, etc., all have a common format. The purpose of the files are described later on. The common syntax is described here.

Each line contains the following:

• A regular expression. This is a basic regular expression in the syntax used by GNU emacs.

• A whitespace separator—one or more spaces and/or tabs.

• A file name or command-line template.

Blank lines are ignored. Lines that start with the character ‘#’ are treated as comments. Long lines unfortunately can not be broken in two parts in any way.

The first regular expression that matches the current directory name in the repository is used. The rest of the line is used as a file name or command-line as appropriate.

C.3.2 Security and the Trigger Scripts

Security is a huge subject, and implementing a secure system is a non-trivial task. This section will barely touch on all the issues involved, but it is well to note that, as with any script you will be allowing an untrusted user to run on your server, there are measures you can take to help prevent your trigger scripts from being abused.

For instance, since the CVS trigger scripts all run in a copy of the user’s sandbox on the server, a naively coded Perl trigger script which attempts to use a Perl module that is not installed on the system can be hijacked by any user with commit access who is checking in a file with the correct name. Other scripting languages may be vulnerable to similar hacks.

One way to make a script more secure, at least with Perl, is to use scripts which invoke the -T, or "taint-check" switch on their #! line. In the most basic terms, this causes Perl to avoid running code that may have come from an external source. Please run the perldoc perlsec command for more on Perl security. Again, other languages may implement other security verification hooks which look more or less like Perl’s "taint-check" mechanism.

C.4 The commit support files

There are three kinds of trigger scripts (see Section C.3 [Trigger Scripts], page 142) that can be run at various times during a commit. They are specified in files in the repository, as described below. The following table summarizes the file names and the purpose of the corresponding programs.

‘commitinfo’

The program is responsible for checking that the commit is allowed. If it exits with a non-zero exit status the commit will be aborted.

‘verifymsg’

The specified program is used to evaluate the log message, and possibly verify that it contains all required fields. This is most useful in combination with the ‘rcsinfo’ file, which can hold a log message template (see Section C.5 [rcsinfo], page 149).

‘editinfo’

The specified program is used to edit the log message, and possibly verify that it contains all required fields. This is most useful in combination with the ‘rcsinfo’ file, which can hold a log message template (see Section C.5 [rcsinfo], page 149). (obsolete)

‘loginfo’ The specified program is called when the commit is complete. It receives the log message and some additional information and can store the log message in a file, or mail it to appropriate persons, or maybe post it to a local newsgroup, or… Your imagination is the limit!

C.4.1 Commitinfo

The ‘commitinfo’ file defines programs to execute whenever ‘cvs commit’ is about to execute. These programs are used for pre-commit checking to verify that the modified, added and removed files are really ready to be committed. This could be used, for instance, to verify that the changed files conform to to your site’s standards for coding practice.

As mentioned earlier, each line in the ‘commitinfo’ file consists of a regular expression and a command-line template. The template can include a program name and any number of arguments you wish to supply to it. The full path to the current source repository is appended to the template, followed by the file names of any files involved in the commit (added, removed, and modified files).

The first line with a regular expression matching the directory within the repository will be used. If the command returns a non-zero exit status the commit will be aborted.

If the repository name does not match any of the regular expressions in this file, the ‘DEFAULT’ line is used, if it is specified.

All occurrences of the name ‘ALL’ appearing as a regular expression are used in addition to the first matching regular expression or the name ‘DEFAULT’.

The command will be run in the root of the workspace containing the new versions of any files the user would like to modify (commit), or in a copy of the workspace on the server (see Section 2.9 [Remote repositories], page 19). If a file is being removed, there will be no copy of the file under the current directory. If a file is being added, there will be no corresponding archive file in the repository unless the file is being resurrected.

Note that both the repository directory and the corresponding Attic (see Section 2.2.4 [Attic], page 11) directory may need to be checked to locate the archive file corresponding to any given file being committed. Much of the information about the specific commit request being made, including the destination branch, commit message, and command line options specified, is not available to the command.

C.4.2 Verifying log messages

Once you have entered a log message, you can evaluate that message to check for specific content, such as a bug ID. Use the ‘verifymsg’ file to specify a program that is used to verify the log message. This program could be a simple script that checks that the entered message contains the required fields.

The ‘verifymsg’ file is often most useful together with the ‘rcsinfo’ file, which can be used to specify a log message template.

Each line in the ‘verifymsg’ file consists of a regular expression and a command-line template. The template must include a program name, and can include any number of arguments. The full path to the current log message template file is appended to the template.

One thing that should be noted is that the ‘ALL’ keyword is not supported. If more than one matching line is found, the first one is used. This can be useful for specifying a default verification script in a directory, and then overriding it in a subdirectory.

If the repository name does not match any of the regular expressions in this file, the ‘DEFAULT’ line is used, if it is specified.

If the verification script exits with a non-zero exit status, the commit is aborted.

In the default configuration, CVS allows the verification script to change the log message. This is controlled via the RereadLogAfterVerify CVSROOT/config option.

   When ‘RereadLogAfterVerify=always’ or ‘RereadLogAfterVerify=stat’, the log mes-

sage will either always be reread after the verification script is run or reread only if the log message file status has changed.

See Section C.11 [config], page 153, for more on CVSROOT/config options.

It is NOT a good idea for a ‘verifymsg’ script to interact directly with the user in the various client/server methods. For the pserver method, there is no protocol support for communicating between ‘verifymsg’ and the client on the remote end. For the ext and server methods, it is possible for CVS to become confused by the characters going along the same channel as the CVS protocol messages. See Section 2.9 [Remote repositories], page 19, for more information on client/server setups. In addition, at the time the ‘verifymsg’ script runs, the CVS server has locks in place in the repository. If control is returned to the user here then other users may be stuck waiting for access to the repository.

This option can be useful if you find yourself using an rcstemplate that needs to be modified to remove empty elements or to fill in default values. It can also be useful if the rcstemplate has changed in the repository and the CVS/Template was not updated, but is able to be adapted to the new format by the verification script that is run by ‘verifymsg’.

An example of an update might be to change all occurrences of ’BugId:’ to be ’DefectId:’ (which can be useful if the rcstemplate has recently been changed and there are still checkedout user trees with cached copies in the CVS/Template file of the older version).

Another example of an update might be to delete a line that contains ’BugID: none’ from the log message after validation of that value as being allowed is made.

The following is a little silly example of a ‘verifymsg’ file, together with the corresponding ‘rcsinfo’ file, the log message template and an verification script. We begin with the log message template. We want to always record a bug-id number on the first line of the log message. The rest of log message is free text. The following template is found in the

file ‘/usr/cvssupport/tc.template’.
     BugId:

The script ‘/usr/cvssupport/bugid.verify’ is used to evaluate the log message.

     #!/bin/sh

     #
     #       bugid.verify filename
     #
     #  Verify that the log message contains a valid bugid
     #  on the first line.
     #
     if sed 1q < $1 | grep '^BugId:[ ]*[0-9][0-9]*$' > /dev/null; then
         exit 0
     elif sed 1q < $1 | grep '^BugId:[ ]*none$' > /dev/null; then
         # It is okay to allow commits with 'BugId: none',
         # but do not put that text into the real log message.
         grep -v '^BugId:[ ]*none$' > $1.rewrite
         mv $1.rewrite $1
         exit 0
     else
         echo "No BugId found."
         exit 1
     fi

The ‘verifymsg’ file contains this line:

     ^tc     /usr/cvssupport/bugid.verify

The ‘rcsinfo’ file contains this line:

     ^tc     /usr/cvssupport/tc.template

The ‘config’ file contains this line:

     RereadLogAfterVerify=always

C.4.3 Editinfo

The ‘editinfo’ feature has been rendered obsolete. To set a default editor for log messages use the CVSEDITOR, EDITOR environment variables (see Appendix D [Environment variables], page 155) or the ‘-e’ global option (see Section A.4 [Global options], page 88). See Section C.4.2 [verifymsg], page 144, for information on the use of the ‘verifymsg’ feature for evaluating log messages.

If you want to make sure that all log messages look the same way, you can use the ‘editinfo’ file to specify a program that is used to edit the log message. This program could be a custom-made editor that always enforces a certain style of the log message, or maybe a simple shell script that calls an editor, and checks that the entered message contains the required fields.

If no matching line is found in the ‘editinfo’ file, the editor specified in the environment variable $CVSEDITOR is used instead. If that variable is not set, then the environment variable $EDITOR is used instead. If that variable is not set a default will be used. See Section 1.3.2 [Committing your changes], page 4.

The ‘editinfo’ file is often most useful together with the ‘rcsinfo’ file, which can be used to specify a log message template.

Each line in the ‘editinfo’ file consists of a regular expression and a command-line template. The template must include a program name, and can include any number of arguments. The full path to the current log message template file is appended to the template.

One thing that should be noted is that the ‘ALL’ keyword is not supported. If more than one matching line is found, the first one is used. This can be useful for specifying a default edit script in a module, and then overriding it in a subdirectory.

If the repository name does not match any of the regular expressions in this file, the ‘DEFAULT’ line is used, if it is specified.

If the edit script exits with a non-zero exit status, the commit is aborted.

Note: when cvs is accessing a remote repository, or when the ‘-m’ or ‘-F’ options to cvs commit are used, ‘editinfo’ will not be consulted. There is no good workaround for this; use ‘verifymsg’ instead.

C.4.3.1 Editinfo example

The following is a little silly example of a ‘editinfo’ file, together with the corresponding ‘rcsinfo’ file, the log message template and an editor script. We begin with the log message template. We want to always record a bug-id number on the first line of the log message. The rest of log message is free text. The following template is found in the

file ‘/usr/cvssupport/tc.template’.
BugId:

The script ‘/usr/cvssupport/bugid.edit’ is used to edit the log message.

#!/bin/sh
#
#       bugid.edit filename
#
#  Call $EDITOR on FILENAME, and verify that the
#  resulting file contains a valid bugid on the first
#  line.
if [ "x$EDITOR" = "x" ]; then EDITOR=vi; fi
if [ "x$CVSEDITOR" = "x" ]; then CVSEDITOR=$EDITOR; fi
$CVSEDITOR $1
until head -1|grep '^BugId:[ ]*[0-9][0-9]*$' < $1
do  echo -n  "No BugId found.  Edit again? ([y]/n)"
    read ans
    case ${ans} in
        n*) exit 1;;
    esac
    $CVSEDITOR $1
done

The ‘editinfo’ file contains this line:

^tc     /usr/cvssupport/bugid.edit

The ‘rcsinfo’ file contains this line:

^tc     /usr/cvssupport/tc.template

C.4.4 Loginfo

The ‘loginfo’ file is used to control where ‘cvs commit’ log information is sent. The first entry on a line is a regular expression which is tested against the directory that the change is being made to, relative to the $CVSROOT. If a match is found, then the remainder of the line is a filter program that should expect log information on its standard input. Note that the filter program must read all of the log information or cvs may fail with a broken pipe signal.

If the repository name does not match any of the regular expressions in this file, the ‘DEFAULT’ line is used, if it is specified.

All occurrences of the name ‘ALL’ appearing as a regular expression are used in addition to the first matching regular expression or ‘DEFAULT’.

The first matching regular expression is used.

See Section C.4 [commit files], page 143, for a description of the syntax of the ‘loginfo’ file.

The user may specify a format string as part of the filter. The string is composed of a ‘%’ followed by a space, or followed by a single format character, or followed by a set of format characters surrounded by ‘{’ and ‘}’ as separators. The format characters are:

s file name

V old version number (pre-checkin)

v new version number (post-checkin)

All other characters that appear in a format string expand to an empty field (commas separating fields are still provided).

For example, some valid format strings are ‘%’, ‘%s’, ‘%{s}’, and ‘%{sVv}’.

The output will be a space separated string of tokens enclosed in quotation marks ("). Any embedded dollar signs ($), backticks (`), backslashes (\), or quotation marks will be preceded by a backslash (this allows the shell to correctly parse it as a single string, reguardless of the characters it contains). For backwards compatibility, the first token will be the repository subdirectory. The rest of the tokens will be comma-delimited lists of the information requested in the format string. For example, if ‘/u/src/master/yoyodyne/tc’ is the repository, ‘%{sVv}’ is the format string, and three files (ChangeLog, Makefile, foo.c) were modified, the output might be:

     "yoyodyne/tc ChangeLog,1.1,1.2 Makefile,1.3,1.4 foo.c,1.12,1.13"

As another example, ‘%{}’ means that only the name of the repository will be generated.

Note: when cvs is accessing a remote repository, ‘loginfo’ will be run on the remote (i.e., server) side, not the client side (see Section 2.9 [Remote repositories], page 19).

C.4.4.1 Loginfo example

The following ‘loginfo’ file, together with the tiny shell-script below, appends all log messages to the file ‘$CVSROOT/CVSROOT/commitlog’, and any commits to the administrative files (inside the ‘CVSROOT’ directory) are also logged in ‘/usr/adm/cvsroot-log’. Commits to the ‘prog1’ directory are mailed to ceder.

     ALL             /usr/local/bin/cvs-log $CVSROOT/CVSROOT/commitlog $USER
     ^CVSROOT        /usr/local/bin/cvs-log /usr/adm/cvsroot-log
     ^prog1          Mail -s %s ceder

The shell-script ‘/usr/local/bin/cvs-log’ looks like this:

     #!/bin/sh
     (echo "------------------------------------------------------";
      echo -n $2"  ";
      date;
      echo;
      cat) >> $1

C.4.4.2 Keeping a checked out copy

It is often useful to maintain a directory tree which contains files which correspond to the latest version in the repository. For example, other developers might want to refer to the latest sources without having to check them out, or you might be maintaining a web site with cvs and want every checkin to cause the files used by the web server to be updated.

The way to do this is by having loginfo invoke cvs update. Doing so in the naive way will cause a problem with locks, so the cvs update must be run in the background. Here is an example for unix (this should all be on one line):

     ^cyclic-pages (date; cat; (sleep 2; cd /u/www/local-docs;
      cvs -q update -d) &) >> $CVSROOT/CVSROOT/updatelog 2>&1

This will cause checkins to repository directories starting with cyclic-pages to update the checked out tree in ‘/u/www/local-docs’.

C.5 Rcsinfo

The ‘rcsinfo’ file can be used to specify a form to edit when filling out the commit log. The ‘rcsinfo’ file has a syntax similar to the ‘verifymsg’, ‘commitinfo’ and ‘loginfo’ files. See Section C.3.1 [syntax], page 143. Unlike the other files the second part is not a command-line template. Instead, the part after the regular expression should be a full pathname to a file containing the log message template.

If the repository name does not match any of the regular expressions in this file, the ‘DEFAULT’ line is used, if it is specified.

All occurrences of the name ‘ALL’ appearing as a regular expression are used in addition to the first matching regular expression or ‘DEFAULT’.

The log message template will be used as a default log message. If you specify a log message with ‘cvs commit -m message’ or ‘cvs commit -f file’ that log message will override the template.

See Section C.4.2 [verifymsg], page 144, for an example ‘rcsinfo’ file.

When cvs is accessing a remote repository, the contents of ‘rcsinfo’ at the time a directory is first checked out will specify a template which does not then change. If you edit ‘rcsinfo’ or its templates, you may need to check out a new working directory.

C.6 Taginfo

The ‘taginfo’ file defines programs to execute when someone executes a tag or rtag command. The ‘taginfo’ file has the standard form for trigger scripts (see Section C.3 [Trigger Scripts], page 142), where each line is a regular expression followed by a command to execute (see Section C.3.1 [syntax], page 143). The arguments passed to the command are, in order, the tagname, operation (add for tag, mov for tag -F, and del for tag -d), repository, and any remaining are pairs of filename revision. A non-zero exit of the filter program will cause the tag to be aborted.

Here is an example of using the ‘taginfo’ file to log tag and rtag commands. In the ‘taginfo’ file put:

     ALL /usr/local/cvsroot/CVSROOT/loggit

Where ‘/usr/local/cvsroot/CVSROOT/loggit’ contains the following script:

     #!/bin/sh
     echo "$@" >>/home/kingdon/cvsroot/CVSROOT/taglog

C.7 Ignoring files via cvsignore

There are certain file names that frequently occur inside your working copy, but that you don’t want to put under cvs control. Examples are all the object files that you get while you compile your sources. Normally, when you run ‘cvs update’, it prints a line for each file it encounters that it doesn’t know about (see Section A.19.2 [update output], page 124).

cvs has a list of files (or sh(1) file name patterns) that it should ignore while running update, import and release. This list is constructed in the following way.

• The list is initialized to include certain file name patterns: names associated with cvs administration, or with other common source control systems; common names for patch files, object files, archive files, and editor backup files; and other names that are usually artifacts of assorted utilities. Currently, the default list of ignored file name patterns is:

             RCS     SCCS    CVS     CVS.adm
             RCSLOG  cvslog.*
             tags    TAGS
             .make.state     .nse_depinfo
             *~      #*      .#*     ,*      _$*     *$
             *.old   *.bak   *.BAK   *.orig  *.rej   .del-*
             *.a     *.olb   *.o     *.obj   *.so    *.exe
             *.Z     *.elc   *.ln
             core

• The per-repository list in ‘$CVSROOT/CVSROOT/cvsignore’ is appended to the list, if that file exists.

• The per-user list in ‘.cvsignore’ in your home directory is appended to the list, if it exists.

• Any entries in the environment variable $CVSIGNORE is appended to the list.

• Any ‘-I’ options given to cvs is appended.

• As cvs traverses through your directories, the contents of any ‘.cvsignore’ will be appended to the list. The patterns found in ‘.cvsignore’ are only valid for the directory that contains them, not for any sub-directories.

In any of the 5 places listed above, a single exclamation mark (‘!’) clears the ignore list. This can be used if you want to store any file which normally is ignored by cvs.

Specifying ‘-I !’ to cvs import will import everything, which is generally what you want to do if you are importing files from a pristine distribution or any other source which is known to not contain any extraneous files. However, looking at the rules above you will see there is a fly in the ointment; if the distribution contains any ‘.cvsignore’ files, then the patterns from those files will be processed even if ‘-I !’ is specified. The only workaround is to remove the ‘.cvsignore’ files in order to do the import. Because this is awkward, in the future ‘-I !’ might be modified to override ‘.cvsignore’ files in each directory.

Note that the syntax of the ignore files consists of a series of lines, each of which contains a space separated list of filenames. This offers no clean way to specify filenames which contain spaces, but you can use a workaround like ‘foo?bar’ to match a file named ‘foo bar’ (it also matches ‘fooxbar’ and the like). Also note that there is currently no way to specify comments.

C.8 The checkoutlist file

It may be helpful to use cvs to maintain your own files in the ‘CVSROOT’ directory. For example, suppose that you have a script ‘logcommit.pl’ which you run by including the following line in the ‘commitinfo’ administrative file:

     ALL   $CVSROOT/CVSROOT/logcommit.pl

To maintain ‘logcommit.pl’ with cvs you would add the following line to the ‘checkoutlist’ administrative file:

     logcommit.pl

The format of ‘checkoutlist’ is one line for each file that you want to maintain using cvs, giving the name of the file, followed optionally by more whitespace and any error message that should print if the file cannot be checked out into CVSROOT after a commit:

     logcommit.pl Could not update CVSROOT/logcommit.pl.

After setting up ‘checkoutlist’ in this fashion, the files listed there will function just like cvs’s built-in administrative files. For example, when checking in one of the files you should get a message such as:

     cvs commit: Rebuilding administrative file database

and the checked out copy in the ‘CVSROOT’ directory should be updated.

Note that listing ‘passwd’ (see Section 2.9.3.1 [Password authentication server], page 21) in ‘checkoutlist’ is not recommended for security reasons.

For information about keeping a checkout out copy in a more general context than the one provided by ‘checkoutlist’, see Section C.4.4.2 [Keeping a checked out copy], page 149.

C.9 The history file

The file ‘$CVSROOT/CVSROOT/history’ is used to log information for the history command (see Section A.13 [history], page 113). This file must be created to turn on logging. This is done automatically if the cvs init command is used to set up the repository (see Section 2.6 [Creating a repository], page 18).

The file format of the ‘history’ file is documented only in comments in the cvs source code, but generally programs should use the cvs history command to access it anyway, in case the format changes with future releases of cvs.

C.10 Expansions in administrative files

Sometimes in writing an administrative file, you might want the file to be able to know various things based on environment cvs is running in. There are several mechanisms to do that.

To find the home directory of the user running cvs (from the HOME environment variable), use ‘~’ followed by ‘/’ or the end of the line. Likewise for the home directory of user, use ‘~user’. These variables are expanded on the server machine, and don’t get any reasonable expansion if pserver (see Section 2.9.3 [Password authenticated], page 21) is in use; therefore user variables (see below) may be a better choice to customize behavior based on the user running cvs.

One may want to know about various pieces of information internal to cvs. A cvs internal variable has the syntax ${variable}, where variable starts with a letter and consists of alphanumeric characters and ‘_’. If the character following variable is a non-alphanumeric character other than ‘_’, the ‘{’ and ‘}’ can be omitted. The cvs internal variables are:

CVSROOT This is the absolute path to the current cvs root directory. See Chapter 2 [Repository], page 7, for a description of the various ways to specify this, but note that the internal variable contains just the directory and not any of the access method information.

RCSBIN In cvs 1.9.18 and older, this specified the directory where cvs was looking for rcs programs. Because cvs no longer runs rcs programs, specifying this internal variable is now an error.

CVSEDITOR
EDITOR

VISUAL These all expand to the same value, which is the editor that cvs is using. See Section A.4 [Global options], page 88, for how to specify this.

USER Username of the user running cvs (on the cvs server machine). When using pserver, this is the user specified in the repository specification which need not be the same as the username the server is running as (see Section 2.9.3.1 [Password authentication server], page 21). Do not confuse this with the environment variable of the same name.

If you want to pass a value to the administrative files which the user who is running cvs can specify, use a user variable. To expand a user variable, the administrative file contains ${=variable}. To set a user variable, specify the global option ‘-s’ to cvs, with argument variable=value. It may be particularly useful to specify this option via ‘.cvsrc’ (see Section A.3 [~/.cvsrc], page 88).

For example, if you want the administrative file to refer to a test directory you might create a user variable TESTDIR. Then if cvs is invoked as

cvs -s TESTDIR=/work/local/tests

and the administrative file contains sh ${=TESTDIR}/runtests, then that string is ex-

panded to sh /work/local/tests/runtests.

All other strings containing ‘$’ are reserved; there is no way to quote a ‘$’ character so that ‘$’ represents itself.

Environment variables passed to administrative files are:

CVS_USER The cvs-specific username provided by the user, if it can be provided (currently just for the pserver access method), and to the empty string otherwise. (CVS_

USER and USER may differ when ‘$CVSROOT/CVSROOT/passwd’ is used to map

cvs usernames to system usernames.)

LOGNAME The username of the system user.

USER Same as LOGNAME. Do not confuse this with the internal variable of the same name.

C.11 The CVSROOT/config configuration file

The administrative file ‘config’ contains various miscellaneous settings which affect the behavior of cvs. The syntax is slightly different from the other administrative files. Variables are not expanded. Lines which start with ‘#’ are considered comments. Other lines consist of a keyword, ‘=’, and a value. Note that this syntax is very strict. Extraneous spaces or tabs are not permitted.

Currently defined keywords are:

RCSBIN=bindir

For cvs 1.9.12 through 1.9.18, this setting told cvs to look for rcs programs in the bindir directory. Current versions of cvs do not run rcs programs; for compatibility this setting is accepted, but it does nothing.

SystemAuth=value

If value is ‘yes’, then pserver should check for users in the system’s user database if not found in ‘CVSROOT/passwd’. If it is ‘no’, then all pserver users must exist in ‘CVSROOT/passwd’. The default is ‘yes’. For more on pserver, see Section 2.9.3 [Password authenticated], page 21.

TopLevelAdmin=value

Modify the ‘checkout’ command to create a ‘CVS’ directory at the top level of the new working directory, in addition to ‘CVS’ directories created within checked-out directories. The default value is ‘no’. This option is useful if you find yourself performing many commands at the top level of your working directory, rather than in one of the checked out subdirectories. The ‘CVS’ directory created there will mean you don’t have to specify CVSROOT for each command. It also provides a place for the ‘CVS/Template’ file (see Section 2.3 [Working directory storage], page 13).

LockDir=directory

Put cvs lock files in directory rather than directly in the repository. This is useful if you want to let users read from the repository while giving them write access only to directory, not to the repository. It can also be used to put the locks on a very fast in-memory file system to speed up locking and unlocking the repository. You need to create directory, but cvs will create subdirectories of directory as it needs them. For information on cvs locks, see Section 10.5 [Concurrency], page 67. Before enabling the LockDir option, make sure that you have tracked down and removed any copies of cvs 1.9 or older. Such versions neither support LockDir, nor will give an error indicating that they don’t support it. The result, if this is allowed to happen, is that some cvs users will put the locks one place, and others will put them another place, and therefore the repository could become corrupted. cvs 1.10 does not support LockDir but it will print a warning if run on a repository with LockDir enabled.

LogHistory=value

Control what is logged to the ‘CVSROOT/history’ file (see Section A.13 [history], page 113). Default of ‘TOEFWUPCGMAR’ (or simply ‘all’) will log all transactions. Any subset of the default is legal. (For example, to only log transactions that modify the ‘*,v’ files, use ‘LogHistory=TMAR’.)

RereadLogAfterVerify=value

Modify the ‘commit’ command such that CVS will reread the log message after running the program specified by ‘verifymsg’. value may be one of ‘yes’ or ‘always’, indicating that the log message should always be reread; ‘no’ or ‘never’, indicating that it should never be reread; or value may be ‘stat’, indicating that the file should be checked with the file system ‘stat()’ function to see if it has changed (see warning below) before rereading. The default value

is ‘always’.

The ‘stat’ mode can cause CVS to pause for up to one extra second per directory committed. This can be less IO and CPU intensive but is not recommended for use with large repositories See Section C.4.2 [verifymsg], page 144, for more information on how verifymsg may be used.

IgnoreUnknownConfigKeys=value

If value is ‘yes’, then cvs should ignore any keywords in ‘CVSROOT/config’ which it does not recognize. This option is intended primarily for transitions between versions of cvs which support more configuration options in an environment where a read-only mirror of the current cvs server may be maintained by someone else who is not yet ready to upgrade to the same version. It is recommended that this option be used only for a short time so that problems with the ‘CVSROOT/config’ file will be found quickly. The default is ‘no’.

Appendix D All environment variables which affect CVS

This is a complete list of all environment variables that affect cvs.

$CVSIGNORE

A whitespace-separated list of file name patterns that cvs should ignore. See Section C.7 [cvsignore], page 150.

$CVSWRAPPERS

A whitespace-separated list of file name patterns that cvs should treat as wrappers. See Section C.2 [Wrappers], page 142.

$CVSREAD If this is set, checkout and update will try hard to make the files in your working directory read-only. When this is not set, the default behavior is to permit modification of your working files.

$CVSUMASK

Controls permissions of files in the repository. See Section 2.2.2 [File permissions], page 9.

$CVSROOT Should contain the full pathname to the root of the cvs source repository (where the rcs files are kept). This information must be available to cvs for most commands to execute; if $CVSROOT is not set, or if you wish to override it for one invocation, you can supply it on the command line: ‘cvs -d cvsroot cvs_command...’ Once you have checked out a working directory, cvs stores the appropriate root (in the file ‘CVS/Root’), so normally you only need to worry about this when initially checking out a working directory.

$CVSEDITOR
$EDITOR

$VISUAL Specifies the program to use for recording log messages during commit. $CVSEDITOR overrides $EDITOR, which overrides $VISUAL. See Section 1.3.2 [Committing your changes], page 4 for more or Section A.4 [Global options], page 88 for alternative ways of specifying a log editor.

$PATH If $RCSBIN is not set, and no path is compiled into cvs, it will use $PATH to try to find all programs it uses.

$HOME
$HOMEPATH
$HOMEDRIVE

Used to locate the directory where the ‘.cvsrc’ file, and other such files, are searched. On Unix, cvs just checks for HOME. On Windows NT, the system will set HOMEDRIVE, for example to ‘d:’ and HOMEPATH, for example to ‘\joe’. On Windows 95, you’ll probably need to set HOMEDRIVE and HOMEPATH yourself.

$CVS_RSH Specifies the external program which cvs connects with, when :ext: access method is specified. see Section 2.9.2 [Connecting via rsh], page 20.

$CVS_SSH Specifies the external program which cvs connects with, when :extssh: access method is specified. see Section 2.9.2 [Connecting via rsh], page 20.

$CVS_SERVER

Used in client-server mode when accessing a remote repository using rsh. It specifies the name of the program to start on the server side (and any necessary arguments) when accessing a remote repository using the :ext:, :fork:, or :server: access methods. The default value for :ext: and :server: is cvs; the default value for :fork: is the name used to run the client. see Section 2.9.2 [Connecting via rsh], page 20

$CVS_PASSFILE

Used in client-server mode when accessing the cvs login server. Default value is ‘$HOME/.cvspass’. see Section 2.9.3.2 [Password authentication client], page 24

$CVS_CLIENT_PORT

Used in client-server mode to set the port to use when accessing the server via Kerberos, GSSAPI, or cvs’s password authentication protocol if the port is not specified in the CVSROOT. see Section 2.9 [Remote repositories], page 19

$CVS_RCMD_PORT

Used in client-server mode. If set, specifies the port number to be used when accessing the rcmd demon on the server side. (Currently not used for Unix clients).

$CVS_CLIENT_LOG

Used for debugging only in client-server mode. If set, everything sent to the server is logged into ‘$CVS_CLIENT_LOG.in’ and everything sent from the server

is logged into ‘$CVS_CLIENT_LOG.out’.
$CVS_SERVER_SLEEP

Used only for debugging the server side in client-server mode. If set, delays the start of the server child process the specified amount of seconds so that you can attach to it with a debugger.

$CVS_IGNORE_REMOTE_ROOT

For cvs 1.10 and older, setting this variable prevents cvs from overwriting the ‘CVS/Root’ file when the ‘-d’ global option is specified. Later versions of cvs

do not rewrite ‘CVS/Root’, so CVS_IGNORE_REMOTE_ROOT has no effect.

$COMSPEC Used under OS/2 only. It specifies the name of the command interpreter and defaults to cmd.exe.

$TMPDIR
$TMP

$TEMP Directory in which temporary files are located. The cvs server uses TMPDIR. See Section A.4 [Global options], page 88, for a description of how to specify this. Some parts of cvs will always use ‘/tmp’ (via the tmpnam function provided by the system). On Windows NT, TMP is used (via the _tempnam function provided by the system). The patch program which is used by the cvs client uses TMPDIR, and if it is not set, uses ‘/tmp’ (at least with GNU patch 2.1). Note that if your server Appendix D: All environment variables which affect CVS 157

and client are both running cvs 1.9.10 or later, cvs will not invoke an external patch program.

158 CVS—Concurrent Versions System v1.11.23

Appendix E Compatibility between CVS Versions

The repository format is compatible going back to cvs 1.3. But see Section 10.6.5 [Watches Compatibility], page 71, if you have copies of cvs 1.6 or older and you want to use the optional developer communication features.

The working directory format is compatible going back to cvs 1.5. It did change between cvs 1.3 and cvs 1.5. If you run cvs 1.5 or newer on a working directory checked out with cvs 1.3, cvs will convert it, but to go back to cvs 1.3 you need to check out a new working directory with cvs 1.3.

The remote protocol is interoperable going back to cvs 1.5, but no further (1.5 was the first official release with the remote protocol, but some older versions might still be floating around). In many cases you need to upgrade both the client and the server to take advantage of new features and bug fixes, however.

160 CVS—Concurrent Versions System v1.11.23

Appendix F Troubleshooting

If you are having trouble with cvs, this appendix may help. If there is a particular error message which you are seeing, then you can look up the message alphabetically. If not, you can look through the section on other problems to see if your problem is mentioned there.

F.1 Partial list of error messages

Here is a partial list of error messages that you may see from cvs. It is not a complete list— cvs is capable of printing many, many error messages, often with parts of them supplied by the operating system, but the intention is to list the common and/or potentially confusing error messages.

The messages are alphabetical, but introductory text such as ‘cvs update: ’ is not considered in ordering them.

In some cases the list includes messages printed by old versions of cvs (partly because users may not be sure which version of cvs they are using at any particular moment).

file:line: Assertion 'text' failed

The exact format of this message may vary depending on your system. It indicates a bug in cvs, which can be handled as described in Appendix H [BUGS], page 171.

cvs command: authorization failed: server host rejected access

This is a generic response when trying to connect to a pserver server which chooses not to provide a specific reason for denying authorization. Check that the username and password specified are correct and that the CVSROOT specified is allowed by ‘--allow-root’ in ‘inetd.conf’. See Section 2.9.3 [Password authenticated], page 21.

cvs command: conflict: removed file was modified by second party

This message indicates that you removed a file, and someone else modified it. To resolve the conflict, first run ‘cvs add file’. If desired, look at the other party’s modification to decide whether you still want to remove it. If you don’t want to remove it, stop here. If you do want to remove it, proceed with ‘cvs remove file’ and commit your removal.

cannot change permissions on temporary directory
     Operation not permitted

This message has been happening in a non-reproducible, occasional way when we run the client/server testsuite, both on Red Hat Linux 3.0.3 and 4.1. We haven’t been able to figure out what causes it, nor is it known whether it is specific to Linux (or even to this particular machine!). If the problem does occur on other unices, ‘Operation not permitted’ would be likely to read ‘Not owner’ or whatever the system in question uses for the unix EPERM error. If you have any information to add, please let us know as described in Appendix H [BUGS], page 171. If you experience this error while using cvs, retrying the operation which produced it should work fine.

cvs [server aborted]: Cannot check out files into the repository itself

The obvious cause for this message (especially for non-client/server cvs) is that the cvs root is, for example, ‘/usr/local/cvsroot’ and you try to check out files when you are in a subdirectory, such as ‘/usr/local/cvsroot/test’. However, there is a more subtle cause, which is that the temporary directory on the server is set to a subdirectory of the root (which is also not allowed). If this is the problem, set the temporary directory to somewhere else, for example ‘/var/tmp’; see TMPDIR in Appendix D [Environment variables], page 155, for how to set the temporary directory.

cannot commit files as 'root'
See ‘'root' is not allowed to commit files’.
cannot open CVS/Entries for reading: No such file or directory

This generally indicates a cvs internal error, and can be handled as with other cvs bugs (see Appendix H [BUGS], page 171). Usually there is a workaround— the exact nature of which would depend on the situation but which hopefully could be figured out.

cvs [init aborted]: cannot open CVS/Root: No such file or directory

This message is harmless. Provided it is not accompanied by other errors, the operation has completed successfully. This message should not occur with current versions of cvs, but it is documented here for the benefit of cvs 1.9 and older.

cvs server: cannot open /root/.cvsignore: Permission denied
cvs [server aborted]: can't chdir(/root): Permission denied

See Section F.2 [Connection], page 167.

cvs [checkout aborted]: cannot rename file file to CVS/,,file: Invalid
argument

This message has been reported as intermittently happening with cvs 1.9 on Solaris 2.5. The cause is unknown; if you know more about what causes it, let us know as described in Appendix H [BUGS], page 171.

cvs [command aborted]: cannot start server via rcmd

This, unfortunately, is a rather nonspecific error message which cvs 1.9 will print if you are running the cvs client and it is having trouble connecting to the server. Current versions of cvs should print a much more specific error message. If you get this message when you didn’t mean to run the client at all, you probably forgot to specify :local:, as described in Chapter 2 [Repository], page 7.

ci: file,v: bad diff output line: Binary files - and /tmp/T2a22651 differ

cvs 1.9 and older will print this message when trying to check in a binary file if rcs is not correctly installed. Re-read the instructions that came with your rcs distribution and the install file in the cvs distribution. Alternately, upgrade to a current version of cvs, which checks in files itself rather than via rcs.

cvs checkout: could not check out file

With cvs 1.9, this can mean that the co program (part of rcs) returned a failure. It should be preceded by another error message, however it has been observed without another error message and the cause is not well-understood. With the current version of cvs, which does not run co, if this message occurs without another error message, it is definitely a cvs bug (see Appendix H [BUGS], page 171).

cvs [login aborted]: could not find out home directory

This means that you need to set the environment variables that cvs uses to locate your home directory. See the discussion of HOME, HOMEDRIVE, and HOMEPATH in Appendix D [Environment variables], page 155.

cvs update: could not merge revision rev of file: No such file or directory

cvs 1.9 and older will print this message if there was a problem finding the rcsmerge program. Make sure that it is in your PATH, or upgrade to a current version of cvs, which does not require an external rcsmerge program.

cvs [update aborted]: could not patch file: No such file or directory

This means that there was a problem finding the patch program. Make sure that it is in your PATH. Note that despite appearances the message is not referring to whether it can find file. If both the client and the server are running a current version of cvs, then there is no need for an external patch program and you should not see this message. But if either client or server is running cvs 1.9, then you need patch.

cvs update: could not patch file; will refetch

This means that for whatever reason the client was unable to apply a patch that the server sent. The message is nothing to be concerned about, because inability to apply the patch only slows things down and has no effect on what cvs does.

dying gasps from server unexpected

There is a known bug in the server for cvs 1.9.18 and older which can cause this. For me, this was reproducible if I used the ‘-t’ global option. It was fixed by Andy Piper’s 14 Nov 1997 change to src/filesubr.c, if anyone is curious. If you see the message, you probably can just retry the operation which failed, or if you have discovered information concerning its cause, please let us know as described in Appendix H [BUGS], page 171.

end of file from server (consult above messages if any)

The most common cause for this message is if you are using an external rsh or ssh program and it exited with an error. In this case the rsh program should have printed a message, which will appear before the above message. For more information on setting up a cvs client and server, see Section 2.9 [Remote repositories], page 19.

cvs [update aborted]: EOF in key in RCS file file,v
cvs [checkout aborted]: EOF while looking for end of string in RCS file file,v

This means that there is a syntax error in the given rcs file. Note that this might be true even if rcs can read the file OK; cvs does more error checking of errors in the RCS file. That is why you may see this message when upgrading from cvs 1.9 to cvs 1.10. The likely cause for the original corruption is hardware, the operating system, or the like. Of course, if you find a case in which cvs seems to corrupting the file, by all means report it, (see Appendix H [BUGS], page 171). There are quite a few variations of this error message, depending on exactly where in the rcs file cvs finds the syntax error.

cvs commit: Executing 'mkmodules'

This means that your repository is set up for a version of cvs prior to cvs 1.8. When using cvs 1.8 or later, the above message will be preceded by

     cvs commit: Rebuilding administrative file database

If you see both messages, the database is being rebuilt twice, which is unnecessary but harmless. If you wish to avoid the duplication, and you have no versions of cvs 1.7 or earlier in use, remove -i mkmodules every place it appears in your modules file. For more information on the modules file, see Section C.1 [modules], page 139.

missing author

Typically this can happen if you created an RCS file with your username set to empty. cvs will, bogusly, create an illegal RCS file with no value for the author field. The solution is to make sure your username is set to a non-empty value and re-create the RCS file.

cvs [checkout aborted]: no such tag tag

This message means that cvs isn’t familiar with the tag tag. Usually this means that you have mistyped a tag name; however there are (relatively obscure) cases in which cvs will require you to try a few other cvs commands involving that tag, before you find one which will cause cvs to update the ‘val-tags’ file; see discussion of val-tags in Section 2.2.2 [File permissions], page 9. You only need to worry about this once for a given tag; when a tag is listed in ‘val-tags’, it stays there. Note that using ‘-f’ to not require tag matches does not override this check; see Section A.5 [Common options], page 90.

cvs [command aborted]: out of memory

There is insufficient (virtual) memory available to continue. In client/server mode, the problem is almost certainly on the server rather than the client; see Section 2.9.1 [Server requirements], page 20 for memory estimates. Many systems have limits on the amount of virtual memory that a single process can use, so a process can run out of virtual memory long before the system itself has run out. The method of increasing the per-process limits varies depending on the operating system.

*PANIC* administration files missing

This typically means that there is a directory named cvs but it does not contain the administrative files which cvs puts in a CVS directory. If the problem is that you created a CVS directory via some mechanism other than cvs, then the answer is simple, use a name other than cvs. If not, it indicates a cvs bug (see Appendix H [BUGS], page 171).

rcs error: Unknown option: -x,v/

This message will be followed by a usage message for rcs. It means that you have an old version of rcs (probably supplied with your operating system), as well as an old version of cvs. cvs 1.9.18 and earlier only work with rcs version 5 and later; current versions of cvs do not run rcs programs.

cvs [server aborted]: received broken pipe signal

This message can be caused by a loginfo program that fails to read all of the log information from its standard input. If you find it happening in any other circumstances, please let us know as described in Appendix H [BUGS], page 171.

'root' is not allowed to commit files

When committing a permanent change, cvs makes a log entry of who committed the change. If you are committing the change logged in as "root" (not under "su" or other root-priv giving program), cvs cannot determine who is actually making the change. As such, by default, cvs disallows changes to be committed by users logged in as "root". (You can disable this option by passing the --enable-rootcommit option to ‘configure’ and recompiling cvs. On some systems this means editing the appropriate ‘config.h’ file before building cvs.)

Terminated with fatal signal 11

This message usually indicates that cvs (the server, if you’re using client/server mode) has run out of (virtual) memory. Although cvs tries to catch the error and issue a more meaningful message, there are many circumstances where that is not possible. If you appear to have lots of memory available to the system, the problem is most likely that you’re running into a system-wide limit on the amount of memory a single process can use or a similar process-specific limit. The mechanisms for displaying and setting such limits vary from system to system, so you’ll have to consult an expert for your particular system if you don’t know how to do that.

Too many arguments!

This message is typically printed by the ‘log.pl’ script which is in the ‘contrib’ directory in the cvs source distribution. In some versions of cvs, ‘log.pl’ has been part of the default cvs installation. The ‘log.pl’ script gets called from the ‘loginfo’ administrative file. Check that the arguments passed in ‘loginfo’ match what your version of ‘log.pl’ expects. In particular, the ‘log.pl’ from cvs 1.3 and older expects the log file as an argument whereas the ‘log.pl’ from cvs 1.5 and newer expects the log file to be specified with a ‘-f’ option. Of course, if you don’t need ‘log.pl’ you can just comment it out of ‘loginfo’.

cvs [update aborted]: unexpected EOF reading file,v
See ‘EOF in key in RCS file’.
cvs [login aborted]: unrecognized auth response from server

This message typically means that the server is not set up properly. For example, if ‘inetd.conf’ points to a nonexistent cvs executable. To debug it further, find the log file which inetd writes (‘/var/log/messages’ or whatever inetd uses on your system). For details, see Section F.2 [Connection], page 167, and Section 2.9.3.1 [Password authentication server], page 21.

cvs commit: Up-to-date check failed for `file'

This means that someone else has committed a change to that file since the last time that you did a cvs update. So before proceeding with your cvs commit you need to cvs update. cvs will merge the changes that you made and the changes that the other person made. If it does not detect any conflicts it will report ‘M file’ and you are ready to cvs commit. If it detects conflicts it will print a message saying so, will report ‘C file’, and you need to manually resolve the conflict. For more details on this process see Section 10.3 [Conflicts example], page 65.

Usage: diff3 [-exEX3 [-i | -m] [-L label1 -L label3]] file1 file2 file3
     Only one of [exEX3] allowed

This indicates a problem with the installation of diff3 and rcsmerge. Specifically rcsmerge was compiled to look for GNU diff3, but it is finding unix diff3 instead. The exact text of the message will vary depending on the system. The simplest solution is to upgrade to a current version of cvs, which does not rely on external rcsmerge or diff3 programs.

warning: unrecognized response `text' from cvs server

If text contains a valid response (such as ‘ok’) followed by an extra carriage return character (on many systems this will cause the second part of the message to overwrite the first part), then it probably means that you are using the ‘:ext:’ access method with a version of rsh, such as most non-unix rsh versions, which does not by default provide a transparent data stream. In such cases you probably want to try ‘:server:’ instead of ‘:ext:’. If text is something else, this may signify a problem with your cvs server. Double-check your installation against the instructions for setting up the cvs server.

cvs commit: [time] waiting for user's lock in directory

This is a normal message, not an error. See Section 10.5 [Concurrency], page 67, for more details.

cvs commit: warning: editor session failed

This means that the editor which cvs is using exits with a nonzero exit status. Some versions of vi will do this even when there was not a problem editing the file. If so, point the CVSEDITOR environment variable to a small script such as:

     #!/bin/sh
     vi $*
     exit 0
cvs update: warning: file was lost

This means that the working copy of file has been deleted but it has not been removed from cvs. This is nothing to be concerned about, the update will just recreate the local file from the repository. (This is a convenient way to discard local changes to a file: just delete it and then run cvs update.)

cvs update: warning: file is not (any longer) pertinent

This means that the working copy of file has been deleted, it has not been removed from cvs in the current working directory, but it has been removed from cvs in some other working directory. This is nothing to be concerned about, the update would have removed the local file anyway.

F.2 Trouble making a connection to a CVS server

This section concerns what to do if you are having trouble making a connection to a cvs server. If you are running the cvs command line client running on Windows, first upgrade the client to cvs 1.9.12 or later. The error reporting in earlier versions provided much less information about what the problem was. If the client is non-Windows, cvs 1.9 should be fine.

If the error messages are not sufficient to track down the problem, the next steps depend largely on which access method you are using.

:ext: Try running the rsh program from the command line. For example: "rsh servername cvs -v" should print cvs version information. If this doesn’t work, you need to fix it before you can worry about cvs problems.

:server: You don’t need a command line rsh program to use this access method, but if you have an rsh program around, it may be useful as a debugging tool. Follow the directions given for :ext:.

:pserver:

Errors along the lines of "connection refused" typically indicate that inetd isn’t even listening for connections on port 2401 whereas errors like "connection reset by peer", "received broken pipe signal", "recv() from server: EOF", or "end of file from server" typically indicate that inetd is listening for connections but is unable to start cvs (this is frequently caused by having an incorrect path in ‘inetd.conf’ or by firewall software rejecting the connection). "unrecognized auth response" errors are caused by a bad command line in ‘inetd.conf’, typically an invalid option or forgetting to put the ‘pserver’ command at the end of the line. Another less common problem is invisible control characters that your editor "helpfully" added without you noticing. One good debugging tool is to "telnet servername 2401". After connecting, send any text (for example "foo" followed by return). If cvs is working correctly, it will respond with

     cvs [pserver aborted]: bad auth protocol start: foo

If instead you get:

     Usage: cvs [cvs-options] command [command-options-and-arguments]
     ...

then you’re missing the ‘pserver’ command at the end of the line in ‘inetd.conf’; check to make sure that the entire command is on one line and that it’s complete. Likewise, if you get something like:

     Unknown command: `pserved'
     CVS commands are:
             add          Add a new file/directory to the repository

     ...

then you’ve misspelled ‘pserver’ in some way. If it isn’t obvious, check for invisible control characters (particularly carriage returns) in ‘inetd.conf’. If it fails to work at all, then make sure inetd is working right. Change the invocation in ‘inetd.conf’ to run the echo program instead of cvs. For example:

     2401  stream  tcp  nowait  root /bin/echo echo hello

After making that change and instructing inetd to re-read its configuration file, "telnet servername 2401" should show you the text hello and then the server should close the connection. If this doesn’t work, you need to fix it before you can worry about cvs problems. On AIX systems, the system will often have its own program trying to use port 2401. This is AIX’s problem in the sense that port 2401 is registered for use with cvs. I hear that there is an AIX patch available to address this problem. Another good debugging tool is the ‘-d’ (debugging) option to inetd. Consult your system documentation for more information. If you seem to be connecting but get errors like:

     cvs server: cannot open /root/.cvsignore: Permission denied
     cvs [server aborted]: can't chdir(/root): Permission denied

then you probably haven’t specified ‘-f’ in ‘inetd.conf’. (In releases prior to cvs 1.11.1, this problem can be caused by your system setting the $HOME environment variable for programs being run by inetd. In this case, you can either have inetd run a shell script that unsets $HOME and then runs cvs, or you can use env to run cvs with a pristine environment.) If you can connect successfully for a while but then can’t, you’ve probably hit inetd’s rate limit. (If inetd receives too many requests for the same service in a short period of time, it assumes that something is wrong and temporarily disables the service.) Check your inetd documentation to find out how to adjust the rate limit (some versions of inetd have a single rate limit, others allow you to set the limit for each service separately.)

F.3 Other common problems

Here is a list of problems which do not fit into the above categories. They are in no particular order.

• On Windows, if there is a 30 second or so delay when you run a cvs command, it may mean that you have your home directory set to ‘C:/’, for example (see HOMEDRIVE and HOMEPATH in Appendix D [Environment variables], page 155). cvs expects the home directory to not end in a slash, for example ‘C:’ or ‘C:\cvs’.

• If you are running cvs 1.9.18 or older, and cvs update finds a conflict and tries to merge, as described in Section 10.3 [Conflicts example], page 65, but doesn’t tell you there were conflicts, then you may have an old version of rcs. The easiest solution probably is to upgrade to a current version of cvs, which does not rely on external rcs programs.

Appendix G Credits

Roland Pesch, then of Cygnus Support <roland@wrs.com> wrote the manual pages which were distributed with cvs 1.3. Much of their text was copied into this manual. He also read an early draft of this manual and contributed many ideas and corrections.

The mailing-list info-cvs is sometimes informative. I have included information from postings made by the following persons: David G. Grubbs <dgg@think.com>.

Some text has been extracted from the man pages for rcs.

The cvs faq by David G. Grubbs has provided useful material. The faq is no longer maintained, however, and this manual is about the closest thing there is to a successor (with respect to documenting how to use cvs, at least).

In addition, the following persons have helped by telling me about mistakes I’ve made:

     Roxanne Brunskill <rbrunski@datap.ca>,
     Kathy Dyer <dyer@phoenix.ocf.llnl.gov>,
     Karl Pingle <pingle@acuson.com>,
     Thomas A Peterson <tap@src.honeywell.com>,
     Inge Wallin <ingwa@signum.se>,
     Dirk Koschuetzki <koschuet@fmi.uni-passau.de>
     and Michael Brown <brown@wi.extrel.com>.

The list of contributors here is not comprehensive; for a more complete list of who has contributed to this manual see the file ‘doc/ChangeLog’ in the cvs source distribution.

170 CVS—Concurrent Versions System v1.11.23

Appendix H Dealing with bugs in CVS or this manual

Neither cvs nor this manual is perfect, and they probably never will be. If you are having trouble using cvs, or think you have found a bug, there are a number of things you can do about it. Note that if the manual is unclear, that can be considered a bug in the manual, so these problems are often worth doing something about as well as problems with cvs itself.

• If you want someone to help you and fix bugs that you report, there are companies

which will do that for a fee. One such company is:

     Ximbiot LLC
     Suite 230
     200 Diversion St.
     Rochester Hills, MI  48307-6636
     USA
     Email: info@ximbiot.com
     Phone: (248) 835-1260
     Fax:   (248) 835-1263
     http://ximbiot.com/

• If you got cvs through a distributor, such as an operating system vendor or a vendor of freeware cd-roms, you may wish to see whether the distributor provides support. Often, they will provide no support or minimal support, but this may vary from distributor to distributor.

• If you have the skills and time to do so, you may wish to fix the bug yourself. If you wish to submit your fix for inclusion in future releases of cvs, see the file hacking in the cvs source distribution. It contains much more information on the process of submitting fixes.

• There may be resources on the net which can help. A good place to start is:

     http://cvs.nongnu.org/

If you are so inspired, increasing the information available on the net is likely to be appreciated. For example, before the standard cvs distribution worked on Windows 95, there was a web page with some explanation and patches for running cvs on Windows 95, and various people helped out by mentioning this page on mailing lists or newsgroups when the subject came up.

• It is also possible to report bugs to bug-cvs@nongnu.org. Note that someone may or may not want to do anything with your bug report—if you need a solution consider one of the options mentioned above. People probably do want to hear about bugs which are particularly severe in consequences and/or easy to fix, however. You can also increase your odds by being as clear as possible about the exact nature of the bug and any other relevant information. The way to report bugs is to send email to bug-cvs@nongnu.org. Note that submissions to bug-cvs@nongnu.org may be distributed under the terms of the gnu Public License, so if you don’t like this, don’t submit them. There is usually no justification for sending mail directly to one of the cvs maintainers rather than to bug-cvs@nongnu.org; those maintainers who want to hear about such bug reports read bug-cvs@nongnu.org. Also note that sending a bug report to other mailing lists or newsgroups is not a substitute for sending it to bug-cvs@nongnu.org. It is fine to discuss cvs bugs on whatever forum you prefer, but there are not necessarily any maintainers reading bug reports sent anywhere except bug-cvs@nongnu.org.

People often ask if there is a list of known bugs or whether a particular bug is a known one. The file bugs in the cvs source distribution is one list of known bugs, but it doesn’t necessarily try to be comprehensive. Perhaps there will never be a comprehensive, detailed list of known bugs.

Index 173

Index

!

!, in modules file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141

#

#cvs.lock, removing . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 #cvs.lock, technical details . . . . . . . . . . . . . . . . . . . . 12 #cvs.rfl, and backups . . . . . . . . . . . . . . . . . . . . . . . . . . 18 #cvs.rfl, removing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 #cvs.rfl, technical details . . . . . . . . . . . . . . . . . . . . . . 12 #cvs.tfl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 #cvs.wfl, removing . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 #cvs.wfl, technical details . . . . . . . . . . . . . . . . . . . . . 12

&

&, in modules file . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140

-

-a, in modules file . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 -d, in modules file . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 -e, in modules file . . . . . . . . . . . . . . . . . . . . . . . . 141, 142 -j (merging branches) . . . . . . . . . . . . . . . . . . . . . . . . . . 45 -j (merging branches), and keyword substitution

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 -k (keyword substitution) . . . . . . . . . . . . . . . . . . . . . . 77 -kk, to avoid conflicts during a merge . . . . . . . . . . 47 -o, in modules file. . . . . . . . . . . . . . . . . . . . . . . . 141, 142 -s, in modules file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 -t, in modules file . . . . . . . . . . . . . . . . . . . . . . . . 141, 142

.

.# files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 .bashrc, setting CVSROOT in . . . . . . . . . . . . . . . . . . 7 .cshrc, setting CVSROOT in . . . . . . . . . . . . . . . . . . . . 7 .cvsrc file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 .profile, setting CVSROOT in . . . . . . . . . . . . . . . . . . . 7 .tcshrc, setting CVSROOT in . . . . . . . . . . . . . . . . . . . 7

/

/usr/local/cvsroot, as example repository . . . . . . . 7

:

:ext:, setting up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 :ext:, troubleshooting . . . . . . . . . . . . . . . . . . . . . . . . . 167 :extssh:, setting up . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 :fork:, setting up. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 :gserver:, setting up. . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 :kserver:, setting up. . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 :local:, setting up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 :pserver:, setting up. . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 :pserver:, troubleshooting . . . . . . . . . . . . . . . . . . . . . 167 :server:, setting up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 :server:, troubleshooting . . . . . . . . . . . . . . . . . . . . . . 167

<

<<<<<<< . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

=

======= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

>

>>>>>>> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

files (VMS) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125

A

Abandoning work. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 Access a branch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 add (subcommand) . . . . . . . . . . . . . . . . . . . . . . . . 53, 93 Adding a tag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 Adding files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Admin (subcommand) . . . . . . . . . . . . . . . . . . . . . . . . . 94 Administrative files (intro). . . . . . . . . . . . . . . . . . . . . 16 Administrative files (reference). . . . . . . . . . . . . . . . 139 Administrative files, editing them . . . . . . . . . . . . . . 17 Alias modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 ALL in commitinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 Ampersand modules . . . . . . . . . . . . . . . . . . . . . . . . . . 140 annotate (subcommand) . . . . . . . . . . . . . . . . . . . . . . . 98 Atomic transactions, lack of . . . . . . . . . . . . . . . . . . . 68 Attic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Authenticated client, using . . . . . . . . . . . . . . . . . . . . 24 Authenticating server, setting up. . . . . . . . . . . . . . . 21 Authentication, stream . . . . . . . . . . . . . . . . . . . . . . . . 88 Author keyword . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Automatically ignored files . . . . . . . . . . . . . . . . . . . 150 Avoiding editor invocation . . . . . . . . . . . . . . . . . . . . . 92

B

174 CVS—Concurrent Versions System v1.11.23

Backing up, repository. . . . . . . . . . . . . . . . . . . . . . . . . 18 Base directory, in CVS directory . . . . . . . . . . . . . . . 16 BASE, as reserved tag name . . . . . . . . . . . . . . . . . . . 34 BASE, special tag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 Baserev file, in CVS directory. . . . . . . . . . . . . . . . . . 16 Baserev.tmp file, in CVS directory . . . . . . . . . . . . . 16 Bill of materials . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 Binary files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 Branch merge example. . . . . . . . . . . . . . . . . . . . . . . . . 45 Branch number . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33, 43 Branch tags, deleting . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Branch tags, moving . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Branch, accessing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 Branch, check out . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 Branch, creating a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 Branch, identifying . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 Branch, retrieving . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 Branch, vendor- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 Branches motivation . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 Branches, copying changes between . . . . . . . . . . . . 41 Branches, sticky . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 Branching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 Bringing a file up to date . . . . . . . . . . . . . . . . . . . . . . 64 Bugs in this manual or CVS . . . . . . . . . . . . . . . . . . 171 Bugs, reporting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 Builds . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

C

Changes, copying between branches . . . . . . . . . . . . 41 Changing a log message. . . . . . . . . . . . . . . . . . . . . . . . 95 Check out a branch . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 Checked out copy, keeping . . . . . . . . . . . . . . . . . . . . 149 Checking out source . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 checkout (subcommand) . . . . . . . . . . . . . . . . . . . . . . . 99 Checkout program . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Checkout, as term for getting ready to edit. . . . . 70 Checkout, example. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 checkoutlist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 Choosing, reserved or unreserved checkouts …. 71 Cleaning up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Client/Server Operation . . . . . . . . . . . . . . . . . . . . . . . 19 Client/Server Operation, port specification …. 19,

21 co (subcommand) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 Command reference . . . . . . . . . . . . . . . . . . . . . . . . . . 127 Command structure . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 Comment leader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 commit (subcommand) . . . . . . . . . . . . . . . . . . . . . . . 101 ‘commitinfo’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 ‘commitinfo’, command environment . . . . . . . . . 144 ‘commitinfo’, working directory . . . . . . . . . . . . . . 144 Commits, precommit verification of . . . . . . . . . . . 144 Committing changes to files . . . . . . . . . . . . . . . . . . . . . 4 Committing, administrative support files. . . . . . 143 Committing, when to . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Common options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 Common syntax of info files . . . . . . . . . . . . . . . . . . 143 Compatibility, between CVS versions . . . . . . . . . 159 Compression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90, 127 COMSPEC, environment variable. . . . . . . . . . . . . 156 config, in CVSROOT . . . . . . . . . . . . . . . . . . . . . . . . . 153 Conflict markers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 Conflict resolution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 Conflicts (merge example) . . . . . . . . . . . . . . . . . . . . . 66 Contributors (CVS program). . . . . . . . . . . . . . . . . . . . 1 Contributors (manual). . . . . . . . . . . . . . . . . . . . . . . . 169 Copying a repository . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Copying changes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 Correcting a log message. . . . . . . . . . . . . . . . . . . . . . . 95 Creating a branch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 Creating a project . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 Creating a repository . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Credits (CVS program) . . . . . . . . . . . . . . . . . . . . . . . . . 1 Credits (manual) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 CVS 1.6, and watches . . . . . . . . . . . . . . . . . . . . . . . . . 71 CVS command structure. . . . . . . . . . . . . . . . . . . . . . . 87 CVS directory, in repository . . . . . . . . . . . . . . . . . . . 11 CVS directory, in working directory . . . . . . . . . . . . 13 CVS passwd file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 CVS, history of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 CVS, introduction to . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 CVS, versions of. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 CVS/Base directory . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 CVS/Baserev file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 CVS/Baserev.tmp file . . . . . . . . . . . . . . . . . . . . . . . . . 16 CVS/Entries file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 CVS/Entries.Backup file . . . . . . . . . . . . . . . . . . . . . . . 15 CVS/Entries.Log file . . . . . . . . . . . . . . . . . . . . . . . . . . 15 CVS/Entries.Static file . . . . . . . . . . . . . . . . . . . . . . . . 16 CVS/Notify file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 CVS/Notify.tmp file . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 CVS/Repository file . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 CVS/Root file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 CVS/Tag file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 CVS/Template file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 CVS CLIENT LOG, environment variable …. 156 CVS CLIENT PORT . . . . . . . . . . . . . . . . . . . . . . . . . 26 CVS IGNORE REMOTE ROOT, environment

variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 CVS PASSFILE, environment variable . . . . . . . . . 25 CVS RCMD PORT, environment variable …. 156 CVS RSH, environment variable . . . . . . . . . . . . . . 155 CVS SERVER, and :fork:. . . . . . . . . . . . . . . . . . . . . . 27 CVS SERVER, environment variable . . . . . . . . . . 21 CVS SERVER SLEEP, environment variable

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 CVS SSH, environment variable . . . . . . . . . . . . . . 155 CVS USER, environment variable . . . . . . . . . . . . 153 cvsadmin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 CVSEDITOR, environment variable . . . . . . . . 4, 155 CVSEDITOR, internal variable . . . . . . . . . . . . . . . 152 cvsignore (admin file), global . . . . . . . . . . . . . . . . . 150 CVSIGNORE, environment variable . . . . . . . . . . 155 CVSREAD, environment variable . . . . . . . . . . . . . 155 CVSREAD, overriding. . . . . . . . . . . . . . . . . . . . . . . . . 90 cvsroot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 CVSROOT (file) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 CVSROOT, environment variable . . . . . . . . . . . . . . . 7 CVSROOT, internal variable . . . . . . . . . . . . . . . . . 152 CVSROOT, module name . . . . . . . . . . . . . . . . . . . . . 16 CVSROOT, multiple repositories . . . . . . . . . . . . . . 17 CVSROOT, overriding. . . . . . . . . . . . . . . . . . . . . . . . . 89 CVSROOT, storage of files . . . . . . . . . . . . . . . . . . . . 13 CVSROOT/config . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 CVSROOT/Emptydir directory. . . . . . . . . . . . . . . . 14 CVSROOT/val-tags file, and read-only access to

projects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 CVSROOT/val-tags file, forcing tags into . . . . . 164 CVSUMASK, environment variable . . . . . . . . . . . . 10 cvswrappers (admin file). . . . . . . . . . . . . . . . . . . . . . 142 CVSWRAPPERS, environment variable . . 142, 155

D

Date keyword . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Dates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 Dead state . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Decimal revision number. . . . . . . . . . . . . . . . . . . . . . . 33 DEFAULT in commitinfo . . . . . . . . . . . . . . . . . . . . . 144 DEFAULT in editinfo . . . . . . . . . . . . . . . . . . . . . . . . 147 DEFAULT in ‘verifymsg’ . . . . . . . . . . . . . . . . . . . . 145 Defining a module . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 Defining modules (intro) . . . . . . . . . . . . . . . . . . . . . . . 16 Defining modules (reference manual) . . . . . . . . . . 139 Deleting branch tags. . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Deleting files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 Deleting revisions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 Deleting sticky tags. . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 Deleting tags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Descending directories . . . . . . . . . . . . . . . . . . . . . . . . . 51 Device nodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 Diff. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 diff (subcommand) . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Differences, merging . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 Directories, moving . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 Directories, removing . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Directory, descending . . . . . . . . . . . . . . . . . . . . . . . . . . 51 Disjoint repositories . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Distributing log messages . . . . . . . . . . . . . . . . . . . . . 148 driver.c (merge example) . . . . . . . . . . . . . . . . . . . . . . 65

E

edit (subcommand) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 editinfo (admin file) . . . . . . . . . . . . . . . . . . . . . . . . . . 146 Editing administrative files . . . . . . . . . . . . . . . . . . . . 17 Editing the modules file . . . . . . . . . . . . . . . . . . . . . . . 31 Editor, avoiding invocation of . . . . . . . . . . . . . . . . . . 92 EDITOR, environment variable . . . . . . . . . . . . 4, 155 EDITOR, internal variable . . . . . . . . . . . . . . . . . . . 152 EDITOR, overriding . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 Editor, specifying per module. . . . . . . . . . . . . . . . . 146 editors (subcommand) . . . . . . . . . . . . . . . . . . . . . . . . . 71 emerge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 Emptydir, in CVSROOT directory . . . . . . . . . . . . . 14 Encryption . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 Entries file, in CVS directory . . . . . . . . . . . . . . . . . . 14 Entries.Backup file, in CVS directory . . . . . . . . . . 15 Entries.Log file, in CVS directory . . . . . . . . . . . . . . 15 Entries.Static file, in CVS directory . . . . . . . . . . . . 16 Environment variables . . . . . . . . . . . . . . . . . . . . . . . . 155 environment variables, passed to administrative

files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 Errors, reporting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 Example of a work-session . . . . . . . . . . . . . . . . . . . . . . 3 Example of merge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 Example, branch merge . . . . . . . . . . . . . . . . . . . . . . . . 45 Excluding directories, in modules file . . . . . . . . . 141 Exit status, of commitinfo . . . . . . . . . . . . . . . . . . . . 144 Exit status, of CVS. . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 Exit status, of editor . . . . . . . . . . . . . . . . . . . . . . . . . 166 Exit status, of ‘verifymsg’. . . . . . . . . . . . . . . . . . . . 145 export (subcommand) . . . . . . . . . . . . . . . . . . . . . . . . 112 Export program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141

F

Fetching source . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 File had conflicts on merge . . . . . . . . . . . . . . . . . . . . 64 File locking. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 File permissions, general . . . . . . . . . . . . . . . . . . . . . . . . 9 File permissions, Windows-specific . . . . . . . . . . . . . 10 File status. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 Files, moving . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Files, reference manual . . . . . . . . . . . . . . . . . . . . . . . 139 Fixing a log message. . . . . . . . . . . . . . . . . . . . . . . . . . . 95 Forcing a tag match . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 fork, access method . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 Form for log message . . . . . . . . . . . . . . . . . . . . . . . . . 149 Format of CVS commands . . . . . . . . . . . . . . . . . . . . . 87

G

Getting started . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Getting the source . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Global cvsignore. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 Global options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 Group, UNIX file permissions, in repository . . . . . 9 gserver (client/server connection method), port

specification . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 21 GSSAPI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Gzip. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90, 127

H

Hard links . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 HEAD, as reserved tag name . . . . . . . . . . . . . . . . . . 34 HEAD, special tag. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 Header keyword . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 history (subcommand). . . . . . . . . . . . . . . . . . . . . . . . 113 History browsing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 History file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 History files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 History of CVS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 HOME, environment variable . . . . . . . . . . . . . . . . . 155 HOMEDRIVE, environment variable . . . . . . . . . 155 HOMEPATH, environment variable . . . . . . . . . . . 155

I

Id keyword . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Ident (shell command) . . . . . . . . . . . . . . . . . . . . . . . . . 76 Identifying a branch . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 Identifying files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Ignored files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 IgnoreUnknownConfigKeys, in CVSROOT/config

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 Ignoring files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 import (subcommand) . . . . . . . . . . . . . . . . . . . . . . . . 115 Importing files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 Importing files, from other version control systems

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 Importing modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 inetd, configuring for pserver. . . . . . . . . . . . . . . . . . . 21 Info files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 Info files (syntax). . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 Info files, security. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 Informing others. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 init (subcommand) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Installed images (VMS) . . . . . . . . . . . . . . . . . . . . . . . . 10 Internal variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 Introduction to CVS . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Invoking CVS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 Isolation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

J

Join . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

K

Keeping a checked out copy. . . . . . . . . . . . . . . . . . . 149 Kerberos, using :gserver: . . . . . . . . . . . . . . . . . . . . . . . 26 Kerberos, using :kserver: . . . . . . . . . . . . . . . . . . . . . . . 26 Kerberos, using kerberized rsh . . . . . . . . . . . . . . . . . 21 Keyword expansion . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Keyword List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Keyword substitution . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Keyword substitution, and merging . . . . . . . . . . . . 47 Keyword substitution, changing modes . . . . . . . . . 77 Kflag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 kinit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Known bugs in this manual or CVS . . . . . . . . . . . 172 kserver (client/server connection method), port

specification . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 21

L

Layout of repository . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Left-hand options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 Linear development . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 Link, symbolic, importing . . . . . . . . . . . . . . . . . . . . 116 List, mailing list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Locally Added . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 Locally Modified. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 Locally Removed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 LockDir, in CVSROOT/config . . . . . . . . . . . . . . . . 153 Locker keyword. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Locking files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 Locks, cvs, and backups . . . . . . . . . . . . . . . . . . . . . . . 18 Locks, cvs, introduction . . . . . . . . . . . . . . . . . . . . . . . 67 Locks, cvs, technical details . . . . . . . . . . . . . . . . . . . . 12 log (subcommand). . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 Log information, saving. . . . . . . . . . . . . . . . . . . . . . . 152 Log keyword. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Log message entry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Log message template . . . . . . . . . . . . . . . . . . . . . . . . 149 Log message, correcting. . . . . . . . . . . . . . . . . . . . . . . . 95 Log message, verifying . . . . . . . . . . . . . . . . . . . . . . . . 144 Log messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 Log messages, editing. . . . . . . . . . . . . . . . . . . . . . . . . 146 LogHistory, in CVSROOT/config . . . . . . . . . . . . . 154 Login (subcommand) . . . . . . . . . . . . . . . . . . . . . . . . . . 24 loginfo (admin file) . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 LOGNAME, environment variable . . . . . . . . . . . . 153 Logout (subcommand). . . . . . . . . . . . . . . . . . . . . . . . . 25

M

Mail, automatic mail on commit . . . . . . . . . . . . . . . 67 Mailing list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Mailing log messages . . . . . . . . . . . . . . . . . . . . . . . . . 148 Main trunk and branches . . . . . . . . . . . . . . . . . . . . . . 41 make . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 Many repositories. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Markers, conflict. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 Merge, an example . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 Merge, branch example . . . . . . . . . . . . . . . . . . . . . . . . 45 Merging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 Merging a branch. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 Merging a file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 Merging two revisions. . . . . . . . . . . . . . . . . . . . . . . . . . 46 Merging, and keyword substitution. . . . . . . . . . . . . 47 mkmodules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 Modifications, copying between branches . . . . . . . 41 Module status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Module, defining. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 Modules (admin file) . . . . . . . . . . . . . . . . . . . . . . . . . 139 Modules file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Modules file program options . . . . . . . . . . . . . . . . . 142 Modules file, changing . . . . . . . . . . . . . . . . . . . . . . . . . 31 modules.db . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 modules.dir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 modules.pag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Motivation for branches. . . . . . . . . . . . . . . . . . . . . . . . 41 Moving a repository . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Moving branch tags. . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Moving directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 Moving files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Moving tags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Multiple developers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 Multiple repositories . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

N

Name keyword . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Name, symbolic (tag) . . . . . . . . . . . . . . . . . . . . . . . . . . 34 Needs Checkout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 Needs Merge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 Needs Patch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 Newsgroups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 notify (admin file) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 Notify file, in CVS directory . . . . . . . . . . . . . . . . . . . 16 Notify.tmp file, in CVS directory. . . . . . . . . . . . . . . 16 Number, branch . . . . . . . . . . . . . . . . . . . . . . . . . . . 33, 43 Number, revision- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

O

Option defaults . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 Options, global . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 Options, in modules file . . . . . . . . . . . . . . . . . . . . . . 141 Outdating revisions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 Overlap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 Overriding CVSREAD. . . . . . . . . . . . . . . . . . . . . . . . . 90 Overriding CVSROOT. . . . . . . . . . . . . . . . . . . . . . . . . 89 Overriding EDITOR. . . . . . . . . . . . . . . . . . . . . . . . . . . 89 Overriding RCSBIN . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 Overriding TMPDIR . . . . . . . . . . . . . . . . . . . . . . . . . . 89 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Ownership, saving in CVS . . . . . . . . . . . . . . . . . . . . . 85

P

Parallel repositories. . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 passwd (admin file) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Password client, using . . . . . . . . . . . . . . . . . . . . . . . . . 24 Password server, setting up . . . . . . . . . . . . . . . . . . . . 21 PATH, environment variable . . . . . . . . . . . . . . . . . . 155 Per-directory sticky tags/dates. . . . . . . . . . . . . . . . . 16 Per-module editor . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 Permissions, general . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Permissions, saving in CVS . . . . . . . . . . . . . . . . . . . . 85 Permissions, Windows-specific . . . . . . . . . . . . . . . . . 10 Policy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 port, specifying for remote repositories . . . . . 19, 21 Precommit checking . . . . . . . . . . . . . . . . . . . . . . . . . . 144 pserver (client/server connection method), port

specification . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 21 pserver (subcommand). . . . . . . . . . . . . . . . . . . . . . . . . 21 PVCS, importing files from . . . . . . . . . . . . . . . . . . . . 30

R

RCS history files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 RCS revision numbers . . . . . . . . . . . . . . . . . . . . . . . . . 34 RCS, importing files from. . . . . . . . . . . . . . . . . . . . . . 30 RCS-style locking. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 RCSBIN, in CVSROOT/config . . . . . . . . . . . . . . . 153 RCSBIN, internal variable . . . . . . . . . . . . . . . . . . . . 152 RCSBIN, overriding . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 RCSfile keyword . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 rcsinfo (admin file) . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 rdiff (subcommand) . . . . . . . . . . . . . . . . . . . . . . . . . . 118 Read-only files, and -r . . . . . . . . . . . . . . . . . . . . . . . . . 89 Read-only files, and CVSREAD. . . . . . . . . . . . . . . 155 Read-only files, and watches . . . . . . . . . . . . . . . . . . . 68 Read-only files, in repository . . . . . . . . . . . . . . . . . . . . 9 Read-only mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 Read-only repository access . . . . . . . . . . . . . . . . . . . . 27 readers (admin file) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 Recursive (directory descending) . . . . . . . . . . . . . . . 51 Reference manual (files) . . . . . . . . . . . . . . . . . . . . . . 139 Reference manual for variables . . . . . . . . . . . . . . . . 155 Reference, commands. . . . . . . . . . . . . . . . . . . . . . . . . 127 Regular expression syntax . . . . . . . . . . . . . . . . . . . . 143 Regular modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 release (subcommand) . . . . . . . . . . . . . . . . . . . . . . . . 120 Releases, revisions and versions . . . . . . . . . . . . . . . . 33 Releasing your working copy . . . . . . . . . . . . . . . . . . . . 5 Remote repositories. . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Remote repositories, port specification. . . . . . 19, 21 remove (subcommand). . . . . . . . . . . . . . . . . . . . . . . . 121 Remove (subcommand) . . . . . . . . . . . . . . . . . . . . . . . . 54 Removing a change . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 Removing branch tags . . . . . . . . . . . . . . . . . . . . . . . . . 37 Removing directories . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Removing files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 Removing tags. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Removing your working copy. . . . . . . . . . . . . . . . . . . . 5 Renaming directories . . . . . . . . . . . . . . . . . . . . . . . . . . 57 Renaming files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Renaming tags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 Replacing a log message . . . . . . . . . . . . . . . . . . . . . . . 95 Reporting bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 Repositories, multiple. . . . . . . . . . . . . . . . . . . . . . . . . . 17 Repositories, remote . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Repositories, remote, port specification . . . . . 19, 21 Repository (intro) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Repository file, in CVS directory. . . . . . . . . . . . . . . 14 Repository, backing up . . . . . . . . . . . . . . . . . . . . . . . . 18 Repository, example . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Repository, how data is stored . . . . . . . . . . . . . . . . . . 8 Repository, moving . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Repository, setting up . . . . . . . . . . . . . . . . . . . . . . . . . 18 RereadLogAfterVerify, in CVSROOT/config . . 154 Reserved checkouts . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 Resetting sticky tags. . . . . . . . . . . . . . . . . . . . . . . . . . . 39 Resolving a conflict . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 Restoring old version of removed file . . . . . . . . . . . 47 Resurrecting old version of dead file. . . . . . . . . . . . 47 Retrieve a branch. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 Retrieving an old revision using tags . . . . . . . . . . . 35 Reverting to repository version . . . . . . . . . . . . . . . . 70 Revision keyword . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Revision management. . . . . . . . . . . . . . . . . . . . . . . . . . 73 Revision numbers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 Revision numbers (branches). . . . . . . . . . . . . . . . . . . 43 Revision tree . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 Revision tree, making branches . . . . . . . . . . . . . . . . 41 Revisions, merging differences between . . . . . . . . . 46 Revisions, versions and releases . . . . . . . . . . . . . . . . 33 Right-hand options . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 Root file, in CVS directory. . . . . . . . . . . . . . . . . . . . . . 7 rsh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 rsh replacements (Kerberized, SSH, &c). . . . . . . . 21 rtag (subcommand). . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 rtag, creating a branch using. . . . . . . . . . . . . . . . . . . 41

S

Saving space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 SCCS, importing files from . . . . . . . . . . . . . . . . . . . . 30 Security, file permissions in repository . . . . . . . . . . . 9 Security, GSSAPI. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Security, Kerberos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Security, of pserver . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Security, setuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Server, CVS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Server, temporary directories . . . . . . . . . . . . . . . . . . 28 Setgid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Setting up a repository . . . . . . . . . . . . . . . . . . . . . . . . 18 Setuid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Source keyword. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 Source, getting CVS source . . . . . . . . . . . . . . . . . . . . . 1 Source, getting from CVS . . . . . . . . . . . . . . . . . . . . . . . 4 Special files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 Specifying dates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 Spreading information . . . . . . . . . . . . . . . . . . . . . . . . . 67 ssh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 SSH (rsh replacement) . . . . . . . . . . . . . . . . . . . . . . . . . 21 Starting a project with CVS . . . . . . . . . . . . . . . . . . . 29 State keyword . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 Status of a file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 Status of a module . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Sticky date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 Sticky tags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 Sticky tags, resetting . . . . . . . . . . . . . . . . . . . . . . . . . . 39 Sticky tags/dates, per-directory . . . . . . . . . . . . . . . . 16 Storing log messages. . . . . . . . . . . . . . . . . . . . . . . . . . 148 Stream authentication . . . . . . . . . . . . . . . . . . . . . . . . . 88 Structure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 Subdirectories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 Support, getting CVS support . . . . . . . . . . . . . . . . 171 Symbolic link, importing . . . . . . . . . . . . . . . . . . . . . 116 Symbolic links . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 Symbolic name (tag) . . . . . . . . . . . . . . . . . . . . . . . . . . 34 Syntax of info files. . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 SystemAuth, in CVSROOT/config. . . . . . . . . . . . 153

T

tag (subcommand) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 Tag file, in CVS directory. . . . . . . . . . . . . . . . . . . . . . 16 Tag program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 tag, command, introduction. . . . . . . . . . . . . . . . . . . . 34 tag, creating a branch using . . . . . . . . . . . . . . . . . . . 41 Tag, example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 Tag, retrieving old revisions. . . . . . . . . . . . . . . . . . . . 35 Tag, symbolic name . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 taginfo (admin file) . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 Tags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 Tags, deleting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Tags, logging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 Tags, moving . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Tags, renaming . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 Tags, sticky . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 Tags, verifying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 tc, Trivial Compiler (example) . . . . . . . . . . . . . . . . . . 3 Team of developers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 TEMP, environment variable . . . . . . . . . . . . . . . . . 156 Template file, in CVS directory . . . . . . . . . . . . . . . . 16 Template for log message . . . . . . . . . . . . . . . . . . . . . 149 Temporary directories, and server . . . . . . . . . . . . . . 28 Temporary files, location of . . . . . . . . . . . . . . . . . . . 156 Third-party sources. . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 Time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 Timezone, in input . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 Timezone, in output . . . . . . . . . . . . . . . . . . . . . . . . . . 116 TMP, environment variable . . . . . . . . . . . . . . . . . . . 156 TMPDIR, environment variable. . . . . . . . . . . . . . . 156 TMPDIR, overriding . . . . . . . . . . . . . . . . . . . . . . . . . . 89 TopLevelAdmin, in CVSROOT/config . . . . . . . . 153 Trace. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 Traceability . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 Tracking sources . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 Transactions, atomic, lack of . . . . . . . . . . . . . . . . . . . 68 Trigger scripts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 Trigger scripts, security. . . . . . . . . . . . . . . . . . . . . . . 143 Trivial Compiler (example) . . . . . . . . . . . . . . . . . . . . . 3 Typical repository . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

U

Umask, for repository files . . . . . . . . . . . . . . . . . . . . . 10 Undoing a change . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 unedit (subcommand) . . . . . . . . . . . . . . . . . . . . . . . . . 70 Unknown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 Unreserved checkouts . . . . . . . . . . . . . . . . . . . . . . . . . . 63 Unresolved Conflict . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 Up-to-date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 update (subcommand). . . . . . . . . . . . . . . . . . . . . . . . 122 Update, introduction . . . . . . . . . . . . . . . . . . . . . . . . . . 64 update, to display file status . . . . . . . . . . . . . . . . . . . 64 Updating a file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 User aliases. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 User variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 USER, environment variable . . . . . . . . . . . . . . . . . . 153 USER, internal variable . . . . . . . . . . . . . . . . . . . . . . 152 users (admin file) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

V

val-tags file, and read-only access to projects … 10 val-tags file, forcing tags into . . . . . . . . . . . . . . . . . 164 Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 Vendor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 Vendor branch. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 ‘verifymsg’ (admin file) . . . . . . . . . . . . . . . . . . . . . . 144 ‘verifymsg’, changing the log message … 145, 154 version (subcommand). . . . . . . . . . . . . . . . . . . . . . . . 138 Versions, of CVS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 Versions, revisions and releases. . . . . . . . . . . . . . . . . 33 Viewing differences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 VISUAL, environment variable . . . . . . . . . . . . . 4, 155 VISUAL, internal variable . . . . . . . . . . . . . . . . . . . . 152

W

watch add (subcommand) . . . . . . . . . . . . . . . . . . . . . 69 watch off (subcommand). . . . . . . . . . . . . . . . . . . . . . . 69 watch on (subcommand) . . . . . . . . . . . . . . . . . . . . . . . 68 watch remove (subcommand) . . . . . . . . . . . . . . . . . . 69 watchers (subcommand) . . . . . . . . . . . . . . . . . . . . . . . 71 Watches . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 wdiff (import example) . . . . . . . . . . . . . . . . . . . . . . . . 79 Web pages, maintaining with CVS . . . . . . . . . . . . 149 What (shell command) . . . . . . . . . . . . . . . . . . . . . . . . 76 What branches are good for . . . . . . . . . . . . . . . . . . . 41 What is CVS not? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 What is CVS? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 When to commit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Windows, and permissions . . . . . . . . . . . . . . . . . . . . . 10 Work-session, example of . . . . . . . . . . . . . . . . . . . . . . . 3 Working copy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 Working copy, removing . . . . . . . . . . . . . . . . . . . . . . . . 5 Wrappers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 writers (admin file) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

X

Ximbiot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 xinetd, configuring for pserver . . . . . . . . . . . . . . . . . 21

Z

Zone, time, in input . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 Zone, time, in output . . . . . . . . . . . . . . . . . . . . . . . . . 116

180 CVS—Concurrent Versions System v1.11.23